A Bibliography of Books about the Structured Query Language (SQL)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

17 May 2023 Version 2.61

Title word cross-reference

*Forms [McC94].

.NET [Sun06].

10 [Intxxb]. 101 [All01]. 10g [Pri04, Urm04, KM07]. 11 [Ran96]. 13249-3 [Int99a]. 1950s [Hai09]. 1960s [Hai09]. 198X. [AC87]. 1995 [Int95a]. 1999 [MS02].

2 [CAE^{+}76, Gab07, M^{+}06]. 2.0 [EFH^{+}11, JH98]. 2000} [Abr03, Bru09, Buc01, Dal01, Del00, DS01, Des01, Eng01, Gri02, GR01, Hau01, H^{+}04, Ise01, LS01, Mar01, McM01, OC01, Pet00, Rab01, RP01, Sei01, Sha00, Sha01, Sha02, Spe02a, SW01, Sun03, TW06, Utl00, Vie00]. 2003 [DC03, DC06]. 2004 [Bar04]. 2005 [Ano06a, Ano07b, Bag06, Bea04, Car07, EC07, GD06, Gib08, GH06, Ham05, Ham06, Jor06, JWO09, KV07, K^{+}07, Lar06b, Lar06c, MKR07, MHTH07, M^{+}07, Mac07, M^{+}06, MM07, MTK06, Nan07, Nie07, Ote06, Pet05, Pet06, R^{+}07, R^{+}05, RDD^{+}06, Sha07, S^{+}06, Sun06, TM05, T^{+}06, Vee08, Vie06, Vie07, Wat06, Wil06, WLT07]. 2008 [AFB11, Col10, Dew08, Hot08, Hot09, LGMM09, LBF08, MCT08, Mis09, Ote10, Pet08, SSS10, Sta10, Ten10, T^{+}08, WF10]. 2016 [VCD16]. 24-hour [K^{+}10]. 250 [Gun05]. 2600 [Ano87, Ano90].

3 [For99]. 3.5 [Ham08]. 370 [Wad12].

4 [Gil03]. 4.1 [DuB05]. 400 [Hov95, MHJ95].

5 [Har04, Tra04]. 5/MySQL [Har04]. 555-1988 [Ame97]. 571 [AC87]. 571-D [AC87].

6 [YŞ01]. 6.5 [Laf97].

7 [Byr99, Cha96, CP99, Des00]. 7.0 [ES99, McG99, She00]. 7.2 [Sze02]. 70-029 [She00]. 70-431 [Ano06a, Car07, Gib08, Jor06, JWO09]. 70-432 [Hot09]. 70-443 [Gib08, MHTH07, Mac07]. 70-444 [Gib08]. 70-445 [Vee08].

8 [Blu07, GT06]. 8.1 [Lar06a]. 8i [DHMT00, YŞ01].

9 [RK10]. 9.0 [Smi10]. 9.1.3 [Ano06b]. 9075 [Intxxb]. 9075-1 [Int99b, ISO11a]. 9075-10 [ISO08a, Intxxb]. 9075-13 [ISO08b]. 9075-14 [ISO11b]. 9075-2 [Int99c, ISO11d, ISO11c]. 9075-3 [Int95c, ISO08c]. 9075-3-1995 [Ame95]. 9075-4 [Int99d, ISO11e]. 9075-4-1996 [Ame96]. 9075-5 [Intxxa]. 9075-9 [ISO08d]. 92 [Pas93]. 95 [Fre98]. 95/NT [Fre98]. 9579-2 [Int98]. 9579-2-1993 [Ame93]. 9579-3 [Int96].

absolute [Har03b, Har04]. Acceptance [Ame97]. Access [Ame93, All05, Bac09, Int96, Int98, San97, SC97, Int00, Ano04, Ano06b, SAS90, SAS93a, SAS93b, SAS94a, SAS94b, AN02, CB00, Sin00]. ACID [Reh03]. ACM [Ash87]. action [Ban12, Col10, Lam11]. active [DC03, DC06]. Ada [Int95b, Int95a]. Adabas [DHMT00, Mag09]. Admin [RK10]. Administering [Byr99, CP99, DD03, DD06, DuB03b, DuB05]. Administration [KL99, McG98, Morxxc, Wei04, AFB11, Ano07b, Cha96, Cho11, Col10, Gib08, Hau01, K^{+}07, Laf97, Lar06a, Lar07, LS01, Mal11, Rab01, She03, WLT07]. administrator [M^{+}07, MyS05, Sta10]. administrators [Spe02a]. ADO.NET [Ham08, M^{+}06]. Advanced [Bur03, Dav99, Feu96, GS02a, Mel03b, Reh03, Cel95, Cel05a, MKR07, Ora90a, Urm00, Wel92]. AG [Mag09]. aggregates [CH07, Kie84]. al [Lah05]. all-in-one [Cog04]. almanach [KK02]. America [Mag09]. Analysis [S^{+}06, Ten10, HQ06, Lin07, Lin08, TL01, Wün13]. analytics [Cel06, HDA11]. annotated [LC99]. ANSI [Ame89, Ame92, Ame93, Ame95, Ame96, Ame97, Dav99]. ANSI/EIA [Ame97]. ANSI/ISO/IEC [Ame93, Ame95, Ame96]. answers [Cel07]. antipatterns [Kar10]. Anwendungsentwicklung [Boe03]. Apache [Cog04, Gab07, G^{+}06, Gla04, HJL00, HG00, LW03, Mel03a, Mel04, Nar05, Reh03, TSC02]. API [JRC97]. APIs [Eis05]. aplikacjami [Feh03]. Applets [Küh97]. Application [Int99a, Gab07, MD92, Ora90b, Say89a, Say90a, Say90c, Utl00, WF10]. Applications [McC98, WL01, AO10, App02, Bac07, Buc01, BGWG04, FL08, GB01, Kho92, LD91, MM07, Ode99, Sno00, VCD16, WL02, WL04, Yad07]. Applied [dHK07]. Applikationen [GGK99]. Applixware [GGK99]. Approach [ABC^{+}76, CAE^{+}76]. apps [Val05]. Arbeiten [GGK99]. architectures [VB07]. archives [LC99]. art [FR06]. AS/400 [Hov95]. ASP [Bar04, Wei00]. ASP.NET [Kie09]. attackers [Che15]. Ausfallsicherheit [DHMT00]. ausführlicher [Eis03]. Automated [Che91]. automation [RC04]. auxiliary [Cel08]. av [HJL00]. Availability [IBM09a, H^{+}04, Ote10]. avoiding [Kar10]. Award [Ash87].

BackOffice [Mic99]. Backup [DHMT00, Pre07, Des00, Des01, McB02]. backups [ZB04]. balancing [ZB04]. Banks [Cod70]. Base [Cod71, Cod72, Hai09, Dup05, Lar06a, Lar07]. based [CO93, PP98, Per98c, SSS10, Che91]. basert [HJL00]. basics [Pas90]. bazodanowymi [Feh03]. Beans [Küh97]. beginner [DeB18, Har03b, Har04, Hou01a, Pet00, Pet05, Pet06, Pet08, SH03]. Beginning [Bal07a, Bal07b, DB05, DBB06, DB08, Dar12, Dew07, Dew08, Gil06, GT06, Gla04, MS05, Nar05, SM05, SM01, TW06, Vie06, WF10, WC05, WLT07]. beherrschen [Eis03]. Berkeley [Yad07]. Beschreibung [Eis03]. best [Ano07a, Feu01a, Feu01b, VB07]. bestselling [Feu00a]. bible [CP04, KT03, KT08, Nie07]. Bibliotheken [Küh97]. Bildbearbeitungsprogramm [GGK99]. Bindings [Intxxa, ISO08a, Intxxb]. biodata [Wün13]. biology [Wün13]. Black [Whi06, Dal01]. blueprint [Mon03, Val06]. błyskawiczny [Feh03]. Book [Abr03, Whi06, Ber04, Dal01, GMUW02, Yad07]. books [Feu00a]. Bookshelf [O’R00, Feu00a, FWS^{+}03, SEG^{+}03]. brief [SF13]. Buch [GGK99]. Build [Yan04]. Building [Owe96, Owe98, Rai08, Bac07, DD03, DD06, Kie09, Stu01]. Built [FDB98, FB98, FBD98]. Built-in [FDB98]. built-ins [FB98, FBD98]. Büropakete [GGK99]. Business [Bra88, K^{+}10, MTK06, LGMM09, Lar06b, Mar05, Vee08].

C [Tar89, WL91]. C-ISAM [Tar89]. CA [Kru09]. CAE [X/O94]. Calculus [Cod71]. California [CK12]. Call [Ame95, Int95c, ISO08c]. Call-Level [Ame95, Int95c, ISO08c]. cascading [Per98a, Per98b]. Cassandra [Hew10]. CD [Feu00a, FWS^{+}03, O’R00, SEG^{+}03]. CD-ROM [Feu00a]. Celko [Cel95, Cel05a, Cel05b, Cel06, Cel07, Cel08]. centric [Wel01]. certification [DHP04]. changes [SAS94b]. CICS [Bra88]. Cincom [Nie09]. Claude [Abr03]. clearly [Har98a, Har03a]. CLI [DD98, Int95c, ISO08c, Ame95, DD98]. Client [Eis05, San97, SC97, Kho92, Pur94]. Client-APIs [Eis05]. Client/server [San97, SC97, Kho92, Pur94]. CLING [Cha90, CR90]. CLING/SQL [Cha90, CR90]. Cloud [PMH10, VCD16]. CLR [CH07]. cluster [Cho11]. clustering [DF06, Ote10]. COBOL [Bra88, CM90]. Code [Hol04, Wel01]. Code-centric [Wel01]. codes [GN06]. ColdFusion [Bar04]. combining [Tho03]. commands [GN06]. commerce [DB05, DBB06, DB08]. Commercialization [BH09]. Common [Cha90, CR90]. community [Gar09]. compact [AO10]. compatibles [Bak89]. Compiling [Wad12]. Complete [Sip98, GMUW02, GW99, GW02, Gro02, MS93, Sha00, Sha01, Sha07, Syb01, WGOD10, vdL89]. Completeness [Cod72]. Compliance [Ame97]. Components [Mel01, MS02]. Comprehensive [JRC97, DD03, DD06, Hes07, vdL07b]. Computational [Wün13]. Computing [PMH10, ES96, ES99]. Concept [Gri12]. Conception [Dup05, Pat09]. Concepts [Mom01b, GR93]. Conformance [Int96]. consultant [Sta10]. content [Tho03]. context [MB06]. Control [CAE^{+}76, AO10]. conversion [Che91, C^{+}04, C^{+}05]. Cookbook [DuB03a, RK10, SG02, DuB06, DuB07, Ham08, Mol06, Spe02b]. Core [Atk02, M^{+}07]. Corporation [Bru09, Row12]. correlation [Kie84, Per98a, Per98b]. course [Gre02, Tay01]. crash [Gre02, Tay01]. create [BP08]. create-modify-reuse [BP08]. Creating [McC98, MM07, Val06, Yad07]. critical [VCD16]. CryptSQLite [WSS^{+}20]. ćwiczenia [Sze02].

D [AC87]. Data [Abr03, Bac09, Bad20, BC73, BCKH75, CAE^{+}76, Cod70, Cod71, Cod72, Hai09, ISO08d, Lin07, Lin08, MCT08, MTK06, Pet19, Sei01, TM05, Tar89, TL01, WSS^{+}20, BG04, Dav99, DeB18, HDA11, HV00, O’G98, Rai08, VH08, VH14]. Database [Ame89, Ame93, Ame95, Ame96, Ano87, Ano90, ABC^{+}76, BH09, Bur03, Dat05, FIP90, FIP93, GMUW02, Gra12, Gri12, Int92, Int95c, Int96, Int98, Int99b, Int99c, Int99a, Int99d, Intxxa, ISO08c, ISO08d, ISO08a, ISO08b, ISO11a, ISO11d, ISO11e, ISO11c, ISO11b, KL99, Kru09, Lus88b, McC98, Nat90b, O’N94, OO01, Pet19, PMH10, Pri04, RG03, Ree97, Sip98, Sto88, Uni87, Urm04, WC12, WL01, Ame92, Bak89, Bre91, BGWG04, Che91, Che15, Cod87, Dat75, Dat77, Dat81, Dat86, Dat90, Dat95, DD97, Dat00, Dat04, ES96, ES99, For99, Gib08, GMT95, GB01, Int00, Kar10, Kuz10, Lah05, Lus88a, MKR07, MHTH07, M^{+}07, Mac07, Mal11, MD92, McM01, Met01, Mon03, New92, Ote10, Pet05, PB93, Rio05, RTRT94, Sno00, Spe02a, Val06, Wal00, WL02, WL04, Yad07, Yan04, dHK07]. database [vdL07a, Intxxb]. Databases [SM01, dR10, All03, AO10, DD03, DD06, EFH^{+}11, MS05, Owe96, Owe98, Pas93, Per98a, Per98b, RW12, She00, Van91]. Datacom [Ste09]. Datacom/DB [Ste09]. Datenbank [GGK99]. Datenbankanwendungen [Klu98]. Datenbanken [DHMT00, EFH^{+}11]. Datentypen [Eis05]. days [But03, LAG97, MB00a, MSM00, Mau01, MB00b, MS97, McG98, Ste00, SP00, WS99]. DB [Ste09, Yad07]. DB2 [C^{+}05, DHMT00, And13, Boh13, C^{+}04, FS90, HS13, IBM09a, Lah05, Mar89, MH89, Say89a, Say90a, Say90c, Say91, SG91, Wip89]. DB2/SQL [Mar89, MH89]. DBA [RC03, Aul01, Ber03, BG04, RC04, SS03]. DBAs [AFB11, Mal11, Nan05, NF05]. dBASE [SH90, vdL91, Car92, Car93, HH89].

DBMS [Bac09, Lar91, Lar93]. Debian [Her04]. Decision [Abr03]. Deep [Pet19, N^{+}10]. defenses [Nys07]. Defensive [Kuz10]. definite [Yad07]. Definition [BC73, CAE^{+}76, ISO11c]. Definitive [Geo11, Owe06, PMH10, AO10, CD10, DuB03b, DuB05, Gar09, Gen99, GM01, Gen05b, Hew10, KK05, Whi12]. Delivering [Lar06b]. demystified [Opp05]. depper [VCD16]. depth [Dat05]. Description [Int95a, Int95b]. Design [Hol04, CM90, Gab07, GMT95, Lah05, McM01, RP01, Say90a, Wal00]. Designing [Rio05, She00, MHTH07, Mac07]. desk [Cog04, Lan04a, Lan04b]. Desktop [PMH10, Wei00, KK01]. Detection [Reh03]. Developer [Atk02, FO00a, GS01b, McC98, CB00, Feu00b, FO00b, GS01a, Hot07, HH91a, HH92, Mar01, MCG03, McC96, Ode99, OC99, OC01, Ote06, Utl00, YŞ01]. Developers [Dew07, And13, Bea04, Dew08, DS05, EC07, KM07, San97, SC97, Say89a, Say90c, vdL07b]. Developing [Pur94, Sno00, App02, Kho92]. Development [Boh13, Morxxa, Che91, Cog04, Gar09, G^{+}06, Gla04, Klu98, LW03, MKR07, MD92, Nar05, Wal00, WL91, WF10, WT01]. développeur [GS^{+}02b]. Direct [Bac09]. Direct-Access [Bac09]. directory [DC03, DC06]. DIS [Intxxb]. Disaster [IBM09a]. distilled [SF13]. Distributed [Dye99, Pet19, WL91]. Dive [Pet19]. dives [N^{+}10]. dla [Har98b, Har00, Har05]. DMO [Mit03, MA03]. DMX [Ten10]. do [Vas05]. Documentation [WA02]. DOM [McC05]. données [Dup05, Lar06a, Lar07]. Draft [AC87]. Dreamweaver [Bar04]. driven [Val06, WF10, Yan04]. DS [DW89, FS90, Gil91a, Gil91b, HN13, SAS90, Say89a, Say90c, SG91, Wip89]. dummies [Cog04, Tay10, Val05, Val07, Wat06]. d’une [Dup05, Lar06a, Lar07]. Dynamic [IBM09b, Bac07, McC05, Tho03, Val06]. e-commerce [DB05, DBB06, DB08]. Early [Boh13, Cha12, WC12]. easy [BG04]. easy-to-find [BG04]. edition [Col00a, GD06]. editor [Gra12]. effective [Rio05]. effektives [Eis05]. Efficient [McC98, Dew08]. EIA [Ame97]. eight [Bac07]. Einführung [Mom01a]. Einsatz [Wei04]. Einstieg [EFH^{+}11]. embed [AO10]. Embedded [Say90a, Say91, Yad07, Tar89]. Embedded-SQL [Tar89]. embedding [AC87]. emerging [SF13]. emphasizing [Gar09]. Empress [DHMT00]. English [BC73, CB74]. enhancement [Int00]. enhancements [SAS94b, Sha02]. enough [Ano08]. enterprises [Sto11]. entity [Gog94]. entity-relationship [Gog94]. Entry [EFH^{+}11]. Entwickler [Eis03]. Entwicklungssoftware [Klu98]. environment [Tar89]. Environments [FIP95, Ano95, Com95]. Era [Dar12]. erfolgreiches [GGK99]. Erweiterungen [DD98]. Essbase [S^{+}06]. Essential [Cop08, KL99, Sti01, GN06, HG04, Kre00a, Kre00b, Pet05, Urm97]. essentials [BR01, Ham05, Ham06]. every [Urm97]. everyday [Val05]. everything [Vas05]. Exam [MHTH07, Mac07, She00, Ano06a, Car07, Gib08, Hot09, Jor06, Vee08]. Example [RS04, GR01, Ris10]. examples [Pas93, PMD05, Rai08, VB07]. exams [Gib08]. Excel [Lin07, Lin08]. executive [Mar05]. Expert [Bel07, HM05, KV07, Lus88b, MKR07, Lus88a]. explained [Har98a, Har03a]. exploitation [Lar06a, Lar07]. express [GD06, Bal06, Dew07]. Expressions [BCKH75]. extended [Gog94]. extension [PP98, Per98c]. Extensions [DD98, RTRT94]. External [ISO08d].

facilities [MD92]. Facility [BC73, SAS93a, SAS94a, SAS94b, SG91]. Fast [McC98, Kie09]. fault [Yad07]. fault-tolerant [Yad07]. FDBC [Sip98]. features [Feu99b, Feu99c, Mel03b, Sha02]. featuring [Cro88, Cro89]. Fibel [Küh97]. file [Bra88]. find [BG04]. First [Bac09, HN13, Ash87, Bea04, Bei07]. flexibility [Gar09]. formation [Cha96]. Formative [Gra12]. forms [KM07]. FORTRAN [Che91]. FORTRAN-based [Che91]. Forum [Kop04]. Foundation [Int99c, ISO11d, Cod87, Ora90c, Ora90d]. Founded [Cod71]. Framework [Int99b, ISO11a, Gar09, O’G98]. functions [CH07, J^{+}05]. Fundamentals [Pat99, BG08, Gog94, MTC00, Pat02]. Funktionen [Eis05].

ge [Eis05]. ge-packt [Eis05]. geographic [RTRT94]. German [DD98, EFH^{+}11]. GIMP [GGK99]. Got [Hai09]. Growth [HS13]. Grundlagen [Boe03]. Guest [Gra12]. Guide [Atk02, Bak89, Cha96, Gen02, Geo11, JRC97, Kho92, Loc00, Lus88b, Owe06, PMH10, Sip98, Wip89, AO10, Car07, C^{+}04, C^{+}05, CB00, CD10, Dat87, DW89, Dat89, DD93, DD97, DT02, DeB18, DC03, DC06, DS05, DD03, DD06, DuB03b, DHP04, DuB05, ES96, ES99, Feu99b, Feu99c, Gar09, Gen99, GM01, Gen04b, Gen05b, Gen06, GS^{+}02b, Gib08, GG89, Har92, HM92, HL95, Hen00, Hen02, HV00, Hes07, Hew10, Hou01a, HH89, HH91a, HH92, Inf95a, Inf95b, Inf00a, Inf00b, Inf00c, JH98, Jor06, KK05, Laf97, Lus88a, Mar01, Mar89, MH89, MHJ95, MCG03, McC96, MD92, MS93, ME00, MyS05, Nat90a, OC99, OC01, Ote06, Pet00, Pet05, Pet06, Pet08, Pra91, Pra05, PL06, RW12, SAS89, SAS95, Sac87, SF13, San97, SC97, Say91]. guide [Sha02, SH03, SS03, SH92, Urm97, Utl00, VB07, VH08, VH14, WL91, Whi12, Yad07, YŞ01, vdL91, vdL92, vdL98]. guidebook [Spe02a]. guru [Hen00, Hen02].

hacks [CR07, Cum07]. Hadoop [Lam11, Whi12]. Handbook [BED96, Bur03, GS01b, BED93, BED01, BCE03, CO93, EDB89, ES96, Eng97, ES99, Eng01, GS01a, Hot07]. Handbuch [Eis03]. Hands [GG97, Nan07, HV00, VH08, VH14]. Hands-on [GG97, Nan07, HV00, VH08, VH14]. Happened [Deu13]. HBase [Geo11]. Head [Bei07]. here [Say89b]. High [IBM09a, SZT^{+}08, Smi10, WSS^{+}20, ZB04, Ano07a, Bur01, Har01b, H^{+}04, Ote10]. high-performance [Bur01, Har01b]. History [Cha12, Eps13, HS13, Kru09, Row12, Ste09]. Hitchhiker [VB07]. Hive [CWR12]. hjelp [HJL00]. Holub [Hol04]. Host [Intxxa]. hour [K^{+}10]. hours [Bal06, Mel02, Mel03a, SP98, SPJ08]. HTML [Hen02]. Hyperion [S^{+}06]. hyperscale [VCD16].

IBM [Bak89, C^{+}05, HS13, HN13, Lar91, Lar93, WC12]. IDL [Küh97]. IDMS [Kru09]. IDS [Bac09, Reh03]. IEC [Ame93, Ame95, Ame96, Int92, Int95a, Int95c, Int95b, Int96, Int98, Int99b, Int99c, Int99a, Int99d, Intxxb, Intxxa, Int00, ISO08c, ISO08d, ISO08a, ISO08b, ISO11a, ISO11d, ISO11e, ISO11c, ISO11b]. ile [Yaz06]. im [GGK99]. Implementation [Int96, Ano06a, Car07, DC03, DC06, Gab07, Hot09, Jor06, JWO09, Vee08]. implementing [For99, She00]. improvement [Gil91a, Gil91b]. IMS [McG09, Pat09]. Inc. [MD92]. includes [Ame89, Gar09]. index [Lah05]. Industry [CK12]. Information [Ame89, Ame93, Ame95, Ame96, Gri12, Hai09, Int92, Int95c, Int95b, Int96, Int98, Int99b, Int99c, Int99a, Int99d, Intxxa, Int00, ISO08c, ISO08d, ISO08a, ISO08b, ISO11a, ISO11d, ISO11e, ISO11c, ISO11b, Sip13, Ame92, RTRT94, Urm97, Int95a, Intxxb, McG09]. Informix [DHMT00, IBM09b, LT91, Lef91, Tar89, HL95, Mar05, McN97, SAS94a, Sip13]. INFORMIX-SQL [LT91, Lef91, Tar89]. infrastructure [MHTH07, Mac07]. INGRES [Cha90, CR90, HH92, Car92, Car93, Row12]. INGRES/SQL [Cha90, CR90]. injection [Nys07]. ins [DHMT00, FB98, FBD98]. Insider [Eis03]. Insider-Wissen [Eis03]. insights [VCD16]. Installation [DHMT00]. Instant [Buc01]. Integrated [Bac09, San97, SC97]. integrating [TSC02]. integration [Ham07, KV07, Nan07]. Intelligence [MTK06, LGMM09, Lar06b, Vee08, K^{+}10]. intelligent [Owe96, Owe98]. Interactive [Hov95, Ros00, RR00]. Interconnection [Int96, Int98]. Interface [Ame95, Int95c, ISO08c, Cha90, CR90, SAS90, SAS93a, SAS93b, SAS94a]. interfaces [Stu01]. Internals [Pet19, BCE03, Pac06]. Internationalization [Kap01]. Internet [Klu98]. Intranet [Klu98]. Introducing [VCD16]. Introduction [All03, Mom01b, vdL88, vdL93, vdL07a, CT02, Dat75, Dat77, Dat81, Dat86, Dat90, Dat95, Dat00, Dat04, DT02, Din94, Gra12, McC94, Ode99, TC89, Wün13]. Intrusion [Reh03]. ISAM [Tar89]. ISO [Ame93, Ame95, Ame96, CO93, Int92, Int95a, Int95c, Int95b, Int96, Int98, Int99b, Int99c, Int99a, Int99d, Intxxb, Intxxa, Int00]. ISO/IEC [Int92, Int95a, Int95c, Int95b, Int96, Int98, Int99b, Int99c, Int99a, Int99d, Intxxb, Intxxa, Int00, ISO08c, ISO08d, ISO08a, ISO08b, ISO11a, ISO11d, ISO11e, ISO11c, ISO11b]. issues [PB93]. IV [HH89, Car93, SH90, vdL91].

jake [HCxx]. Java [DHMT00, ISO08b, Klu98, Küh97, HJL00, MCG03, McC98, ME00, Pri01, Ree97, Sip98, Tho03]. JavaScript [Gab07, Gar09]. JDBC [JRC97, Klu98, ME00, Ree97, Sip98]. JDeveloper [KM07]. JDK [Klu98]. jezyka [KK02]. job [BG04]. Joe [Cel95, Cel05a, Cel05b, Cel06, Cel07, Cel08]. JRT [ISO08b]. JSDK [Klu98]. JSP [Har01a]. jumpstart [PMD05]. Just [Ano08].

każdego [Har98b, Har00, Har05]. kit [Ano06a, GD06, Hot09, Klu98, MHTH07, M^{+}07, Mac07, Mic99, Rab01, Vee08]. Klient [HJL00]. Knight [K^{+}10]. Konfiguration [DHMT00]. Konzepte [Mom01a]. kurs [Feh03].

LAMP [G^{+}06, LW03]. Language [Ame89, BC73, Din94, FIP90, FIP93, Int95b, Intxxb, Intxxa, ISO08a, ISO08b, Mel01, Wad12, Ame92, Ano87, Ano90, Ano05, BED93, CB74, Dat87, Dat89, DD93, DD97, DuB04, Feu99a, FPD99, Feu03, FPD03, FPD04, FPD08, GG97, HH88, HH91b, MS02, Met01, Nat90b, New92, Per98a, Per98b, PB93, Uni87, vdL07a, BED96, EDB89, Int95a]. Languages [Ame95, AC87, Ame96, Int92, Int95a, Int95c, Int95b, Int99b, Int99c, Int99a, Int99d, Intxxb, Intxxa, ISO08c, ISO08d, ISO08a, ISO08b, ISO11a, ISO11d, ISO11e, ISO11c, ISO11b]. Large [Cod70]. leadership [Mar05]. Learn [Dic02, Bal07b]. Learning [Bag06, Bea05, DP07, Hol04, PF02, TW07, Wel91, Wel92]. Lectures [Ash87]. Lessons [Mar05]. Level [Ame95, Ame97, Int95c, ISO08c]. library [Ise01, Yad07]. lightweight [Gar09]. like [Kie84]. limited [PP98, Per98c]. LINUX [GGK99, DW06, DHMT00, Wei04, Dup05, FWS^{+}03, Gab07, G^{+}06, IBM09a, IBM09b, LW03, MS^{+}00, TSC02, Wal00, Wün13]. LINUX-Applikationen [GGK99]. LINUX-Büropakete [GGK99]. Linux/Unix [Gab07]. LISP [Cha90, CR90]. load [ZB04]. look [Bea04]. Looking [Hol04]. løsning [HJL00]. Lot [Ame97].

LPP [Per98a, Per98b, PP98, Per98c].

Machine [Wad12]. machines [Has96]. Macromedia [Bar04]. maintaining [O’G98]. maintenance [Ano06a, Car07, DC03, DC06, Hot09, Jor06, JWO09, Vee08]. Management [ABC^{+}76, BH09, Gra12, ISO08d, Kru09, McG09, RG03, Sip13, Bak89, Che91, MD92, Mic97, Mon03, SG91, SSS10]. Managing [Boh13, RYW02]. Manipulation [CAE^{+}76, HV00, VH08, VH14]. Manual [Morxxb, WA02, Pos07]. marketplace [GG97]. Mastering [Cro88, Cro89, Gil03, Gun05, GH06, dH05, McD03, MB02, MB04, dHN05, vdL07a]. mathematics [dHK07]. MCITP [Gib08, MHTH07, Mac07]. MCSE [Byr99, DC03, DC06, Hau01, Laf97, McG98, McM01, Rab01, She00]. MCTS [Ano06a, Car07, Hot09, Jor06, Vee08]. MDX [HQ06, SC09, S^{+}06]. MED [ISO08d]. mere [HV00, VH08, VH14]. Microsoft [All05, AN02, AFB11, Ano06a, Bal06, BG08, Byr99, C^{+}05, CB00, Del00, DS01, ES96, ES99, Eng01, GR01, Gun05, GH06, HDA11, H^{+}04, Hot08, Hot09, Ise01, Jor06, JWO09, K^{+}10, LGMM09, Lar06b, Lar06c, LBF08, MCT08, MHTH07, M^{+}07, Mac07, Mal11, Mar01, McG99, Mis09, MTK06, Nan07, Ote06, Ote10, PS99, Pet05, Pet06, Pet08, Rab01, R^{+}07, R^{+}11, Sei01, Sha07, Sha02, She00, SC09, Sol96, SR98, SS03, Spe02a, S^{+}06, Sta10, Sun06, Ten10, Ten11, T^{+}08, VCD16, Vee08, Wat06, WS99, Wyn99]. Migrating [IBM09b]. Mini [JH98]. Mining [Abr03, MCT08, Sei01, TM05, TL01]. MiniSQL [DHMT00]. minutes [For13]. mirroring [Ote10]. mission [VCD16]. mission-critical [VCD16]. Model [Cod70, Dar12, Bal07b, Gog94, WF10]. model-driven [WF10]. modeling [Dav99, WF10]. modern [RW12]. modify [BP08]. Module [Int95a, Int95b]. Modules [Ame96, Int99d, ISO11e]. MongoDB [Ban12, CD10, Cho11, PMH10]. mortals [HV00, VH08, VH14]. movement [RW12]. MRI [Bru09]. mSQL [YRK99]. Multimedia [Int99a]. MVP [N^{+}10]. MX [Bar04]. MySQL [DHMT00, HG00, Ano06b, Ano08, DuB04, Mel04, Ano04, Ano05, Atk02, Bac07, Bel07, BP08, BR01, BGWG04, But03, C^{+}04, Cog04, CP04, DB05, DB08, DF06, DP07, DuB99, DuB01, DuB02, DuB03a, DuB03b, DHP04, DuB05, DuB06, DuB07, DuB13, Dye05, Dye08, Gab07, G^{+}06, Gil03, Gil06, Gla04, GN06, GB01, Gre02, HJL00, Har01a, Har03b, Har04, HF06, HG04, Kof01, Kop04, LW03, MB00a, MCG03, Max00, Max01, MB00b, M^{+}06, Mel02, Mel03a, Mon03, Nar05, Pac06, PL06, RYW02, Ree03, Ree07, Reh03, SZT^{+}08, SM05, Stu01, TW07, TSC02, Val05, Val06, Val07, Vas05, WT01, WA02, WL01, WL02, WL04, Wün13, Yan04, YRK99, ZB04, vdL07b]. MySQL/PHP [BGWG04, GB01]. MySQL4 [DuB03b]. MySQL5 [KK05].

Nachschlagen [Eis05]. neuen [Küh97]. nichtrelationaler [EFH^{+}11]. nonrelational [EFH^{+}11]. North [Mag09]. Northern [CK12]. NoSQL [EFH^{+}11, PMH10, RW12, SF13, Sto11, Tiw11]. Novice [Dew07, Bal07a, Bal07b, DB05, DBB06, DB08, Dew08, Gil06, GT06, MS05]. NT [Fre98]. Nutshell [Wei00, Dye05, Dye08, KK01, KKH04, KHK08].

Object [ISO08a, For99, Mel03b, Intxxb]. object-relational [For99, Mel03b]. OCA [DT02]. OCA/OCP [DT02]. OCP [DT02]. ODBC [SAS93a, Sip98]. Office [GGK99]. Official [JH98]. offizielle [Eis03]. OLAP [Cel06]. OLB [ISO08a, Intxxb]. on-the-job [BG04]. one [Cog04, Mel04]. Open [Int96, Int98, LW03, Mon03, Yad07]. Operatoren [Eis05]. opis [KK02]. optimalen [DHMT00]. Optimization [Has96, Eng97, Eng01, ZB04]. optimizer [Aul03]. optimizers [Lah05]. Optimizing [GP94, MH03]. Options [IBM09a]. Oracle [DHMT00, Urm97, vdL92, CG93, Cro89, Har92, All01, AFB11, Aul03, Bur01, BCE03, Car92, Car93, Cha96, Cro88, Dye99, Feu95a, Feu95b, Feu96, Feu97a, Feu97b, FDB98, FB98, FBD98, Feu99a, Feu99b, FPD99, Feu99c, Feu00b, FO00a, FO00b, Feu00a, Feu01a, Feu01b, Feu02, FP02, Feu03, FPD03, FPD04, FP05a, FP05b, FPD08, Gen99, Gen00, GM01, Gen02, Gen05b, Gen05a, GC96, Gur02, dH05, HM05, Har01b, HM92, Hot07, HH91a, IBM09b, KM07, KL99, Kre00a, Kre00b, Lah05, LK98, LC99, Mal11, McC96, M^{+}06, McC94, McD03, Men13, MH03, MB02, MB04, Nan05, NF05, O’R00, Ode99, Owe96, Owe98, PMD05, Pre12, PF02, Pri01, Pri04, Ris10, Ros00, RR00, RS04, RC03, RC04, Say90b, Tho03, Tre99]. Oracle [Urm96a, Urm96b, Urm04, WL91, YŞ01, dHN05, vdL98]. ORACLE/SQL [Har92, HM92]. Oracle8 [PH98, Urm97]. Oracle8i [Feu99b, Aul01, Feu99c, Urm00]. Oracle9i [DT02]. oriented [Sno00]. Origin [Bac09]. other [Mel03b]. outline [MTC00]. outlines [Aul03]. own [Yan04].

P [Bre91]. P-SQL [Bre91]. paced [Ano06a, Hot09, MHTH07, M^{+}07, Mac07, Vee08]. Packages [FDB98, Int99a, Feu96]. packt [Eis05]. parallel [Has96, O’G98]. Part [Ame93, Ame95, Ame96, Int95c, Int96, Int98, Int99b, Int99c, Int99a, Int99d, Intxxa, ISO08c, ISO08d, ISO08a, ISO08b, ISO11a, ISO11d, ISO11e, ISO11c, ISO11b, Pre12, Intxxb]. pass [SAS93a, SAS94a, SAS94b]. pass-through [SAS93a, SAS94a, SAS94b]. path [Dew08]. patience [PP98, Per98c]. Patterns [Hol04, PP98, Per98c]. PC [Ste05]. PCs [Bak89]. Performance [CG93, Eis05, GC96, Smi10, Ano07a, Bur01, Eng97, Eng01, Gil91a, Gil91b, GP03, Har01b, MH03, O’N94, OO01, SZT^{+}08, Urm97, ZB04].

Perl [DuB01, DuB02, LW03, She03]. persistence [SF13]. Persistent [Ame96, Int99d, ISO11e]. Perspective [CK12]. Phil [Mar05]. PHP [DHMT00, Ano08, Bac07, Bar04, Boe03, BP08, BGWG04, Cog04, DB05, DBB06, DB08, DP07, G^{+}06, GS02a, Gil06, GT06, Gla04, GB01, HG00, Har03b, Har04, Kop04, LW03, Mar02, Max00, Max01, Mel03a, Mel04, Reh03, Tra04, Val05, Val06, Val07, Vas05, WT01, WL01, WL02, WL04, Yan04]. PHP/MySQL [Har03b, Kop04]. PHP/MySQL/Apache [HG00]. PHP/XML [Mar02]. PHP5 [CP04, Gab07, Nar05]. phrasebook [GN06]. Physical [GMT95]. PICS [Int96]. pitfalls [Kar10]. PL [All01, Feu00a, O’R00, Aul01, Bal07a, Bal07b, Feu95a, Feu95b, Feu96, Feu97a, Feu97b, FB98, FBD98, Feu99a, Feu99b, FPD99, Feu99c, Feu00b, FO00a, FO00b, Feu01a, Feu01b, Feu02, FP02, Feu03, FPD03, FPD04, FP05a, FP05b, FPD08, HM05, KM07, LC99, LAG97, McC94, McD03, Nan05, NF05, Ode99, Owe96, Owe98, PF02, Ros00, RR00, RS04, Say91, Tre99, Urm96a, Urm96b, Urm97, Urm00, Urm04, YŞ01]. PL/I [Say91]. PL/SQL [All01, Feu00a, O’R00, Aul01, Bal07a, Bal07b, Feu95a, Feu95b, Feu96, Feu97a, Feu97b, FB98, FBD98, Feu99a, Feu99b, FPD99, Feu99c, Feu00b, FO00a, FO00b, Feu01a, Feu01b, Feu02, FP02, Feu03, FPD03, FPD04, FP05a, FP05b, FPD08, HM05, KM07, LC99, LAG97, McC94, McD03, Nan05, NF05, Ode99, Owe96, Owe98, PF02, Ros00, RR00, RS04, Tre99, Urm96a, Urm96b, Urm97, Urm00, Urm04, YŞ01]. plan [Aul03]. planning [DC03, DC06]. Pocket [Gen02, Gur02, FB98, FBD98, Feu99a, FPD99, Feu03, FPD03, FPD04, FPD08, Gen00, Gen04b, Gen05a, Gen06, Ree03, Ree07, Sta10]. poleceń [KK02]. policy [SSS10]. policy-based [SSS10].

Polish [Sze02, WDM02]. polyglot [SF13]. Portable [Ber03, BG04, Bre91]. Postfix [Blu01]. PostgreSQL [DHMT00, Eis03, GGK99, Klu98, Ber04, Blu07, Boe03, DBB06, DD03, DD06, Dup05, Eis03, Eis05, GS01a, GS01b, GS^{+}02b, GS02a, GT06, Har01c, HCxx, OH12, Lar06a, Lar07, Loc00, Mar02, MS05, Mom01a, Mom01b, Morxxa, Morxxb, Morxxc, Per01, POS02, RK10, Smi10, Sti01, SM01, Sze02, Wal00, Wei04, WD02c, WD02a, WDM02, WD02b, Yaz06]. PostgreSQL-Entwickler [Eis03]. PostgreSQL-Server [Eis03]. power [Hes07]. powerful [AO10]. PPM [Ame97]. Practical [Bac07, BED96, Bow01, DeB18, McD03, Ten10, Ten11, WD02c, BED93, BED01, Cod87, Din94, DS05, EDB89, Wün13, YŞ01, WDM02]. practice [VB07]. practices [Feu01a, Feu01b]. practitioners [Dat05]. pragmatics [Gog94]. praktisk [HG00]. praktyczne [Sze02]. praktyczny [WDM02]. pratique [WD02a, WD02b]. Praxis [Boe03, Klu98]. praxisnah [Har01c]. principles [O’N94, OO01]. Pro [R^{+}05, RDD^{+}06, SSS10]. Procedure [Ame97, HF06, Owe96, SAS89, SAS93a, SAS94a, SAS94b, SW01, Sun03, Sun06]. procedures [CH07, Hen02, Owe98, Wel01]. processing [Bra88, Dav99, GR93, Tar89, Wün13]. Product [Boh13, McG09, MD92]. productivity [Cod87]. Professional [Dew07, G^{+}06, MS^{+}00, M^{+}06, SH90, Tiw11, T^{+}08, Bal07a, Bal07b, DB05, DBB06, DB08, Dew08, Gil06, GT06, Har92, HM92, Mar01, Mar89, MH89, MHJ95, MS05, Ano07b, CH07, HQ06, Kle06, K^{+}07, T^{+}04, T^{+}06, Vie00, Vie07]. professionals [MKR07, Pet05, dHK07]. professionell [Har01c]. professionnels [Mar02]. proficiency [Dew08]. proforma [Int96]. Program [McG09]. Programlama [Yaz06]. Programme [Eis05]. Programmer [Bur03, Loc00, Bha98, Fre98, GG89, Har92, HM92, J^{+}05, Mar89, MH89, MHJ95, Say91, Urm97, WL91]. programmering [HJL00]. programmers [San97, SC97]. Programmierung [Küh97]. Programming [CWR12, Car92, Car93, FP02, FP05a, FP05b, GS02a, Int95b, ISO08b, LBF08, McC98, Morxxa, Pri01, Ree97, Sip98, Urm04, Wil06, AC87, All03, Ano08, Cel95, Cel05a, Cel05b, CH07, DD03, DD06, DuB03b, DuB05, Feu95a, Feu95b, Feu96, Feu97a, Feu97b, Feu99b, Feu99c, Feu02, GR01, Har03b, Har04, HF06, Kar10, KGZ99b, KGZ99a, Kuz10, MTC00, MS^{+}00, M^{+}06, O’N94, OO01, Ora90b, PS99, RP01, San97, SC97, Say90a, SEG^{+}03, SH90, SW01, Sun03, Sun06, Syb01, TSC02, Urm96a, Urm96b, Urm97, Urm00, Vie00, Vie06, Vie07, Wel01, Int95a]. Project [AC87]. projects [AN02]. proposed [AC87]. protecting [Che15]. Protocol [Int96]. przewodnik [WDM02]. PSM [Int99d, ISO11e, Ame96, DD98]. publications [Syb90]. publish [Tho03]. puzzles [Cel07]. Pylons [Gar09]. Python [Gar09, TSC02].

QL [ISO08b]. QMF [FS90, SG91]. Quality [Ame97]. Quantification [Met01]. QUEL [Kie84]. QUEL-like [Kie84]. Queries [BCKH75, Has96, HV00, Kie84, Ten10, Ten11, VH08, VH14]. Query [BED96, BC73, Din94, EDB89, SAS95, BED93, CB74, HH88, HH91b, Met01, Per98a, Per98b, SAS93c, SG91]. Querying [MB06, GG97]. Quick [Wei00, Dew08, KK01, RC03, RC04]. quicksteps [Ste05].

R [ABC^{+}76, Wün13]. R1997 [Ame97]. R2 [HDA11, R^{+}11, Ten11]. rapid [Gar09]. RDBMS [CK12, HN13]. readiness [She00]. Readings [Sto88]. Real [Mit03, MA03, She03, Bal07b, Mar05]. Real-world [Mit03, MA03]. Recovery [DHMT00, IBM09a, Des00, Des01, McB02, Pre07]. Refactoring [FL08]. Reference [Gen02, Gur02, KL99, Morxxb, Sti01, Wei00, WA02, Ano05, Bha98, Cog04, DuB04, FB98, FBD98, Feu99a, FPD99, Feu03, FPD03, FPD04, FPD08, Fre98, Gen00, Gen05a, GW99, GW02, Gro02, HL95, Inf95a, Inf00a, Inf00b, Ise01, J^{+}05, KK01, Kre00a, Kre00b, LT91, Lan04a, Lan04b, Pos07, Ree03, Ree07, RC03, RC04, SAS89, SAS90, SAS93b, SAS95, Sei01, Sha00, Sha01, Sha07, WGOD10, vdL89, vdL07b]. Related [ISO11b, ME00]. Relational [ABC^{+}76, BCKH75, Cod70, Cod71, Cod72, Cod87, Dar12, Gra12, Gri12, Lah05, Mel01, WC12, All03, Bre91, Dat87, Dat89, DD93, Dat05, For99, MD92, MS02, Mel03b, New92, Pas90, Pas93, Van91, vdL07a]. relationship [Gog94]. release [SAS93c, SAS94a]. Remote [Ame93, Int96, Int98, Int00]. replication [ZB04]. report [SAS94a]. reporting [GG97, Ham05, Ham06, Lar06c, Mis09, T^{+}04, T^{+}06, T^{+}08]. requirements [M^{+}07]. resource [Mic99]. Resultatsystem [HG00]. reuse [BP08]. Review [Abr03, She00]. revisited [Kie84]. RMI [Küh97]. ROM [Feu00a]. Routines [ISO08b]. Running [DW06, OH12].

Sam [McG98, SP98]. SAMeDL [Int95b, Int95a]. Sams [Mau01, MB00b, Mel03a, Bal06, But03, For13, MB00a, MSM00, Mel02, Mel04, SP00, SPJ08, WS99]. SAS [Ano04, Ano06b]. SAS/ACCESS [Ano04, Ano06b, SAS90, SAS93a, SAS93b, SAS94a, SAS94b]. scalable [Kie09, Yad07]. Scaling [Cho11]. Schaum [MTC00]. Schemas [ISO11c]. Schemata [ISO11c]. schnelles [Eis05]. Science [Bad20]. scratch [Max00, Max01]. Scripting [Mar01]. Scripts [LK98]. Search [HCxx]. Securing [Che15]. Security [Küh97, PB93, WSS^{+}20, And03, Int00]. Seeing [Abr03]. Seidman [Abr03]. Select [New92]. self [Ano06a, Hot09, MHTH07, M^{+}07, Mac07, SH92, Vee08]. self-paced [Ano06a, Hot09, MHTH07, M^{+}07, Mac07, Vee08]. self-teaching [SH92]. Sequel [Bow01, CB74, CAE^{+}76]. series [Per98a, Per98b, PP98, Per98c]. Serious [Atk02]. Server [Abr03, Ano06a, Ano07b, Bag06, Bal06, Bea04, BG08, Buc01, Byr99, Car07, CP99, Col10, CH07, Dal01, Del00, DS01, Des00, Des01, DC03, DC06, Dew07, Dew08, Eis03, ES99, Eng01, EC07, GD06, Gib08, GMT95, Gri02, GR01, Gun05, GH06, Ham05, Ham06, HDA11, H^{+}04, Hot08, Hot09, Ise01, Jor06, JWO09, Kle06, KV07, K^{+}07, Laf97, Lar06b, Lar06c, LS01, LBF08, MCT08, MKR07, M^{+}07, Mal11, Mar01, M^{+}06, McG98, McG99, McM01, Mis09, MM07, MTK06, Nan07, Nie07, N^{+}10, OC01, Ote06, Ote10, Pet00, Pet05, Pet06, Pet08, Rab01, Ran96, R^{+}07, R^{+}11, RP01, R^{+}05, RDD^{+}06, SAS93b, Sei01, Sha00, Sha01, Sha07, Sha02, She00, SSS10, SC09, SS03, Spe02a, S^{+}06, Sta10, SW01, Sun03, Sun06, TM05, Ten10]. Server [Ten11, TW06, T^{+}06, T^{+}08, Utl00, VCD16, Vee08, Vie00, Vie06, Vie07, Wat06, WF10, Wil06, WLT07, AN02, AFB11, AER96, Che15, FWS^{+}03, HJL00, HCxx, Kho92, Kuz10, LGMM09, MHTH07, Mac07, Mic97, PS99, Pur94, San97, SC97, Sol96, SR98, VB07, WS99, Wyn99, And03, Ano07a, Ber03, BG04, BB04, C^{+}05, CB00, ES96, Eng97, Ham07, HQ06, Hen02, IBM09b, Kap01, Kie09, Lah05, McB02, MD92, Mit03, MA03, Nat90a, OC99, Pur94, Rai08, SA01, She03, Sin00, Syb01, T^{+}04, WF10]. Server2005 [Whi06]. Services [S^{+}06, Ten10, Ham07, HQ06, KV07, Lar06c, Mar02, Mis09, Nan07, T^{+}04, T^{+}06, T^{+}08]. servlet [Klu98]. servlets [Har01a]. sets [Cel08]. setup [Cho11]. Seven [RW12]. sharding [Cho11]. Shared [Cod70]. sicher [Eis03]. site [SA01]. sites [Kie09, Val06]. skills [HG04, Ora90a, Pet05]. skyting [HG00]. Smart [LGMM09]. smarties [Cel95, Cel05a]. SNORT [Reh03]. Software [CM90, Hai09, SAS94b, Mag09, Mar05]. Solutions [McD03, LGMM09, Lar91, Lar93, S^{+}06]. sophisticated [AO10]. Source [WA02, Che91, LW03, Mon03, Yad07]. sous [Dup05]. Spatial [Int99a]. Special [Col00a]. Specialization [Ame93, Int98, Int96]. specification [X/O94]. Specifications [ISO11b]. Specified [Ame97]. Specifying [BCKH75]. spoken [Say89b]. SQL [Abr03, All01, AFB11, Ano06a, Ano07b, Bal06, BG08, Byr99, Car07, C^{+}05, CH07, DD98, Del00, DS01, Din94, Eis03, Eng01, Feu00a, Gib08, GR01, Gun05, GH06, HQ06, HDA11, Hau01, H^{+}04, Hot08, Hot09, Int95c, Int95b, Int99b, Int99c, Int99a, Int99d, Intxxa, Int00, ISO08c, ISO08d, ISO08a, ISO08b, ISO11a, ISO11d, ISO11e, ISO11c, ISO11b, Ise01, JH98, Jor06, JWO09, JRC97, Kle06, K^{+}07, LGMM09, Lar06b, Lar06c, LBF08, MCT08, MHTH07, M^{+}07, Mac07, Mal11, Mar01, McG99, McM01, Mis09, Nan07, O’R00, Ote06, Ote10, Pet05, Pet06, Pet08, Rab01, Ran96, R^{+}07, R^{+}11, Sei01, Sha07, Sha02, She00, SSS10, SC09, SS03, Spe02a, S^{+}06, Sta10, Sun06, Ten10, Ten11, T^{+}04, T^{+}06, T^{+}08, VCD16, Vee08, Vie00, Vie07, Wat06, AC87, Ame89, Ame92]. SQL [Ame93, Ame95, Ame96, Ame97, AN02, All03, AO10, All05, And03, And13, Ano87, Ano90, Ano95, Ano07a, App02, AER96, Aul01, Aul03, Bad20, Bag06, Bak89, Bal07a, Bal07b, Bea04, Bea05, Bei07, BG08, Ber03, BG04, Bha98, BB04, BED93, BED96, BED01, Bow01, Bra88, Bre91, Buc01, Bur01, Bur03, BCE03, Com95, CO93, Car92, Car93, CM90, Cel95, Cel05a, Cel05b, Cel06, Cel07, Cel08, Cha12, CT02, Cha90, CR90, CP99, Che91, Che15, CB00, dC90, Col00a, Col00b, Col10, Cro88, Cro89, CR07, Cum07, Dal01, Dat87, DW89, Dat89, DD93, DD97, DD98, Dav99, DT02, DeB18, Des00, Des01, Deu13, Dew07, Dew08, Dic02, DS05, EDB89, ES96, Eng97, ES99, EC07, FIP90, FIP93, FIP95, FS90, FR06, FL08, Feh03, Feh08, Feu95a]. SQL [Feu95b, Feu96, Feu97a, Feu97b, FB98, FBD98, Feu99a, Feu99b, FPD99, Feu99c, Feu00b, FO00a, FO00b, Feu01a, Feu01b, Feu02, FP02, Feu03, FPD03, FPD04, FP05a, FP05b, FPD08, For13, For99, Fre98, Gen04a, Gen04b, Gen06, GD06, GMT95, Gil91a, Gil91b, Gri02, GW90, GW99, GW02, Gro02, GG97, Gru90, Gru96, GP94, GP03, GG89, Gur02, dH05, Ham05, Ham06, Ham07, HM05, Har98a, Har98b, Har00, Har03a, Har05, HN13, Har01b, Har92, HM92, Has96, HL95, Hen00, Hen02, HV00, Hes07, Hot07, Hou01a, Hou01b, Hov95, HH88, HH89, HH91a, HH91b, HH92, Int92, Int95a, Int96, Int98, Intxxb, Inf95a, Inf95b, Inf00a, Inf00b, Inf00c, J^{+}05, Kap01, Kar10, Kho92, Kie09, Kie84, KJ02, KGZ99b, KGZ99a, KK01]. SQL [KK02, KKH04, KHK08, KV07, KM07, Kre00a, Kre00b, KT03, KT08, Küh97, Kuz10, LT91, Lad00, Laf97, Lah05, Lan04a, Lan04b, Lar91, Lar93, Lef91, Lin07, Lin08, LS01, Lis07, LC99, LD91, LAG97, Lus88a, Lus88b, Lyn90, MKR07, Mar89, MH89, MHJ95, MTC00, MSM00, Mau01, McB02, McC98, McC05, M^{+}06, McC94, McD03, MS97, McG98, MD92, MS93, ME00, Mel01, MS02, Mel03b, MB06, Met01, MB02, MB04, Mit03, MA03, Mol06, MM07, MTK06, Nat90b, Nan05, NF05, Nat90a, New92, Nie07, N^{+}10, Nys07, O’G98, Ode99, Opp05, Ora90b, Ora90a, Ora90c, Ora90d, OC99, OC01, Owe96, Owe98, PS99, Pas90, Pas93, Pat99, Pat02, Per98a, Per98b, PP98, Per98c, Pet00, PB93, PMD05, Pra91, Pra05, PF02, Pri04, Pur94, Rai08, RP01, Ris10, R^{+}05]. SQL [RDD^{+}06, RTRT94, Ros00, RR00, RS04, RC03, SAS89, SAS90, SAS93a, SAS93b, SAS93c, SAS94a, SAS94b, SAS95, San97, SC97, Say89a, Say89b, Say90a, Say90c, Say91, SG91, SA01, Sha00, Sha01, She03, SH03, Sin00, Sip98, Sno00, Sol96, SR98, Spe02b, SG02, SH90, SP98, Ste00, SP00, SPJ08, SH92, SW01, Sun03, Sun06, Syb01, TM05, Tar89, Tay01, Tay10, Ten11, Tho03, Tow03a, Tow03b, Tre99, TC89, TL01, TW06, Uni87, Urm96a, Urm96b, Urm97, Urm00, Urm04, Utl00, Van91, VB07, Vie06, VH08, VH14, Wad12, WS99, WGOD10, WF10, Wel91, Wel92, Wel01, Whi06, Wil06, WC05, Wip89, Woo95, WLT07, Wyn99, YŞ01, dHN05, dR10, vdL88, vdL89, vdL91, vdL92, vdL93, vdL98, vdL07a, vdL07b]. SQL*Forms [McC94]. SQL*Loader [GM01]. SQL*Plus [Gen99, Gen00, Gen02, Gen05a, Say90b, dH05, Sac87, dHN05, Gen05b]. SQL-3 [For99]. SQL-92 [Pas93]. SQL-Beschreibung [Eis03]. SQL-DMO [Mit03, MA03]. SQL-like [Kie84]. SQL/400 [MHJ95]. SQL/92 [DD98, DD98]. SQL/Ada [Int95b, Int95a]. SQL/Bindings [Intxxa]. SQL/CLI [Int95c, ISO08c, Ame95]. SQL/DS [DW89, FS90, Gil91a, Gil91b, HN13, Say89a, Say90c, SG91, Wip89, SAS90]. SQL/Foundation [Int99c, ISO11d]. SQL/Framework [Int99b, ISO11a]. SQL/JRT [ISO08b]. SQL/LPP [Per98a, Per98b, PP98, Per98c]. SQL/MED [ISO08d]. SQL/OLB [ISO08a, Intxxb]. SQL/PSM [Int99d, ISO11e, Ame96]. SQL/Schemata [ISO11c]. SQL/XML [ISO11b, MB06]. SQLAlchemy [Cop08, Gar09]. SQLite [AO10, Hal07, New05, Owe06, WSS^{+}20]. SQLJ [ME00, Pri01]. SQLServerCentral.com [Ano07a]. SQUARE [BCKH75]. stability [Aul03]. Standard [DD93, DD98, Deu13, FIP90, FIP93, FIP95, Ame92, CO93, Cro88, Cro89, Dat87, Dat89, DD97, Gar09, vdL89, AC87]. Staroffice [GGK99]. starter [GD06]. Statement [Int96]. statements [AC87, McC05]. statistical [TL01]. step [Hot08, Mis09, SC09]. Stonebraker [Sto11]. Storage [Hai09]. Store [Bac09]. Stored [Ame96, Int99d, ISO11e, Aul03, CH07, HF06, Hen02, Owe96, Owe98, SW01, Sun03, Sun06, Wel01]. Story [Men13, Pre12, Mar05]. storytelling [DeB18]. strongly [McC05]. structure [Dav99]. Structured [BC73, Din94, BED93, CB74, HH88, HH91b, BED96, EDB89]. Studio [MM07, VB07]. study [Car07, DT02, DC03, DC06, DHP04, Jor06]. style [Cel05b]. Sublanguage [BCKH75, Cod71]. Sublanguages [Cod72]. supplement [Ano04, Ano06b]. Support [HJL00]. survival [SS03]. Sybase [DHMT00, GMT95, MD92, SAS93b, AER96, Eps13, Pur94, Ran96]. Syntax [Gen02, Inf00a, Inf00b]. System [Gri12, Nie09, Bre91, Che91, HJL00, Kop04, MD92, Pur94, Rab01, ABC^{+}76, Bru09, Kru09, McG09, Wad12]. System/370 [Wad12]. Systems [Ame89, BH09, Bru09, Dye99, Gra12, Int96, Int98, Pet19, RG03, Sto88, WC12, Ame92, Dat75, Dat77, Dat81, Dat86, Dat90, Dat95, Dat00, Dat04, GMUW02, Mic97, Pur94, Rio05, RTRT94, WL91].

T [BG08, RP01, Sun06, Wel01]. T-SQL [BG08, RP01, Sun06, Wel01]. tables [Cel08]. take [AO10]. tasks [BG04]. Teach [LAG97, MS97, Bal06, But03, For13, MB00a, MSM00, Mau01, MB00b, McG98, Mel02, Mel03a, Mel04, SP98, SP00, SPJ08, WS99]. teaching [SH92]. team [Mar05]. technical [SAS94a, Sei01, Syb90]. Techniques [Reh03, Gil91a, Gil91b, GR93, KJ02, MKR07, McB02, TSC02, Tre99]. technologies [ME00]. Technology [Ame93, Ame95, Ame96, Int92, Int95a, Int95c, Int95b, Int96, Int98, Int99b, Int99c, Int99a, Int99d, Intxxb, Intxxa, Int00, ISO08c, ISO08d, ISO08a, ISO08b, ISO11a, ISO11d, ISO11e, ISO11c, ISO11b]. temporal [Cel08, Per98a, Per98b]. text [Tho03]. theory [Dat05]. thinking [Cel08]. those [McC94]. Threads [Küh97]. time [Per98a, Per98b, PP98, Per98c, Sno00]. time-oriented [Sno00]. Tipps [DHMT00]. tips [And13, KJ02, Tre99]. Title [Int92]. tjener [HJL00]. TM [ISO08b]. together [ME00]. tolerant [Yad07]. tool [AO10]. Toolkit [MTK06]. Tools [Klu98, Gar09, McB02]. Toolset [MTK06]. Total [Nie09]. trainer [K^{+}10]. training [Ano06a, Hot09, Laf97, MHTH07, M^{+}07, Mac07, Mic97, Rab01, Vee08]. Transact [Hen00, KGZ99b, KGZ99a, Lan04a, Lan04b, MSM00, Mau01, MS97, Spe02b, SG02, TW06]. Transact-SQL [Hen00, KGZ99b, KGZ99a, Lan04a, Lan04b, MSM00, Mau01, MS97, Spe02b, SG02, TW06]. Transaction [GR93]. Trees [Abr03]. triggers [CH07, HL95, Inf95a, Owe96, Owe98, Wel01]. Troubleshooting [Hou01b, Ste05]. Tuning [CG93, GC96, Gur02, And13, Bur01, Eng97, Eng01, GP03, Har01b, Tow03a, Tow03b, Urm97]. TURBO [Tar89]. Turing [Ash87]. tutorial [Inf95b, Inf00c, LT91, vdL07b]. twenty [Ash87]. tworzenia [Feh03]. typed [McC05]. Types [ISO08b, CH07].

u.v.m [Klu98]. UDB [C^{+}04, C^{+}05]. Ultra [Kie09]. Ultra-fast [Kie09]. ultra-scalable [Kie09]. Understanding [Gru90, Lyn90, MS93, ME00, Mel01, Pac06, Pas93, MS02, Mel03b]. Unified [CAE^{+}76]. UNIX [Fre98, IBM09b, Tar89, Gab07, IBM09a, Sip13]. unleashed [PS99, McN97, Ran96, R^{+}07, R^{+}11, Sol96, SR98]. Upgrading [Tra04, Ste05]. usage [SAS89, SAS90, SAS93b, SAS95]. Use [Morxxc, Say90b, SG91]. user [Dat87, Dat89, DD93, DD97, HH89, MD92, Sac87, Stu01]. Using [Aul03, BED96, BC73, dC90, Col00b, GW90, ISO08b, Inf95a, Lef91, McG99, PH98, RYW02, Reh03, SAS93c, Woo95, Wyn99, BED93, BED01, Col00a, DuB03b, DuB05, EDB89, Gar09, HL95, Kie09, LW03, Lin07, Lin08, MHTH07, Mac07, McC94, Pur94, TL01, Yad07, Yan04]. utilisation [Cha96].

variants [BED01]. ved [HJL00]. verification [Per98a, Per98b]. Verifying [Ame97]. version [McC94, SAS89, SAS90, SAS93b, SAS94b, SAS95, Sac87]. views [O’G98]. virtual [Cel08]. visual [CT02, Mon03, TC89, Val06, McC98, MM07, VB07]. visualization [HDA11].

Warehouse [MTK06, O’G98, Rai08]. Watcom [Woo95]. Web [Cog04, DHMT00, EFH^{+}11, Kie09, Mar02, Gar09, HJL00, Kop04, Ano08, App02, AER96, Bac07, DuB01, DuB02, EFH^{+}11, FWS^{+}03, Gab07, G^{+}06, GS02a, Gla04, LW03, McC98, Nar05, Ode99, SEG^{+}03, TSC02, Tho03, Utl00, Val06, WT01, WL01, WL02, WL04]. web-system [Kop04]. Weblogg [Kop04]. Weblogg-Forum [Kop04]. website [Yan04]. weekend [Dic02, Gre02, Tay01]. weeks [RW12]. Welt [EFH^{+}11]. White [Mar05]. wichtigen [GGK99]. Wide [AER96]. window [SAS93c, SAS95]. Windows [DC03, DC06, Fre98, Wei04, Blu07, IBM09a, IBM09b]. Wissen [Eis03]. within [AO10]. WMI [Mar01]. Woods [Abr03]. Work [Pet19]. Workbook [FO00a, Feu00b, FO00b, Ros00, RR00]. Working [FS90]. world [Bal07b, EFH^{+}11, Mit03, MA03, SF13, She03, AER96]. Wrox [GD06]. WSG [Gar09].

X3.135 [Ame89, Ame92]. X3.135-1992 [Ame89, Ame92]. X3.168 [AC87, Ame89]. X3.168-1989 [Ame89]. X3.168-198X. [AC87]. XHTML [Ano08]. XML [ISO11b, App02, Gri02, Hen02, Kle06, Mar02, Mar01, MB06, Sun03]. XML-Related [ISO11b]. XPath [MB06]. XQuery [MB06]. XSLT [Tho03]. XSQL [Tho03].

Years [Gra12, WC12, Ash87]. yourself [Bal06, But03, For13, LAG97, MB00a, MSM00, Mau01, MB00b, MS97, McG98, Mel02, Mel03a, Mel04, SP98, SP00, SPJ08, WS99].

z [And13]. z/OS [And13]. zarzadzania [Feh03]. zur [DHMT00].

References

Astrahan:1976:SRR

[ABC^{+}76] M. M. Astrahan, M. W. Blasgen, D. D. Chamberlin, K. P. Eswaran, J. N. Gray, P. P. Griffiths, W. F. King, R. A. Lorie, P. R. McJones, J. W. Mehl, G. R. Putzolu, I. L. Traiger, B. W. Wade, and V. Watson. System R: a relational approach to database management. ACM Transactions on Database Systems, 1(2):97–137, June 1976. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

                                                 URL http://www.acm.org/pubs/
                                                 articles/journals/tods/1976-
                                                 1-2/p97-astrahan/p97-astrahan.
                                                 pdf;   http://www.acm.org/pubs/
                                                 citations/journals/tods/1976-
                                                 1-2/p97-astrahan/. Also pub-

lished in/as: IBM, San Jose, Research Report. No. RJ-1738, Feb. 1976. Reprinted in [Sto88].

Abrahams:2003:BRS

[Abr03] Alan S. Abrahams. Book review: Seeing the woods for the decision trees: A review of Data Mining with SQL Server 2000 by Claude Seidman. IEEE Distributed Systems Online, 4(1), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

dsonline.computer.org/0301/d/
bks_c.htm.

ANSI-X3H2:1987:DPA

[AC87] American National Standards Institute. Technical Committee X3H2. Data Base Committee X3 Information Processing Systems and Computer and Business Equipment Manufacturers Association. Draft proposed American National Standard embedding of SQL statements into programming languages: X3.168-198X. X3 Project 571-D. Computer and Business Equipment Manufacturers Association, Washington, DC, USA, 1987. 31 pp.

Ashley:1996:SSS

[AER96] Ed Ashley, Beth Epperson, and Kerri De Rosier. Sybase SQL server on the World Wide Web. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1996. ISBN 1-85032-815-3. xviii + 366 pp. LCCN QA76.9.C55 A84 1996.

Anderson:2011:MSS

[AFB11] Mark Anderson, James Fox, and Christian Bolton. Microsoft SQL server 2008 administration for Oracle DBAs. McGraw-Hill, New York, NY, USA, 2011. ISBN 0-07-170064-1 (paperback). xviii + 571 pp. LCCN QA76.9.D3 A6434 2011.

Allen:2001:OPS

[All01] Christopher Allen. Oracle PL/SQL 101. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-219229-1, 0-07-212606-X. xx + 420 pp. LCCN ????

Allen:2003:IRD

[All03] Christopher Allen. Introduction to relational databases and SQL programming. Osborne/McGraw-Hill, Berkeley, CA, USA, 2003. ISBN 0-07-222924-1. 560 pp. LCCN ????

Allison:2005:SMA

[All05] Cecelia Allison. SQL for Microsoft Access. Wordware Pub., Plano, TX, 2005. ISBN 1-4237-2624-3 (electronic book). ???? pp. LCCN ????

ANSI:1989:AXI

[Ame89] American National Standards Institute. ANSI X3.135-1992: Information Systems — Database Language — SQL (includes ANSI X3.168-1989). American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1989. ISBN ???? ???? pp. LCCN ????

                                        US$220. URL http://www.ansi.
                                        org/.

ANSI:1992:ANSd

[Ame92] American National Standards Institute. American national standard for information systems: database language — SQL: ANSI X3.135-1992. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1992. xiv + 580 pp. Revision and consolidation of ANSI X3.135-1989 and ANSI X3.168-1989, Approved October 3, 1989.

ANSI:1993:AII

[Ame93] American National Standards Institute. ANSI/ISO/IEC 9579-2-1993: Information Technology — Remote Database Access — Part 2: SQL Specialization. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1993. ISBN ???? ???? pp. LCCN ???? US$72. URL http://www.

ansi.org/.

ANSI:1995:AIIa

[Ame95] American National Standards Institute. ANSI/ISO/IEC 9075-3-1995: Information Technology — Database Languages — SQL — Part 3: Call-Level Interface (SQL/ CLI). American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1995. ISBN ???? ???? pp. LCCN ????

US$110. URL http://www.ansi.
org/.

ANSI:1996:AII

[Ame96] American National Standards Institute. ANSI/ISO/IEC 9075-4-1996: Information Technology — Database languages — SQL — Part 4: Persistent Stored Modules (SQL/PSM). American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1996. ISBN ???? ???? pp. LCCN

???? URL http://www.ansi.org/

.

ANSI:1997:AER

[Ame97] American National Standards Institute. ANSI/EIA 555-1988 (R1997): Lot Acceptance Procedure for Verifying Compliance with the Specified Quality Level (SQL) in PPM. EIA (CEMA), ????, 1997. ISBN ???? ???? pp. LCCN ????

US$32. URL http://www.ansi.
org/.

Albrecht:2002:MAP

[AN02] Ralf Albrecht and Natasha Nicol. Microsoft Access projects with Microsoft SQL server. Microsoft Press, Bellevue, WA, USA, 2002. ISBN 0-7356-1002-9. xiv + 637 pp. LCCN QA76.9.C55 A515 2002. US$.

Andrews:2003:SSS

[And03] Chip Andrews. SQL Server security. McGraw-Hill/Osborne, Emeryville, CA, USA, 2003. ISBN 0-07-222515-7. xxiv + 322 pp. LCCN ????

Andrews:2013:DST

[And13] Tony Andrews. DB2 SQL tuning tips for z/OS developers. IBM Press, Upper Saddle River, NJ, USA, 2013. ISBN 0-13-303846-7 (paperback). xv + 138 pp. LCCN QA76.9.D3 A6435 2013.

Anonymous:1987:DLS

[Ano87] Anonymous. Database language SQL 2600. Technical report, U.S. Department of Commerce, Gaithersburg, MD, USA, March 10, 1987. 5 pp. Federal information processing standards publication, FIPS PUB, vol. 127. Shipping list number 87-330-P.

Anonymous:1990:DLS

[Ano90] Anonymous. Database language SQL 2600. Technical report, National Computer Systems Laboratory, National Institute of Standards and Technology, Gaithersburg, MD, USA, February 2, 1990. 8 pp. Federal information processing standards publication, vol.

FIPS pub 127-1. Supersedes FIPS PUB 127 - 1987 March 10. Shipping list number 90-353-P.

Anonymous:1995:SE

[Ano95] Anonymous. SQL environments.

Technical report, Computer Systems Laboratory, National Institute of Standards and Technology, U.S. Department of Commerce, Technology Administration, Gaithersburg, MD, USA, February 3, 1995. vi + 58 pp. Shipping list number 95-0126-P.

Anonymous:2004:SAS

[Ano04] Anonymous. SAS/ACCESS 9.1.2 supplement for MySQL: SAS. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 2004. ISBN 1-59047-458-9 (paperback). iv + 23 pp. LCCN QA76.9.D3 S25463 2004.

Anonymous:2005:MLR

[Ano05] Anonymous. MySQL: language reference. MySQL Press, Indianapolis, IN, USA, 2005. ISBN 0-672-32633-7. xiii + 481 pp. LCCN QA76.73.S67 M97 2005.

Anonymous:2006:MSP

[Ano06a] Anonymous. MCTS self-paced training kit (exam 70-431): Microsoft SQL Server 2005–implementation and maintenance. Microsoft Press, Bellevue, WA, USA, 2006. ISBN 0-7356-2271-X. xxxviii + 950 pp. LCCN QA76.3 .M33486 2006.

Anonymous:2006:SAS

[Ano06b] Anonymous. SAS/ACCESS 9.1.3 supplement for MYSQL: SAS. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 2006. ISBN 1-59047-924-6. 23 pp. LCCN MLCM 2006/42439 (T).

Anonymous:2007:BSC

[Ano07a] Anonymous. The best of SQLServer-Central.com : high performance SQL Server. Red Gate Books, Cambridge, UK, 2007. ISBN 1-906434-00-X (vol. 5 : paperback), 1-906434-07-7 (vol. 6 : paperback), 1-906434-36-0 (vol. 7 : paperback).

Anonymous:2007:PSS

[Ano07b] Anonymous. Professional SQL Server 2005 administration. Wrox professional guides. Wiley/Wrox, Indianapolis, IN, USA, 2007. ISBN 0-470-05520-0. xxx + 735 pp. LCCN QA76.9.D3. URL http://

www.loc.gov/catdir/enhancements/
fy0741/2006032940-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0741/2006032940-d.html;   http:
//www.loc.gov/catdir/toc/ecip072/
2006032940.html.

Anonymous:2008:JEW

[Ano08] Anonymous. Just enough Web programming with XHTML, PHP, and MYSQL. Cengage Learning, Course Technology PTR, Boston, MA, USA, 2008. ISBN 1-59863-481-X. ???? pp. LCCN ????

Allen:2010:DGS

[AO10] Grant Allen and Michael Owens. The definitive guide to SQLite: [take control of this compact but powerful tool to embed sophisticated SQL databases within your applications!]. The expert’s voice in open source. Apress, Berkeley, CA, USA, second edition, 2010. ISBN 1-4302-3225-0 (paperback), 1-4302-3226-9 (e-book). xix + 347 pp. LCCN QA76.73.S67 O94 2010.

URL http://benaiahbooks.com/
assets/upload/booklet/small.
pdf;   http://link.springer.
com/openurl?genre=book\%26isbn=
978-1-4302-3225-4.

Appelquist:2002:XSD

[App02] Daniel K. Appelquist. XML and SQL: developing Web applications. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-65796-1. xxiii + 224 pp. LCCN QA76.76.H94 A67 2002.

Ashenhurst:1987:ATA

[Ash87] Robert L. Ashenhurst, editor. ACM Turing Award Lectures: the first twenty years, 1966–1985. ACM Press anthology series. ACM Press and Addison-Wesley, New York, NY 10036, USA and Reading, MA, USA, 1987. ISBN 0-201-07794-9. xviii + 483 pp. LCCN QA76.24 .A33 1987.

Atkinson:2002:CMS

[Atk02] Leon Atkinson. Core MySQL:

The Serious Developer’s Guide. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-066190-2. xxiv + 721 pp. LCCN QA76.3.S67 A855 2002. US$44.99. URL http:

//www.phptr.com/ptrbooks/ptr_
0130661902.html.

Ault:2001:ODS

[Aul01] Michael R. Ault. Oracle8i DBA:

SQL and PL/SQL. Coriolis Group Books, Scottsdale, AZ, USA, 2001.

ISBN 1-57610-037-7 (paperback). xxxii + 428 pp. LCCN QA76.3 .A96 2001eb.

Ault:2003:UOS

[Aul03] Mike Ault. Using Oracle SQL stored outlines and optimizer plan stability. Rampant TechPress, Kittrell, NC, USA, 2003. ISBN 0-9740716-8-4. 36 pp. LCCN ????

Bacon:2007:PPM

[Bac07] Jono Bacon. Practical PHP and MySQL: building eight dynamic Web applications. Negus live Linux series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2007. ISBN 0-13-223997-3 (paperback). xiii + 512 pp. LCCN TK5105.888 .B325 2007. URL http://www.

                                        loc.gov/catdir/toc/ecip0619/
                                        2006027701.html.

Bachman:2009:OID

[Bac09] Charles W. Bachman. The origin of the integrated data store (IDS): The first direct-access DBMS. IEEE Annals of the History of Computing, 31(4):42–54, October/ December 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Badia:2020:SDS

[Bad20] Antonio Badia. SQL for Data Science. Data-Centric Systems and Applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2020. ISBN 3-030-57591-8, 3-030-57592-6 (e-book), 3-030-57593-4. ISSN 2197-9723 (print), 2197-974X (electronic). xi + 285 + 16 pp. URL

http://www.springerlink.com/
content/978-3-030-57592-2.

Bagui:2006:LSS

[Bag06] Sikha Bagui. Learning SQL on SQL Server 2005. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN ???? xvi + 324 pp. LCCN ????

Baker:1989:GSD

[Bak89] Richard H. Baker. Guide to SQL:

database management for IBM PCs and compatibles. Scott, Foresman and Company, Glenview, IL, USA, 1989. ISBN 0-673-38356-3. xxviii + 417 pp. LCCN QA76.73.S67 B35 1989. US$22.95.

Balter:2006:STY

[Bal06] Alison Balter. Sams teach yourself Microsoft SQL Server 2005 Express in 24 hours. Sams teach yourself. Howard W. Sams, Indianapolis, IN 46268, USA, 2006. ISBN 0-672-32741-4. xv + 436 pp. LCCN QA76.9.D3 B3262

2006. URL http://www.loc.gov/
catdir/toc/fy0702/2005937217.
html.

Bales:2007:BPN

[Bal07a] Donald Bales. Beginning PL/SQL:

from novice to professional. The expert’s voice in Oracle. Apress, Berkeley, CA, USA, 2007. ISBN 1-59059-882-2. xxvi + 468 pp. LCCN QA76.73.S67 B357 2007.

Bales:2007:BPS

[Bal07b] Donald J. Bales. Beginning PL/ SQL: from novice to professional; [learn to model the real world through PL/SQL]. The expert’s voice in Oracle. Apress, Berkeley, CA, USA, 2007. ISBN 1-59059-882-2. xxvi + 468 pp. LCCN QA76.73.S67 B357 2007.

Banker:2012:MA

[Ban12] Kyle Banker. MongoDB in action. Manning, Greenwich, CT, USA, 2012. ISBN 1-935182-87-0. xxi + 287 pp. LCCN QA76.9.D3 B354 2012.

Bardzell:2004:MDM

[Bar04] Jeffrey Bardzell. Macromedia Dreamweaver MX 2004: with ASP, ColdFusion and PHP. Macromedia in association with Peachpit Press, Berkeley, CA, USA, 2004. ISBN 0-321-24157-6. viii + 502 pp. LCCN TK5105.8885.D74 B364 2004.

Bordoloi:2004:SSS

[BB04] Bijoy Bordoloi and Douglas Brian Bock. SQL for SQL Server. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, 2004. ISBN 0-13-113299-7 (paperback). xix + 391 pp. LCCN QA76.73.S58 2004.

Boyce:1973:USE

[BC73] R. Boyce and D. Chamberlin. Using a structured English query language as a data definition facility. IBM Research Report RJ1318, IBM Corporation, San Jose, CA, USA, December 1973. This is one of the earliest reports that describes the relational database language that later became standardized as SQL; see [Cha12] for that history, and [CB74] for the conference paper that made early SQL well known.

Burleson:2003:OSI

[BCE03] Donald K. Burleson, Joe Celko, and Dave Ensor. Oracle SQL internals handbook. Rampant TechPress, Kittrell, NC, USA, 2003. ISBN 0-9744355-1-1. 195 pp. LCCN ????

Boyce:1975:SQR

[BCKH75] Raymond F. Boyce, Donald D.

Chamberlin, W. Frank King III, and Michael M. Hammer. Specifying queries as relational expressions: The SQUARE data sublanguage. Communications of the ACM, 18(11):621–628, November 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://dblp.uni-
trier.de/db/journals/cacm/BoyceCKH75.

html. Also published in/as: in Douque74 (IFIP TC-2), 1974. Also published in/as: ACM SIGPLAN-SIGIR 1974, Nance (ed), pp. 31–

47. Also published in/as: IBM, San Jose Research Report RJ1291, October 1973.

Beauchemin:2004:FLS

[Bea04] Bob Beauchemin. A first look at SQL Server 2005 for developers. Microsoft .NET development series. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-321-18059-3 (paperback). xxxviii + 693 pp. LCCN QA76.9.C55 B43 2004.

Beaulieu:2005:LS

[Bea05] Alan Beaulieu. Learning SQL.

O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00727-2 (paperback). xiii + 289 pp. LCCN QA76.73.S67

                                        B43 2005. URL http://www.
                                        loc.gov/catdir/enhancements/
                                        fy0715/2006274865-d.html.

Bowman:1993:PSH

[BED93] Judith S. Bowman, Sandra L. Emerson, and Marcy Darnovsky. The practical SQL handbook: using structured query language. Addison-Wesley, Reading, MA, USA, second edition, 1993. ISBN 0-201-62623-3. xvii + 453 pp. LCCN QA76.73.S67 E54 1993. US$29.95; CAN$38.95.

Bowman:1996:PSH

[BED96] Judith S. Bowman, Sandra L. Emerson, and Marcy Darnovsky. The Practical SQL Handbook: Using Structured Query Language. Addison-Wesley, Reading, MA, USA, third edition, 1996. ISBN 0-201-44787-8. xxvi + 454 pp. LCCN QA76.73.S67 B69 1996. US$39.95. Includes CD-ROM.

Bowman:2001:PSH

[BED01] Judith S. Bowman, Sandra L. Emerson, and Marcy Darnovsky. The practical SQL handbook: using SQL variants. Addison-Wesley, Reading, MA, USA, fourth edition, 2001. ISBN 0-201-70309-2. xxxvi + 469 pp. LCCN QA76.73.S67 B688 2001.

Beighley:2007:HFS

[Bei07] Lynn Beighley. Head first SQL. Head first series. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2007. ISBN 0-596-52684-9. xxxv + 571 pp. LCCN QA76.73.S67 B435

Bell:2007:EM

[Bel07] Charles A. (Charles Andrew) Bell.

Expert MySQL. Expert’s voice in open source. Apress, Berkeley, CA, USA, 2007. ISBN 1-59059-741-9 (paperback), (paperback). xix + 577 pp. LCCN QA76.73.S67 B44 2007.

Bersinic:2003:PDS

[Ber03] Damir Bersinic. Portable DBA:

SQL Server. Osborne/McGraw-Hill, Berkeley, CA, USA, 2003. ISBN 0-07-223016-9. 400 pp. LCCN QA76.9.D3 B477 2003.

Bernier:2004:BP

[Ber04] Robert Bernier. The book of PostgreSQL. No Starch Press, San Francisco, CA, USA, 2004. ISBN 1-59327-045-3. ???? pp. LCCN QA76.9.D3 B479 2004. URL

ftp://uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;   http:
//www.loc.gov/catdir/toc/ecip0417/
2004008803.html.

Bersinic:2004:PDS

[BG04] Damir Bersinic and Stephen Giles.

Portable DBA: SQL Server ; [easyto-find data for on-the-job tasks]. McGraw-Hill/Osborne, Emeryville, CA, USA, 2004. ISBN 0-07-223016-9 (paperback). xii + 339 s pp. LCCN ???? US$19.99.

Ben-Gan:2008:MSS

[BG08] Itzik Ben-Gan. Microsoft SQL Server 2008 T-SQL fundamentals. Microsoft Press, Bellevue, WA, USA, 2008. ISBN 0-7356-2601-4. ???? pp. LCCN ????

Bulger:2004:MPD

[BGWG04] Brad Bulger, Jay Greenspan, David Wall, and Jay Greenspan. MySQL/PHP database applications. John Wiley and Sons, Inc., New York, NY, USA, second edition, 2004. ISBN 0-7645-4963-4. xliii + 764 pp. LCCN QA76.73.S67 G73 2004. URL

ftp://uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;   http:
//www.loc.gov/catdir/bios/wiley045/
2002114859.html;   http://www.
loc.gov/catdir/description/wiley038/
2002114859.html;   http://www.
loc.gov/catdir/toc/wiley041/
2002114859.html.

Bergin:2009:CDM

[BH09] Thomas J. Bergin and Thomas Haigh. The commercialization of database management systems, 1969–1983. IEEE Annals of the History of Computing, 31(4):26–41, October/December 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bhamidipati:1998:SPR

[Bha98] Kishore Bhamidipati. SQL programmer’s reference. Osborne/ McGraw-Hill, Berkeley, CA, USA, 1998. ISBN 0-07-882460-5. xxix + 337 pp. LCCN QA76.73.S67 B45 1998. US$16.99.

Blum:2001:P

[Blu01] Richard Blum. Postfix. Howard W. Sams, Indianapolis, IN 46268, USA, 2001. ISBN 0-672-32114-9. xix + 593 pp. LCCN TK5105.73 .B578 2001. EUR$56.65.

Blum:2007:PW

[Blu07] Richard Blum. PostgreSQL 8 for Windows. Database professional’s library. McGraw-Hill, New York, NY, USA, 2007. ISBN 0-07-148562-

7. xv + 384 pp. LCCN QA76.9.D3 B583 2007. URL http://www.

loc.gov/catdir/enhancements/
fy0710/2007007201-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0710/2007007201-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0710/2007007201-t.html.

Boenigk:2003:PGP

[Boe03] Cornelia Boenigk. PostgreSQL:

Grundlagen — Praxis — Anwendungsentwicklung mit PHP. dpunkt-Verlag, Heidelberg, Germany, 2003. ISBN 3-89864-175-9. xiv + 412 pp. LCCN ???? EUR

39.00. URL http://www.gbv.de/
du/services/toc/bs/352938609.

Bohl:2013:PMD

[Boh13] Marilyn Bohl. Product managing DB2’s early development. IEEE Annals of the History of Computing, 35(2):67–68, April/June 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bowman:2001:PSS

[Bow01] Judith S. Bowman. Practical SQL:

the Sequel. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-61638-6 (paperback). xv + 329 pp. LCCN QA76.73.S67 B695 2001. Includes CD-ROM.

Boronczyk:2008:PMC

[BP08] Tim Boronczyk and Martin E.

Psinas. PHP and MySQL: create-modify-reuse. John Wiley and Sons, Inc., New York, NY, USA, 2008. ISBN 0-470-19242-9. ???? pp. LCCN QA76.73.P224

                                        B64 2008. URL http://www.
                                        loc.gov/catdir/enhancements/
                                        fy0827/2008011996-d.html;   http:
                                        //www.loc.gov/catdir/enhancements/
                                        fy0827/2008011996-t.html.

Brindle:2001:ME

[BR01] Joshua Brindle and Chris Reighard. MySQL essentials. Linux. Prima Publishing, Roseville, CA, USA, 2001. ISBN 0-7615-2445-2. ???? pp.

Bradley:1988:BFP

[Bra88] James Bradley. Business file processing!: COBOL, CICS, and SQL. Mitchell Publishing, Inc., Santa Cruz, CA, USA, 1988. ISBN 0-394-39217-5. 400 pp. LCCN QA76.9.F5 B725 1988. US$28.00. Includes two chapters on hash file processing.

Breaden:1991:PPR

[Bre91] Matthew Alexis Breaden. P-SQL: a portable relational database system. Thesis (m.s.), University of California, Santa Cruz, Santa Cruz, CA, USA, 1991. ii + 35 pp.

Brueck:2009:SMS

[Bru09] Robert L. Brueck. System 2000: The MRI Systems Corporation. IEEE Annals of the History of Computing, 31(4):76–86, October/ December 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Buczek:2001:ISS

[Buc01] Greg Buczek. Instant SQL Server 2000 applications. Osborne/Mc-Graw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-213320-1. xxv + 932 pp. LCCN QA76.9.D3 B823 2001. Includes CD-ROM.

Burleson:2001:OHP

[Bur01] Donald K. Burleson. Oracle highperformance SQL tuning. Osborne/ McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-219058-2. xxvi + 575 pp. LCCN QA76.9.D3 B87525 2001.

Burleson:2003:ASD

[Bur03] Donald K. Burleson. Advanced SQL Database Programmer Handbook. Rampant TechPress, Kittrell, NC, USA, 2003. ISBN 0-9744355-2-X. x + 112 pp. LCCN ????

Butcher:2003:STY

[But03] Tony Butcher. Sams teach yourself MySQL in 21 days. Howard W. Sams, Indianapolis, IN 46268, USA, second edition, 2003. ISBN 0-672-32392-3. xxiii + 608 pp. LCCN QA76.73.S67 B88 2003.

Byrne:1999:MAM

[Byr99] Jeffrey Byrne. MCSE: Administering Microsoft SQL Server 7. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-010795-6. xxi + 567 pp. LCCN QA76.3 .B88 1998. URL http:/

/www.phptr.com/bookseri/mcse.
html.

Chen:2004:MDU

[C^{+}04] Whei-Jen Chen et al., editors. MySQL to DB2 UDB conversion guide. IBM Corporation, San Jose, CA, USA, 2004. ISBN 0-7384-9825-4. xx + 374 pp. LCCN QA76.73.S67 M976 2004eb.

Chen:2005:MSS

[C^{+}05] Whei-Jen Chen et al., editors. Microsoft SQL Server to IBM DB2 UDB conversion guide. Redbooks. IBM Corp., International Technical Support Organization, San Jose, CA, USA, 2005. ISBN 0-7384-9119-5. xxviii + 524 pp. LCCN QA76.76.S64 M53 2005eb.

Chamberlin:1976:SUA

[CAE^{+}76] Donald D. Chamberlin, Morton M. Astrahan, Kapali P. Eswaran, Patricia Priest Griffiths, Raymond A. Lorie, James W. Mehl, Phyllis Reisner, and Bradford Warren Wade. SEQUEL 2: a unified approach to data definition, manipulation, and control. IBM Journal of Research and Development, 20(6): 560–575 (or 560–565??), November 1976. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Carter:1992:PSO

[Car92] John (John R.) Carter. Programming in SQL with Oracle, Ingres, and dBase. Computer studies series. Blackwell Scientific Publications, Oxford, UK, 1992. ISBN 0-632-03136-0. xi + 308 pp. LCCN QA76.73.S67C37 1992.

Carter:1993:PSO

[Car93] John (John R.) Carter. Programming in SQL with Oracle, Ingres, and dBase IV. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1993. ISBN 0-13-014325-1. xi + 308 pp. LCCN QA76.73.S67 C37 1993.

Carpenter:2007:MSS

[Car07] Tom Carpenter. MCTS SQL Server 2005: implementation and maintenance, study guide (exam 70-431). McGraw-Hill, New York, NY, USA, 2007. ISBN 0-07-226321-0 (set), 0-07-226322-9 (paperback), 0-07-226323-7 (CD-ROM). xxxi + 606 pp. LCCN HD8039.D37

C37 2007. URL http://www.
loc.gov/catdir/enhancements/
fy0814/2007280940-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0814/2007280940-d.html;   http:
//www.loc.gov/catdir/toc/fy0802/
2007280940.html.

Chamberlin:1974:SSE

[CB74] Donald D. Chamberlin and Raymond F. Boyce. SEQUEL: a structured English query language. In Randall Rustin, editor, FIDET 74’: Data models: data-structure-set versus relational: Workshop on Data Description, Access, and Control, May 1–3, 1974, Ann Arbor, Michigan, pages 249–264. ACM Press, New York, NY 10036, USA,

1974. URL http://www.almaden.
ibm.com/cs/people/chamberlin/
sequel-1974.pdf.

Chipman:2000:MAD

[CB00] Mary Chipman and Andy Baron.

Microsoft Access developer’s guide to SQL Server. Howard W. Sams, Indianapolis, IN 46268, USA, 2000. ISBN 0-672-31944-6. xiii + 840 pp. LCCN QA76.9.D3 C45 2000.

Chodorow:2010:MDG

[CD10] Kristina Chodorow and Michael Dirolf. MongoDB: the definitive guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2010. ISBN 1-4493-8156-1 (paperback), 1-4493-8157-X. xvii + 193 pp. LCCN HF5548.38.O64 .C462 2010.

Celko:1995:JCS

[Cel95] Joe Celko. Joe Celko’s SQL for smarties: advanced SQL programming. Morgan Kaufmann series in data management systems. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1995. ISBN 1-55860-323-9. xxiv + 467 pp. LCCN QA76.73.S67 C44 1995.

Celko:2005:JCSa

[Cel05a] Joe Celko. Joe Celko’s SQL for smarties: advanced SQL programming. The Morgan Kaufmann series in data management systems. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, third edition, 2005. ISBN 0-12-369379-9 (paperback). xxviii + 808 pp. LCCN QA76.73.S67

                                        C44 2005. URL http://www.
                                        loc.gov/catdir/enhancements/
                                        fy0626/2005279919-d.html;   http:
                                        //www.loc.gov/catdir/enhancements/
                                        fy0626/2005279919-t.html.

Celko:2005:JCSb

[Cel05b] Joe Celko. Joe Celko’s SQL programming style. Elsevier, Amsterdam, The Netherlands, 2005. ISBN 1-4237-2245-0 (electronic book). xviii + 216 pp. LCCN ????

Celko:2006:JCA

[Cel06] Joe Celko. Joe Celko’s analytics and OLAP in SQL. The Morgan Kaufmann series in data management systems. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2006. ISBN 0-12-369512-0 (paperback). xx + 181 pp. LCCN QA76.73.S67 C432 2006. URL http://www.

loc.gov/catdir/enhancements/
fy0704/2006287375-d.html;   http:
//www.loc.gov/catdir/toc/fy0705/
2006287375.html.

Celko:2007:JCS

[Cel07] Joe Celko. Joe Celko’s SQL puzzles and answers. Morgan Kaufmann series in data management systems. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2007. ISBN 0-12-373596-3 (paperback), 0-08-049168-5 (electronic book). xvii + 325 pp. LCCN ????

Celko:2008:JCT

[Cel08] Joe Celko. Joe Celko’s thinking in sets: auxiliary, temporal, and virtual tables in SQL. The Morgan Kaufmann series in data management systems. Elsevier Morgan Kaufmann, Amsterdam, The Netherlands, 2008. ISBN 0-12-374137-8. xix + 362 pp. LCCN QA76.73.S67 C463 2008.

Corrigan:1993:OPT

[CG93] Peter Corrigan and Mark Gurry. ORACLE Performance Tuning. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, September 1993. ISBN 1-56592-048-1. xxxv + 603 pp. LCCN QA76.9.D3 C674 1993. US$34.95. URL http://www.

oreilly.com/catalog/9781565920484.

Comingore:2007:PSS

[CH07] Derek Comingore and Douglas Hinson. Professional SQL Server 2005 CLR programming: with stored procedures, functions, triggers, aggregates, and types. Programmer to programmer; Wrox professional guides. Wrox/Wiley Technology Pub., Indianapolis, IN, USA, 2007. ISBN 0-470-05403-4 (paper/website). xxii + 409 pp. LCCN QA76.73.S67 C655 2007. URL http://www.

loc.gov/catdir/enhancements/
fy0741/2006031028-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0741/2006031028-d.html;   http:
//www.loc.gov/catdir/toc/ecip071/
2006031028.html.

Charness:1990:CCL

[Cha90] Dave Charness. CLING/SQL: Common LISP to INGRES/SQL interface. Memorandum UCB/ERL M90/40, Electronics Research Laboratory, College of Engineering, University of California, Berkeley, CA, USA, May 8, 1990. 28 pp.

Chapuis:1996:OUA

[Cha96] Roger Chapuis. Oracle 7: utilisation and administration. Guide de formation. Editions Laser, Saint Dénis, France, 1996. ISBN 2-910525-24-4. 347 pp. LCCN ????

Chamberlin:2012:EHS

[Cha12] Donald D. Chamberlin. Early history of SQL. IEEE Annals of the History of Computing, 34(4):78–82, October/December 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Chen:1991:ACS

[Che91] Haobo Chen. Automated conversion system for development of SQL database management system from a FORTRAN-based source. Thesis (m.s.), Department of Computer Science, University of California, Davis, Davis, CA, USA, 1991. 143 pp.

Cherry:2015:SSS

[Che15] Denny Cherry. Securing SQL server: protecting your database from attackers. Elsevier/Syngress, Waltham, MA, third edition, 2015. ISBN 0-12-801275-7. xxii + 440 pp. LCCN QA76.73.S67 C478 2015. URL http://www.

loc.gov/catdir/enhancements/
fy1606/2014953492-d.html.

Chodorow:2011:SMS

[Cho11] Kristina Chodorow. Scaling MongoDB: [sharding, cluster setup, and administration]. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2011. ISBN 1-4493-0321-8 (paperback), 1-4493-0412-5 (e-book), 1-4493-0387-0 (e-book). ix + 49 pp. LCCN QA76.9.D3 C463 2011.

Campbell-Kelly:2012:RIN

[CK12] Martin Campbell-Kelly. The RDBMS industry: A Northern California perspective. IEEE Annals of the History of Computing, 34(4): 18–29, October/December 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Casimir-Mrowczynski:1990:SDC

[CM90] Piers Casimir-Mrowczynski. Software design with COBOL and SQL. Computer studies series. Publishers’ Business Services, Brookline Village, MA, USA, 1990. ISBN 0-632-02604-9. xii + 158 pp. LCCN QA76.73.C25 C3724 1990.

Cannan:1993:SSH

[CO93] Stephen J. (Stephen John) Cannan and G. A. M. Otten. SQL

— the standard handbook: based on the new SQL standard (ISO 9075:1922(E)). McGraw-Hill, New York, NY, USA, 1993. ISBN 0-07-707664-8. 584 pp. LCCN QA76.73.S67 C35 1993.

Codd:1970:RMD

[Cod70] E. F. Codd. A relational model of data for large shared data banks. Communications of the ACM, 13 (6):377–387, June 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://dblp.uni-trier.de/db/
                                        journals/cacm/Codd70.html. Reprinted

in [Sto88, pp. 5–15].

Codd:1971:DBS

[Cod71] Edgar F. Codd. A data base sublanguage founded on the relational calculus. In Edgar F. Codd and Albert L. Dean, editors, Data description, access and control: proceedings of 1971 ACM-SIGFIDET workshop, San Diego, California, November 11–12, 1971, pages 35–

68. ACM Press, New York, NY 10036, USA, 1971. LCCN ????

Codd:1972:RCD

[Cod72] E. F. Codd. Relational completeness of data base sublanguages. IBM Research Report RJ987, IBM Corporation, San Jose, CA, USA, 1972.

Codd:1987:RDP

[Cod87] E. F. Codd. Relational database:

a practical foundation for productivity. In Ashenhurst [Ash87], page ?? ISBN 0-201-07794-9. LCCN QA76.24 .A33 1987. ACM Turing Award lecture.

Cogswell:2004:AMP

[Cog04] Jeffrey M. Cogswell. Apache, MySQL, and PHP Web development: all-in-one desk reference for dummies. –For dummies. John Wiley and Sons, Inc., New York, NY, USA, 2004. ISBN 0-7645-4969-3. xxvi + 734 pp. LCCN QA76.73.S67 C63 2004. URL

ftp://uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;   http:
//www.loc.gov/catdir/bios/wiley046/
2003115007.html;   http://www.
loc.gov/catdir/description/wiley041/
2003115007.html;   http://www.
loc.gov/catdir/toc/wiley041/
2003115007.html.

Colburn:2000:SEU

[Col00a] Rafe Colburn. Special edition using SQL. Que Corporation, Indianapolis, IN, USA, 2000. ISBN 0-7897-1974-6. xii + 654 pp. LCCN QA76.73.S67 C65 2000.

Colburn:2000:USH

[Col00b] Rafe Colburn. Using SQL. Que Corporation, Indianapolis, IN, USA, 2000. ISBN 0-7897-1974-6. xii + 654 pp. LCCN QA76.73.S67 C65 2000.

Colledge:2010:SSA

[Col10] Rod Colledge. SQL Server 2008 administration in action. Manning, Greenwich, CT, USA, 2010. ISBN 1-933988-72-X (paperback). xxiv + 440 pp. LCCN ????

CSL:1995:SE

[Com95] Computer Systems Laboratory (U.S.). SQL environments. National Technical Information Service, Springfield, VA, USA, February 3, 1995. vi + 58 pp. Shipping list number 95-0126-P.

Copeland:2008:ES

[Cop08] Rick Copeland. Essential SQLAlchemy. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51614-2. xi + 215 pp. LCCN QA76.9.W43 C67 2008.

Chaturvedi:1999:ASS

[CP99] Divya Chaturvedi and Paritosh Pathak. Administering SQL Server 7. McGraw-Hill, New York, NY, USA, 1999. ISBN 0-07-134168-4. xxv + 515 pp.

Converse:2004:PMB

[CP04] Tim Converse and Joyce Park.

PHP5 and MySQL bible. John Wiley and Sons, Inc., New York, NY, USA, 2004. ISBN 0-7645-5746-7. xxxvii + 1042 pp. URL

ftp://uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;   http:
//www.loc.gov/catdir/bios/wiley047/
2004103176.html;   http://www.
loc.gov/catdir/description/wiley041/
2004103176.html;   http://www.
loc.gov/catdir/toc/wiley041/
2004103176.html.

Charness:1990:CSC

[CR90] Dave Charness and Lawrence A.

Rowe. CLING/SQL: Common LISP to INGRES/SQL interface. Technical report, Electronics Research Laboratory, College of Engineering, University of California, Berkeley, Berkeley, CA, USA, May 8, 1990. 28 pp.

Cumming:2007:SH

[CR07] Andrew Cumming and Gordon Russell. SQL hacks. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2007. ISBN 0-596-52799-3. xxi + 386 pp. LCCN QA76.73.S67 C85 2007eb; QA76.73.S67. URL http://www.

oreilly.com/catalog/9780596527990.

Cronin:1988:MOF

[Cro88] Daniel J. Cronin. Mastering Oracle:

featuring Oracle’s SQL standard. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1988. ISBN 0-672-48419-6. xxi + 542 pp. LCCN QA76.9.D3 C76 1989. US$24.95.

Cronin:1989:MOF

[Cro89] Daniel J. Cronin. Mastering OR-ACLE: featuring ORACLE’s SQL standard. Howard W. Sams, Indianapolis, IN 46268, USA, 1989. ISBN 0-672-48419-6. xxi + 543 pp. LCCN QA76.9.D3 C76 1989. US$24.95.

Chappell:2002:VIS

[CT02] David (David Wayne) Chappell and J. Harvey Trimble. A visual introduction to SQL. John Wiley and Sons, Inc., New York, NY, USA, 2002. ISBN 0-471-41276-7. xii + 291 pp. LCCN QA76.73.S67 C477 2002.

Cumming:2007:SHH

[Cum07] Andrew Cumming. SQL hacks. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2007. ISBN ???? xxi + 386 pp. LCCN ????

Capriolo:2012:PH

[CWR12] Edward Capriolo, Dean Wampler, and Jason Rutherglen. Programming Hive. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2012. ISBN 1-4493-1933-5 (paperback). xvii + 328 pp. LCCN ????

Dalton:2001:SSB

[Dal01] Patrick Dalton. SQL Server 2000 black book. Coriolis Group Books, Scottsdale, AZ, USA, 2001. ISBN 1-57610-770-1. xxxvii + 969 pp. LCCN ????

Darwen:2012:RMB

[Dar12] Hugh Darwen. The relational model: Beginning of an era. IEEE Annals of the History of Computing, 34(4):30–37, October/ December 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Date:1975:IDS

[Dat75] C. J. Date. An introduction to database systems. The systems programming series. Addison-Wesley, Reading, MA, USA, 1975. ISBN 0-201-14452-2. xvii + 366 pp. LCCN QA76.9.D3 D37.

Date:1977:IDS

[Dat77] C. J. Date. An introduction to database systems. The Systems programming series. Addison-Wesley, Reading, MA, USA, second edition, 1977. ISBN 0-201-14456-5. xxiii + 536 pp. LCCN QA76.9.D3 D37 1977.

Date:1981:IDS

[Dat81] C. J. Date. An introduction to database systems. Addison-Wesley systems programming series. Addison-Wesley, Reading, MA, USA, third edition, 1981. ISBN 0-201-14471-9, 0-201-14439-5 (paperback). xxviii + 574 pp. LCCN QA76.9.D3 D37 1981. Two volumes.

Date:1986:IDS

[Dat86] C. J. Date. An introduction to database systems. Addison-Wesley systems programming series. Addison-Wesley, Reading, MA, USA, fourth edition, 1986. ISBN 0-201-14201-5 (vol. 1). ???? pp. LCCN QA76.9.D3 D37 1986. US$31.95.

Date:1987:GSS

[Dat87] C. J. Date. A guide to the SQL standard: a user’s guide to the standard relational language SQL. Addison-Wesley, Reading, MA, USA, 1987. ISBN 0-201-05777-8 (paperback). xiv + 205 pp. LCCN QA76.9.D3 D36951 1987. US$23.95.

Date:1989:GSS

[Dat89] C. J. Date. A guide to the SQL standard: a user’s guide to the standard relational language SQL. Addison-Wesley, Reading, MA, USA, second edition, 1989. ISBN 0-201-50209-7. xii + 228 pp. LCCN QA76.9.D3 D24 1989.

Date:1990:IDS

[Dat90] C. J. Date. An introduction to database systems. Addison-Wesley systems programming series. Addison-Wesley, Reading, MA, USA, fifth edition, 1990. ISBN 0-201-51381-1 (vol. 1). ???? pp. LCCN QA76.9.D3 D37 1990.

Date:1995:IDS

[Dat95] C. J. Date. An introduction to database systems. Addison-Wesley systems programming series. Addison-Wesley, Reading, MA, USA, sixth edition, 1995. ISBN 0-201-54329-X. xxiii + 839 pp. LCCN QA76.9.D3 D3659 1995.

Date:2000:IDS

[Dat00] C. J. Date. An introduction to database systems. Addison-Wesley, Reading, MA, USA, seventh edition, 2000. ISBN 0-201-38590-2. xxii + 938 pp. LCCN QA76.9.D3 D3659 2000.

Date:2004:IDS

[Dat04] C. J. Date. An introduction to database systems. Pearson/Addison Wesley, Boston, MA, USA, eighth edition, 2004. ISBN 0-321-19784-4. xxvii + 983 + 22 pp. LCCN QA76.9.D3 D3659 2003.

Date:2005:DDR

[Dat05] C. J. Date. Database in depth:

relational theory for practitioners. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-10012-4. xviii + 208 pp. LCCN QA76.9.D3 D384 2005.

David:1999:AAS

[Dav99] Michael M. David. Advanced ANSI SQL data modeling and structure processing. The Artech House computer science library. Artech House Inc., Norwood, MA, USA, 1999. ISBN 1-58053-038-9. xx + 244 pp. LCCN QA76.73.S67 D39 1999.

Darie:2005:BPM

[DB05] Cristian Darie and Mihai Bucica. Beginning PHP 5 and MySQL e-commerce: from novice to professional. Apress, Berkeley, CA, USA, 2005. ISBN 1-59059-392-8 (paperback). xxiii + 537 pp. LCCN HF5548.32 .D353 2005. URL http://www.

loc.gov/catdir/enhancements/
fy0663/2005280433-d.html;   http:
//www.loc.gov/catdir/toc/fy0601/
2005280433.html.

Darie:2008:BPM

[DB08] Cristian Darie and Emilian Balanescu. Beginning PHP and MySQL e-commerce: from novice to professional. Apress, Berkeley, CA, USA, second edition, 2008. ISBN 1-59059-864-4 (paperback). xxvii + 707 pp. LCCN HF5548.32 .D353 2008.

Darie:2006:BPP

[DBB06] Cristian Darie, Mihai Bucica, and Emilian Balanescu. Beginning PHP and PostgreSQL e-commerce: from novice to professional. Apress, Berkeley, CA, USA, 2006. ISBN 1-59059-648-X (paperback). xxii + 616 pp. LCCN QA76.73.P224 D37 2006.

Chou:1990:US

[dC90] George Tsu der Chou. Using SQL. Que Corporation, Indianapolis, IN, USA, 1990. ISBN 0-88022-507-6. xx + 474 pp. LCCN QA76.73.S67 C48 1990. US$24.95 (CDN$31.95, UK£22.95 U.K.).

Desai:2003:MWS

[DC03] Anil Desai and James Chellis. MCSE Windows Server 2003 active directory planning, implementation, and maintenance study guide. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2003. ISBN 0-7821-4263-X. xli + 451 pp. LCCN QA76.3 .D4685

                                        2003. URL http://www.loc.gov/
                                        catdir/enhancements/fy0615/2003106720-
                                        b.html;   http://www.loc.gov/
                                        catdir/enhancements/fy0615/2003106720-
                                        d.html;   http://www.loc.gov/

catdir/enhancements/fy0615/2003106720-
t.html.

Desai:2006:MWS

[DC06] Anil Desai and James Chellis.

MCSE Windows Server 2003 active directory planning, implementation, and maintenance study guide. John Wiley and Sons, Inc., New York, NY, USA, second edition, 2006. ISBN 0-7821-4451-9. 492 pp. LCCN QA76.3 .D4685 2006. URL http://www.

loc.gov/catdir/enhancements/
fy0650/2006280392-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0826/2006280392-t.html.

Date:1993:GSS

[DD93] C. J. Date and Hugh Darwen. A guide to the SQL Standard: a user’s guide to the standard relational language SQL. Addison-Wesley, Reading, MA, USA, third edition, 1993. ISBN 0-201-55822-X. xvii + 414 pp. LCCN QA76.9.D3 D3695 1993.

Date:1997:GSS

[DD97] C. J. Date and Hugh Darwen. A guide to the SQL standard: a user’s guide to the standard database language SQL. Addison-Wesley, Reading, MA, USA, fourth edition, 1997. ISBN 0-201-96426-0. xxii + 522 pp. LCCN QA76.73.S67 D38 1997.

Date:1998:SSS

[DD98] Chris J. Date and Hugh Darwen.

SQL — der Standard: SQL/92 mit den Erweiterungen CLI und PSM. (German) [SQL — The Standard: SQL/92 with the CLI and PSM Extensions]. Addison-Wesley-Longman, Bonn, Germany, 1998.

ISBN 3-8273-1345-7. 582 pp. LCCN ????

Douglas:2003:PCG

[DD03] Korry Douglas and Susan (Susan P.) Douglas. PostgreSQL: a comprehensive guide to building, programming, and administering PostgreSQL databases. Developer’s library. Howard W. Sams, Indianapolis, IN 46268, USA, 2003. ISBN 0-7357-1257-3 (paperback). xvii + 790 pp. LCCN QA76.9.D3 D685 2003.

Douglas:2006:PCG

[DD06] Korry Douglas and Susan (Susan P.) Douglas. PostgreSQL: the comprehensive guide to building, programming, and administering PostgreSQL databases. Developer’s library. Howard W. Sams, Indianapolis, IN 46268, USA, second edition, 2006. ISBN 0-672-32756-2 (paperback). xxi + 1006 pp. LCCN QA76.9.D3 D685 2006.

DeBarros:2018:PSB

[DeB18] Anthony DeBarros. Practical SQL: a beginner’s guide to storytelling with data. No Starch Press, San Francisco, CA, USA, 2018. ISBN 1-59327-827-6 (paperback). xxxiii + 351 pp. LCCN QA76.73.S67 D44 2018.

Delaney:2000:IMS

[Del00] Kalen Delaney. Inside Microsoft SQL Server 2000. Microsoft programming series. Microsoft Press, Bellevue, WA, USA, 2000. ISBN 0-7356-0998-5. xxxv + 1046 pp. LCCN QA76.9.C55 D43 2001.

Desai:2000:SSB

[Des00] Anil Desai. SQL Server 7: backup and recovery. Database professional’s library. Osborne/McGraw-Hill, Berkeley, CA, USA, 2000. ISBN 0-07-212410-5. xxix + 480 pp. LCCN QA76.9.C55 D465 2000.

Desai:2001:SSB

[Des01] Anil Desai. SQL Server 2000:

backup and recovery. Database professional’s library. Osborne/Mc-Graw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-213027-X. xxi + 698 pp. LCCN QA76.9.C55 D466 2001.

Deutsch:2013:SSH

[Deu13] Donald R. Deutsch. The SQL standard: How it happened. IEEE Annals of the History of Computing, 35(2):72–75, April/June 2013. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Dewson:2007:BSS

[Dew07] Robin Dewson. Beginning SQL Server 2005 Express for Developers: From Novice to Professional. Springer eBook Collection Professional and Applied Computing [Dig. Serial]; Springer-12059 [Dig. Serial]. Robin Dewson, Berkeley, CA, USA, 2007. ISBN 1-4302-0329-3. LCCN QA76.73.S67.

Dewson:2008:BSS

[Dew08] Robin Dewson. Beginning SQL Server 2008 for developers: from novice to professional; [the quick and efficient path to proficiency in SQL Server 2008]. The expert’s voice in SQL Server. Apress, Berkeley, CA, USA, 2008. ISBN 1-59059-958-6. xxiv + 471 pp. LCCN ????

Davies:2006:MC

[DF06] Alex Davies and Harrison Fisk. MySQL clustering. MySQL Press, Indianapolis, IN, USA, 2006. ISBN 0-672-32855-0 (paperback). xiii + 202 pp. LCCN QA76.73.S67 D395 2006.

Haan:2005:MOS

[dH05] Lex de Haan. Mastering Oracle SQL and SQL*plus. The expert’s voice in Oracle. Apress, Berkeley, CA, USA, 2005. ISBN 1-59059-448-7. xx + 468 pp. LCCN QA76.73.S7 H33 2005.

deHaan:2007:AMD

[dHK07] Lex de Haan and Toon Koppelaars. Applied mathematics for database professionals. Apress, Berkeley, CA, USA, 2007. ISBN 1-59059-745-1. xxviii + 376 pp. LCCN QA76.9.D3 H239

2007. URL http://www.loc.gov/
catdir/toc/fy0804/2008299427.

html. This book provides the mathematical background of logic and set theory to relational databases.

Dicken:2000:DLO

[DHMT00] Hans Dicken, Gunther Hipper, and Peter Müßig-Trapp. Datenbanken unter Linux: Oracle 8i, MySQL, Adabas, Informix, Sybase, DB2, PostgreSQL, MiniSQL, Empress; [Tipps zur optimalen Installation und Konfiguration; Backup, Recovery, Ausfallsicherheit; mit PHP und Java ins Web]. MITP-Verlag, Bonn, Germany, 2000.

ISBN 3-8266-0555-1. 516 (est.) pp. LCCN ????

deHaan:2005:MOS

[dHN05] Lex de Haan and Mogens Nørgaard.

Mastering Oracle SQL and SQL*plus. The expert’s voice in oracle. Apress, Berkeley, CA, USA, 2005. ISBN 1-59059-448-7. xx + 468 pp. LCCN QA76.73.S7

H33 2005. URL http://www.
loc.gov/catdir/enhancements/
fy0663/2005299972-d.html.

DuBois:2004:MCS

[DHP04] Paul DuBois, Stefan Hinz, and Carsten Pedersen. MySQL certification study guide. MySQL Press, Indianapolis, IN, USA, 2004. ISBN 0-672-32632-9 (paperback). xv + 629 pp. LCCN QA76.3 .D688 2004.

Dicken:2002:LSW

[Dic02] Deanna Dicken. Learn SQL in a weekend. In a weekend. Premier Press, Inc., Indianapolis, IN, USA, 2002. ISBN 1-931841-62-4. xxi + 497 pp. LCCN QA76.73.S67 D53 2002.

Din:1994:SQL

[Din94] Akeel I. Din. Structured Query Language (SQL): a practical introduction. NCC Blackwell, Cambridge, MA, USA, 1994. ISBN 1-85554-357-5. 190 pp. LCCN QA76.73.S67 D56 1994.

Davis:2007:LPM

[DP07] Michele E. Davis and Jon A.

(John Andrew) Phillips. Learning PHP and MySQL. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2007. ISBN 0-596-51401-8 (paperback). xiii + 411 pp. LCCN QA76.73.P224 D38

2007. URL http://www.loc.gov/
catdir/toc/fy0803/2008271806.
html.

deRooij:2010:DS

[dR10] Ton de Rooij. Databases en SQL. Academic Service, Postbus 81, 2870 AB Schoonhoven, The Netherlands, fourth edition, 2010. ISBN 90-395-2617-6. viii + 156 pp. LCCN ????

Delaney:2001:IMS

[DS01] Kalen Delaney and Ron Soukup. Inside Microsoft SQL Server 2000. Microsoft programming series. Microsoft Press, Bellevue, WA, USA, 2001. ISBN 0-7356-0998-5. xxxv + 1046 pp. LCCN QA76.9.C55 D43 2001.

Donahoo:2005:SPG

[DS05] Michael J. Donahoo and Gregory D. (Gregory David) Speegle. SQL: practical guide for developers. The Morgan Kaufmann practical guides series. Elsevier Morgan Kaufmann, Amsterdam, The Netherlands, 2005. ISBN 0-12-220531-6. xiv + 249 pp. LCCN QA76.73.S67 D576 2005. URL http://www.

loc.gov/catdir/enhancements/
fy0625/2005049161-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0625/2005049161-t.html.

Dawes:2002:OOI

[DT02] Chip Dawes and Biju Thomas. OCA/OCP: introduction to Oracle9i SQL: study guide. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2002.

ISBN 0-7821-4062-9. 539 pp. LCCN ????

DuBois:1999:M

[DuB99] Paul DuBois. MySQL. New Riders Publishing, Carmel, IN, USA, 1999. ISBN 0-7357-0921-1. xxix + 756 pp. LCCN QA76.73.S67 D58 1999.

DuBois:2001:MPW

[DuB01] Paul DuBois. MySQL and Perl for the Web. New Riders Publishing, Carmel, IN, USA, 2001. ISBN 0-7357-1054-6. xvi + 529 pp. LCCN ????

DuBois:2002:MPW

[DuB02] Paul DuBois. MySQL and Perl for the Web. New Riders Publishing, Carmel, IN, USA, 2002. ISBN 0-7357-1054-6. xvi + 529 pp. LCCN QA76.73.S67 D585 2002.

DuBois:2003:MC

[DuB03a] Paul DuBois. MySQL Cookbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00145-2. xxvii + 992 pp. LCCN QA76.73.S67 D587 2003.

DuBois:2003:MDG

[DuB03b] Paul DuBois. MySQL: the definitive guide to using, programming, and administering MySQL4. Developer’s library. Howard W. Sams, Indianapolis, IN 46268, USA, second edition, 2003. ISBN 0-7357-1212-3. xxiv + 1219 pp. LCCN QA76.73.S67 D588 2003.

DuBois:2004:MLR

[DuB04] Paul DuBois. MYSQL language reference. Howard W. Sams, Indianapolis, IN 46268, USA, 2004. ISBN 0-672-32633-7. ???? pp.

DuBois:2005:MDG

[DuB05] Paul DuBois. MySQL: the definitive guide to using, programming, and administering MySQL 4.1 and 5.0. Developer’s library. Howard W. Sams, Indianapolis, IN 46268, USA, third edition, 2005. ISBN 0-672-32673-6 (paperback). xxii + 1294 pp. LCCN QA76.73.S67 D588 2005.

DuBois:2006:MC

[DuB06] Paul DuBois. MySQL cookbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-52708-X. LCCN ????

DuBois:2007:MC

[DuB07] Paul DuBois. MySQL cookbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2007. ISBN 0-596-52708-X (paperback). xxvi + 948 pp. LCCN QA76.73.S67 D82 2007.

DuBois:2013:M

[DuB13] Paul DuBois. MySQL. Addison-Wesley, Reading, MA, USA, fifth edition, 2013. ISBN 0-13-303855-6. xx + 1153 pp. LCCN QA76.73.S67.

Dupuy:2005:CBD

[Dup05] Stéphane Dupuy. Conception d’une base de données PostgreSQL sous Linux. Mémoire de diplôme d’ingénieur, Centre d’enseignement, Université de soutenance, Bordeaux, France, 2005. 99 pp.

Date:1989:GSD

[DW89] C. J. Date and Colin J. White. A guide to SQL/DS. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-14688-6. xix + 505 pp. LCCN QA76.9.D3 D36954 1989.

Dalheimer:2006:RL

[DW06] Matthias Kalle Dalheimer and Matt Welsh. Running Linux. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fifth edition, 2006. ISBN 0-596-00760-4 (paperback). xviii + 951 pp. LCCN QA76.76.O63 R855 2006. URL http://www.

loc.gov/catdir/enhancements/
fy0715/2007270110-d.html.

Dye:1999:ODS

[Dye99] Charles Dye. Oracle Distributed Systems. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-432-0. xv + 530 pp. LCCN QA76.73.S674 D94 1999. US$39.95.

URL http://www.oreilly.com/
catalog/9781565924321;   http:
//www.oreilly.com/catalog/ordistsys/

.

Dyer:2005:MN

[Dye05] Russel J. T. Dyer. MySQL in a nutshell. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00789-2. xxiv + 321 pp. LCCN QA76.73.S67 D93 2005.

Dyer:2008:MN

[Dye08] Russell J. T. Dyer. MySQL in a nutshell. In a nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2008. ISBN 0-596-51433-6. xvi + 545 pp. LCCN QA76.73.S67 D93 2008.

Ericsson:2007:SSD

[EC07] Robert Ericsson and Jason Cline. SQL Server 2005 for developers. Charles River Media programming series. Charles River Media, Inc., 403 VFW Drive, PO Box 417, Rockland, MA 02370, USA, 2007. ISBN 1-58450-388-2 (paperback with CD). x + 405 pp. LCCN QA76.9.C55

E75 2007. URL http://www.
loc.gov/catdir/toc/ecip0613/
2006016156.html.

Emerson:1989:PSH

[EDB89] Sandra L. Emerson, Marcy Darnovsky, and Judith S. Bowman. The practical SQL handbook: using Structured Query Language. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-51738-8. xiii + 393 pp. LCCN QA76.73.S67 E54 1989. US$21.95.

Edlich:2011:NEW

[EFH^{+}11] Stefan Edlich, Achim Friedland, Jens Hampe, Benjamin Brauer, and Markus Brückner. NoSQL: Einstieg in die Welt nichtrelationaler Web 2.0 Datenbanken. (German) [NoSQL: Entry to the world of nonrelational Web 2.0 databases]. Carl Hanser, München, second edition, 2011. ISBN 3-446-42753-8 (print), 3-446-42855-0 (e-book). xi + 392 pp. LCCN ????

Eisentraut:2003:POH

[Eis03] Peter Eisentraut, editor. PostgreSQL: das offizielle Handbuch; [das Insider-Wissen der PostgreSQL-Entwickler; mit ausführlicher SQL-Beschreibung; den PostgreSQL-Server sicher beherrschen]. Datenbanken. mitp-Verlag, Bonn, Germany, 2003. ISBN 3-8266-1337-6. 814 (est.) pp. LCCN ???? EUR 34.95.

Eisentraut:2005:PGP

[Eis05] Peter Eisentraut. PostgreSQL gepackt: [schnelles und effektives Nachschlagen; Datentypen, Funktionen und Operatoren; Performance, Programme und Client-APIs]. Die ge-packte Referenz. mitp-Verlag, Bonn, Germany, 2005. ISBN 3-8266-1493-3. 479 pp. LCCN ???? EUR 14.95.

England:1997:SSP

[Eng97] Ken England. The SQL Server 6.5 performance optimization and tuning handbook. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1997. ISBN 1-55558-180-3 (paperback). xiii + 259 pp. LCCN QA76.9.C55 E64 1997.

England:2001:MSS

[Eng01] Ken England. Microsoft SQL Server 2000: performance optimization and tuning handbook. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 2001. ISBN 1-55558-241-9 (paperback). xiv + 390 pp. LCCN QA76.9.C55 E635 2001.

Epstein:2013:HS

[Eps13] Bob Epstein. History of Sybase.

IEEE Annals of the History of Computing, 35(2):31–41, April/ June 2013. CODEN IAHCEX.

ISSN 1058-6180 (print), 1934-1547 (electronic).

England:1996:SSH

[ES96] Ken England and Nigel Stanley. The SQL Server handbook: a guide to Microsoft database computing. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1996. ISBN 1-55558-152-8. xiv + 441 pp. LCCN QA76.73.S67 E55 1996.

England:1999:SSH

[ES99] Ken England and Nigel Stanley. The SQL Server 7.0 handbook: a guide to Microsoft database computing. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1999. ISBN 1-55558-201-X (paperback). xi + 451 pp. LCCN QA76.9.D3 E64 1999.

Feuerstein:1998:OPB

[FB98] Steven Feuerstein and John Beresniewicz. Oracle PL/SQL built-ins pocket reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1998. ISBN ???? vi + 72 pp. LCCN ????

Feuerstein:1998:OPS

[FBD98] Steven Feuerstein, John Beresniewicz, and Chip Dawes. Oracle PL/SQL built-ins pocket reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1998. ISBN 1-56592-456-8. vi + 72 pp. LCCN QA76.73.P258 F487 1998.

Feuerstein:1998:OBP

[FDB98] Steven Feuerstein, Charles Dye, and John Beresniewicz. Oracle Built-in Packages. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, April 1998. ISBN 1-56592-375-8. xxi + 931 pp. LCCN QA76.73.S67 F48 1998. US$46.95.

URL http://www.oreilly.com/
catalog/oraclebip. Covers Ora-

cle8.

Fehily:2003:SBK

[Feh03] Chris Fehily. SQL: błyskawiczny kurs tworzenia i zarza̧dzania aplikacjami bazodanowymi. Szybki Start. Helion, Gliwice, Poland, 2003. ISBN 83-7361-037-5. 416 pp. LCCN ????

Fehily:2008:S

[Feh08] Chris Fehily.  SQL.  Visual

quickstart guide. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, third edition, 2008. ISBN 0-321-55357-8 (paperback). xx + 483 pp. LCCN QA76.73.S67 F445

2008. URL http://www.loc.gov/
catdir/toc/fy0805/2008274673.
html.

Feuerstein:1995:OPP

[Feu95a] Steven Feuerstein. Oracle PL/SQL programming. A Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1995. ISBN 1-56592-142-9. xxvii + 885 pp. LCCN ???? “Database management systems”– Cover.. Includes 3 1/2 in. disk covering PL/SQL v1.1-2.3..

Feuerstein:1995:OPS

[Feu95b] Steven Feuerstein. Oracle PL/ SQL programming. Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1995. ISBN 1-56592-142-9. xxvii + 885 pp. LCCN QA76.73.S67 F48 1995.

Feuerstein:1996:AOP

[Feu96] Steven Feuerstein. Advanced Oracle PL/SQL: programming with packages. Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1996. ISBN 1-56592-238-7. xxvi + 661 pp. LCCN QA76.73.P258F48

                                        1996. URL http://www.ora.com/
                                        catalog/advoracle/;   http://
                                        www.oreilly.com/catalog/advoracle.

Feuerstein:1997:OPP

[Feu97a] Steven Feuerstein. Oracle PL/SQL programming. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997. ISBN ???? xxxiii + 987 pp. LCCN ????

Feuerstein:1997:OPS

[Feu97b] Steven Feuerstein. Oracle PL/SQL programming. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1997. xxxiii + 987 pp. LCCN QA76.73.S67 F49 1997.

Feuerstein:1999:OPL

[Feu99a] Steven Feuerstein. Oracle PL/SQL language: pocket reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN ???? v + 94 pp. LCCN ????

Feuerstein:1999:OPP

[Feu99b] Steven Feuerstein. Oracle PL/SQL programming: guide to Oracle8i features. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN ???? xvi + 248 pp. LCCN ????

Feuerstein:1999:OPSb

[Feu99c] Steven Feuerstein. Oracle PL/SQL programming: guide to Oracle8i features. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-675-7. xvi + 248 pp. LCCN QA76.73.P258 F489 1999.

Feuerstein:2000:OPSb

[Feu00a] Steven Feuerstein. The Oracle PL/ SQL CD bookshelf: 7 bestselling books on CD-ROM. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, version 1.0 edition, 2000. ISBN 1-56592-675-7 (paperback). LCCN Disk 1608 Protect.

Feuerstein:2000:OPD

[Feu00b] Steven Feuerstein. Oracle PL/ SQL developer’s workbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2000. ISBN 1-56592-674-9. xix + 569 pp. LCCN ????

Feuerstein:2001:OPB

[Feu01a] Steven Feuerstein. Oracle PL/SQL best practices. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2001. ISBN ???? xix + 182 pp. LCCN ????

Feuerstein:2001:OPS

[Feu01b] Steven Feuerstein. Oracle PL/SQL best practices. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2001. ISBN 0-596-00121-5. xix + 182 pp. LCCN QA76.9.D3 F49 2001.

Feuerstein:2002:OPP

[Feu02] Steven Feuerstein. Oracle PL/SQL programming. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN ???? xxiii + 989 pp. LCCN ????

Feuerstein:2003:OPL

[Feu03] Steven Feuerstein. Oracle PL/SQL language: pocket reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN ???? vi + 121 pp. LCCN ????

FIPS:1990:SDL

[FIP90] FIPS. Standard for database language SQL. Technical Report FIPS PUB 127, Federal Information Processing Standards, National Institute of Standards and Technology, Washington, DC, USA, February 2, 1990.

FIPS:1993:SDL

[FIP93] FIPS. Standard for database language SQL. Technical Report FIPS PUB 127-2, Federal Information Processing Standards, National Institute of Standards and Technology, Washington, DC, USA, June 2,

                                        1993. URL http://www.itl.nist.
                                        gov/fipspubs/fip127-2.htm.

FIPS:1995:SSE

[FIP95] FIPS. Standard for SQL environments. Technical Report FIPS PUB 193, Federal Information Processing Standards, National Institute of Standards and Technology, Washington, DC, USA, February 3,

1995. URL http://www.itl.nist.
gov/fipspubs/fip193.htm.

Faroult:2008:RSA

[FL08] Stéphane Faroult and Pascal L’Hermite. Refactoring SQL applications. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51497-2. xv + 277 pp. LCCN QA76.73.S67 F37 2008eb; QA76.73.S67 F37 2008; QA76.73.S67; Internet.

Feuerstein:2000:OPS

[FO00a] Steven Feuerstein and Andrew Odewahn. Oracle PL/SQL Developer’s Workbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, May 2000. ISBN 1-56592-674-9. xix + 569 pp. LCCN QA76.73.P258 F4875 2000. US$36.95. URL http://www.

oreilly.com/catalog/ordevworkbook.

Feuerstein:2000:OPSa

[FO00b] Steven Feuerstein and Andrew Odewahn. Oracle PL/SQL developer’s workbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2000. ISBN 1-56592-674-9. xix + 569 pp. LCCN QA76.73.P258 F4875 2000.

Fortier:1999:SIO

[For99] Paul J. Fortier. SQL-3: implementing the object-relational database. McGraw-Hill enterprise computing series. McGraw-Hill, New York, NY, USA, 1999. ISBN 0-07-022062-X (paperback). xiii + 414 pp. LCCN QA76.73.S67 F67 1999.

Forta:2013:STY

[For13] Ben Forta. Sams teach yourself SQL in 10 minutes. Howard W. Sams, Indianapolis, IN 46268, USA, fourth edition, 2013. ISBN 0-672-33607-3 (paperback). 274 pp. LCCN QA76.73.S67 F67 2013.

Feuerstein:2002:OPS

[FP02] Steven Feuerstein and Bill Pribyl. Oracle PL/SQL Programming. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2002. ISBN 0-596-00381-1. xxiii + 989 pp. LCCN QA76.73.S67 F49 2002.

                                        US$54.95. URL http://safari.
                                        oreilly.com/0596003811;   http:
                                        //www.oreilly.com/catalog/oraclep3.

Feuerstein:2005:OP

[FP05a] Steven Feuerstein and Bill Pribyl. Oracle PL/SQL Programming. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fourth edition, 2005. ISBN 0-596-00977-1 (paperback). xxiv + 1171 pp. LCCN QA76.73.S67

                                        F49 2005. URL http://www.
                                        oreilly.com/catalog/9780596009779;
                                         http://www.oreilly.com/catalog/
                                        oraclep4/.

Feuerstein:2005:OPS

[FP05b] Steven Feuerstein and Bill Pribyl. Oracle PL/SQL Programming. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fourth edition, 2005. ISBN 0-596-00977-1 (paperback).

1104 (est.) pp. LCCN ????

URL http://www.oreilly.com/
catalog/oraclep4/.

Feuerstein:1999:OPSa

[FPD99] Steven Feuerstein, Bill Pribyl, and Chip Dawes. Oracle PL/SQL language: pocket reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-457-6. v + 94 pp. LCCN QA76.73.P258 F488 1999.

Feuerstein:2003:OPS

[FPD03] Steven Feuerstein, Bill Pribyl, and Chip Dawes. Oracle PL/SQL language: pocket reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2003. ISBN 0-596-00472-9. vi + 121 pp. LCCN QA76.9.D3 F493 2003.

Feuerstein:2004:OPS

[FPD04] Steven Feuerstein, Bill Pribyl, and Chip Dawes. Oracle PL/SQL language: pocket reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2004. ISBN 0-596-00680-2. viii + 147 pp. LCCN ????

Feuerstein:2008:OPS

[FPD08] Steven Feuerstein, Bill Pribyl, and Chip Dawes. Oracle PL/SQL language: pocket reference. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51404-2 (paperback). vii + 170 pp. LCCN ????

Faroult:2006:AS

[FR06] Stéphane Faroult and Peter Robson. The art of SQL. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-00894-5. xvi + 349 pp. LCCN QA76.73.S67 F376 2006eb; QA76.73.S67. URL http://www.

                                        oreilly.com/catalog/9780596008949.

Freeze:1998:SPR

[Fre98] Wayne S. Freeze. The SQL programmer’s reference: Windows 95/ NT and UNIX. Ventana Press, Chapel Hill, NC, USA, 1998. ISBN 1-56604-760-9. liv + 298 pp. LCCN QA76.73.S67 F74 1998.

Fabbri:1990:WDS

[FS90] Tony Fabbri and A. Robert Schwab. Working with DB2, SQL/DS, SQL, and QMF. Tab Professional and Reference Books, Blue Ridge Summit, PA, USA, 1990. ISBN 0-8306-9508-7 (hardcover), 0-8306-3508-4 (paperback). xiv + 218 pp. LCCN QA76.9.D3 F23 1990. US$28.95 (hardcover), US$18.60 (paperback).

Figgin:2003:LWS

[FWS^{+}03] Stephen Figgin, Aaron Weber, Ellen Siever, Matt Welsh, George Reese, Ben Laurie, Peter Laurie, Rasmus Lerdorf, Stas Bekman, and Eric Cholet, editors. The Linux Web server CD bookshelf. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, version 2.0. edition, 2003. ISBN 0-596-00529-6. LCCN QA76.76.O63.

Gerner:2006:PLL

[G^{+}06] Jason Gerner et al., editors. Professional LAMP: Linux, Apache, MySQL, and PHP Web development. John Wiley and Sons, Inc., New York, NY, USA, 2006. ISBN 0-7645-9723-X. xxii + 379 pp. LCCN TK5105.888 .P677 2006. URL http://www.

loc.gov/catdir/enhancements/
fy0623/2005026487-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0654/2005026487-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0654/2005026487-t.html.

Gabarro:2007:WAD

[Gab07] Steven A. Gabarró. Web application design and implementation: Apache 2, PHP5, MySQL, JavaScript, and Linux/Unix. Quantitative software engineering series. John Wiley and Sons, Inc., New York, NY, USA, 2007. ISBN 0-471-77391-3 (cloth). xv + 295 pp. LCCN TK5105.8883 .G33 2007. URL http://www.

loc.gov/catdir/toc/ecip0613/
2006014999.html.

Gardner:2009:DGP

[Gar09] James Gardner. The definitive guide to Pylons: [Pylons is a lightweight web framework emphasizing flexibility and rapid development using standard tools from the Python community ; includes SQLAlchemy, JavaScript, and WSG!]. The expert’s voice in web development; Books for professionals by professionals. Apress, Berkeley, CA, USA, 2009. ISBN 1-59059-934-9 (paperback). xxv + 536 pp. LCCN ???? US$46,99.

Greenspan:2001:MPD

[GB01] Jay Greenspan and Brad Bulger. MySQL/PHP database applications. Professional mindware. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 2001. ISBN 0-7645-3537-4 (paperback). xxi + 596 pp. LCCN QA76.73.S67 G73 2001. URL

                                        ftp://uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;   http:
                                        //www.loc.gov/catdir/bios/wiley044/
                                        00053995.html;   http://www.
                                        loc.gov/catdir/description/wiley038/
                                        00053995.html;   http://www.
                                        loc.gov/catdir/toc/wiley021/
                                        00053995.html.

Gurry:1996:OPT

[GC96] Mark Gurry and Peter Corrigan. Oracle Performance Tuning. Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, November 1996. ISBN 1-56592-237-9. xxii + 940 pp. LCCN QA76.9.D3 G873 1996. US$44.95.

                                        URL http://www.oreilly.com/
                                        catalog/9781565922372;   http:
                                        //www.oreilly.com/catalog/oracle2.

George:2006:WSS

[GD06] Rajesh George and Lance Delano. Wrox’s SQL Server 2005: express edition starter kit. Programmer to programmer. John Wiley and Sons, Inc., New York, NY, USA, 2006. ISBN 0-7645-8923-7. xxiv + 353 pp. LCCN QA76.9.C55 G46 2006. US$15.99. URL http://www.

loc.gov/catdir/enhancements/
fy0623/2005025426-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0740/2005025426-b.html;   http:
//www.loc.gov/catdir/toc/ecip0518/
2005025426.html.

Gennick:1999:OSD

[Gen99] Jonathan Gennick. Oracle SQL*Plus:

the definitive guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-578-5. xxi + 502 pp. LCCN QA76.73.S67 G46 1999.

Gennick:2000:OSP

[Gen00] Jonathan Gennick. Oracle SQL*Plus:

pocket reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2000. ISBN 1-56592-941-1. iv + 90 pp. LCCN QA76.73.S67 G46 2000.

Gennick:2002:OSP

[Gen02] Jonathan Gennick. Oracle SQL*Plus Pocket Reference: a Guide to SQL*Plus Syntax. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 0-596-00441-9. vi + 113 pp. LCCN QA76.73.S675 G462 2002.

US$12.95. URL http://safari.
oreilly.com/0596004419;   http:
//www.oreilly.com/catalog/orsqlpluspr2.

Gennick:2004:S

[Gen04a] Jonathan Gennick. SQL. Leksykon Kieszonkowy. Helion, Gliwice, Poland, 2004. ISBN 83-7361-596-

2. 191 pp. LCCN ????

Gennick:2004:SPG

[Gen04b] Jonathan Gennick. SQL pocket guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2004. ISBN 0-596-00512-1. vi + 154 pp. LCCN QA76.73.S67 G47 2004.

Gennick:2005:OSP

[Gen05a] Jonathan Gennick. Oracle SQL*Plus: pocket reference. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2005. ISBN 0-596-00885-6. 146 pp. LCCN QA76.9.D3 G455 2000.

URL http://www.oreilly.com/
catalog/9780596008857.

Gennick:2005:OSD

[Gen05b] Jonathan Gennick. Oracle SQL*Plus: the definitive guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2005. ISBN 0-596-00746-9. xxvi + 554 pp. LCCN QA76.73.S67 G46 2005eb; QA76.73.S67; QA76.73.S67 G46

2005. URL http://www.oreilly.
com/catalog/9780596007461.

Gennick:2006:SPG

[Gen06] Jonathan Gennick. SQL pocket guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2006. ISBN 0-596-52688-1 (paperback). viii + 184 pp. LCCN QA76.73.S67

G47 2006. URL http://www.
loc.gov/catdir/enhancements/
fy0715/2007270115-d.html;   http:
//www.loc.gov/catdir/toc/fy0801/
2007270115.html.

George:2011:HDG

[Geo11] Lars George. HBase: The Definitive Guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2011. ISBN 1-4493-9610-0 (paper), 1-4493-1577-1 (e-book). xxviii + 522 pp. LCCN QA76.9.D5 G46 2011. URL http:

//proquest.safaribooksonline.
com/9781449314682.

Gupta:1989:SPG

[GG89] Umang Gupta and William Gietz.

SQL programmer’s guide. Programming series. Que Corporation, Indianapolis, IN, USA, 1989. ISBN 0-88022-390-1. xiv + 302 pp. LCCN QA76.73.S67 G87 1989. US$29.95, CDN$39.95.

Groth:1997:HSL

[GG97] Robert Groth and David Gerber. Hands-on SQL: the language, querying, reporting, and the marketplace. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-486143-4. various pp. LCCN QA76.73.S67 G77 1997.

Glinsky:1999:LIO

[GGK99] Peter Glinsky, Frank Gehrke, and Günter Klappheck. Das LINUX im Office Buch: [erfolgreiches Arbeiten mit wichtigen LINUX-Applikationen; die LINUX-Büropakete Staroffice und Applixware; das Bildbearbeitungsprogramm GIMP; die Datenbank PostgreSQL]. Sybex, Düsseldorf, Germany, 1999. ISBN 3-8155-0310-8. xx + 694 pp. LCCN ????

Gunderloy:2006:MMS

[GH06] Mike Gunderloy and Susan Sales Harkins. Mastering Microsoft SQL Server 2005. John Wiley and Sons, Inc., New York, NY, USA, express edition, 2006. ISBN 0-7821-4402-0. xxii + 576 pp. LCCN QA76.9.D26

G85 2006. URL http://www.
loc.gov/catdir/enhancements/
fy0661/2004117505-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0661/2004117505-d.html;   http:
//www.loc.gov/catdir/toc/fy0705/
2004117505.html.

Gibson:2008:MSS

[Gib08] Darril Gibson. MCITP SQL Server 2005 database administration exam guide (exams 70-431, 70-443, and 70-444). All-in-one. McGraw-Hill, New York, NY, USA, 2008. ISBN 0-07-149609-2. xxxvii + 752 pp. LCCN QA76.3 .G525 2008. URL http://www.

loc.gov/catdir/enhancements/
fy0801/2007042217-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0801/2007042217-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0801/2007042217-t.html.

Gilor:1991:SDP

[Gil91a] Dov Gilor. SQL/DS performance: techniques for improvement. Wiley professional computing. John Wiley and Sons, Inc., New York, NY, USA, 1991. ISBN 0-471-52624-X. xxi + 325 pp. LCCN QA76.9.D3 G524 1991.

Gilor:1991:SPT

[Gil91b] Dov Gilor. SQL/DS performance: techniques for improvement. John Wiley and Sons, Inc., New York, NY, USA, 1991. ISBN 0-471-52624-X. xxi + 325 pp. LCCN ????

Gilfillan:2003:MM

[Gil03] Ian Gilfillan. Mastering MySQL 4. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2003. ISBN 0-7821-4162-5. xxxiii + 729 pp. LCCN QA76.73.S67 G55 2003. US$37.99.

Gilmore:2006:BPM

[Gil06] W. J. Gilmore. Beginning PHP and MySQL 5: from novice to professional. Apress, Berkeley, CA, USA, second edition, 2006. ISBN 1-59059-552-1. xxxiv + 913 pp. LCCN QA76.73.S67 G56 2006.

Glass:2004:BPA

[Gla04] Michael (Michael K.) Glass, editor. Beginning PHP, Apache, MySQL Web development. John Wiley and Sons, Inc., New York, NY, USA, 2004. ISBN 0-7645-5744-0 (paperback). xx + 700 pp. LCCN TK5105.888 .B43 2004. US$26.99. URL

ftp://uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;   http:
//www.loc.gov/catdir/bios/wiley047/
2004101426.html;   http://www.
loc.gov/catdir/description/wiley041/
2004101426.html;   http://www.
loc.gov/catdir/toc/wiley041/
2004101426.html.

Gennick:2001:OSD

[GM01] Jonathan Gennick and Sanjay Mishra. Oracle SQL*Loader: the definitive guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2001. ISBN 1-56592-948-9. xvii + 251 pp. LCCN QA76.73.S67 G46 2001.

Gillette:1995:PDD

[GMT95] Rob Gillette, Dean Muench, and Jean Tabaka. Physical database design for SYBASE SQL Server. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-161523-8. xxix + 225 pp. LCCN QA76.9.D3 G5235 1995.

Garcia-Molina:2002:DSC

[GMUW02] Hector Garcia-Molina, Jeffrey D. Ullman, and Jennifer Widom. Database systems: the complete book. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-031995-3. xxvii + 1119 pp. LCCN QA76.9.D3 G3653 2002.

Greant:2006:MPE

[GN06] Zak Greant and Chris Newman. MySQL phrasebook: essential codes and commands. Developer’s library. Howard W. Sams, Indianapolis, IN 46268, USA, 2006. ISBN 0-672-32839-9 (paperback). ix + 181 pp. LCCN QA76.73.S67 G725 2006.

Gogolla:1994:EER

[Gog94] Martin Gogolla. An extended entity-relationship model: fundamentals and pragmatics, volume 767 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNCSD9. ISBN 0-387-57648-7 (New York), 3-540-57648-7 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 136 pp. LCCN QA76.9.D3 G63 1994. DM39.00.

Gulutzan:1994:OS

[GP94] Peter Gulutzan and Trudy Pelzer.

Optimizing SQL. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-100215-5. x + 230 pp. LCCN QA76.73.S67 G85 1994.

Gulutzan:2003:SPT

[GP03] Peter Gulutzan and Trudy Pelzer.

SQL performance tuning. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-201-79169-2. xix + 495 pp. LCCN QA76.73.S67 G853 2003.

Gray:1993:TPC

[GR93] Jim Gray and A. (Andreas) Reuter.

Transaction processing: concepts and techniques. Morgan Kaufmann series in data management systems. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1993. ISBN 1-55860-190-2. xxxii + 1070 pp. LCCN QA76.545 .G73 1993.

Guerrero:2001:MSS

[GR01] Fernando G. Guerrero and Carlos Eduardo Rojas. Microsoft SQL Server 2000 programming by example. By example series. Que Corporation, Indianapolis, IN, USA, 2001. ISBN 0-7897-2449-9. xvii + 771 pp. LCCN QA76.9.C55 G84 2001.

Grad:2012:RDM

[Gra12] Burton Grad. Relational database management systems: The formative years [Guest editor’s introduction]. IEEE Annals of the History of Computing, 34(4):7–8, October/ December 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Greenspan:2002:MWC

[Gre02] Jay Greenspan. MySQL weekend crash course. Hungry Minds, Cleveland, OH, USA, 2002. ISBN 0-7645-3634-6 (paperback). xxii + 358 pp. LCCN QA76.73.S67 G733 2002. URL

ftp://uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;   http:
//www.loc.gov/catdir/bios/wiley045/
2001093437.html;   http://www.
loc.gov/catdir/description/wiley039/
2001093437.html;   http://www.
loc.gov/catdir/toc/wiley023/
2001093437.html.

Griffin:2002:XSS

[Gri02] John Griffin. XML and SQL Server 2000. New Riders Publishing, Carmel, IN, USA, 2002. ISBN 0-7357-1112-7. xvi + 363 pp. LCCN QA76.9.C55 G77 2002.

Grier:2012:RDC

[Gri12] David Alan Grier. The relational database and the concept of the information system. IEEE Annals of the History of Computing, 34(4):9– 17, October/December 2012. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Groff:2002:SHE

[Gro02] James R. Groff. SQL: the complete reference. McGraw-Hill/ Osborne, Emeryville, CA, USA, 2002. ISBN 0-07-222817-2 (electronic book). xxiv + 1050 pp. LCCN QA76.73.S67 G76 2002eb.

Gruber:1990:US

[Gru90] Martin Gruber. Understanding SQL. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1990. ISBN 0-89588-644-8. xxvi + 434 pp. LCCN QA76.73.S67 G67 1990. US$24.95.

Gruber:1996:S

[Gru96] Martin Gruber. SQL. Helion, Gliwice, Poland, 1996. ISBN 83-86718-32-3. 279 + 1 pp. LCCN ????

Geschwinde:2001:PDH

[GS01a] Ewald Geschwinde and Hans-Jürgen Schönig. PostgreSQL developer’s handbook. Howard W. Sams, Indianapolis, IN 46268, USA, 2001. ISBN 0-672-32260-9. xi + 753 pp.

Geschwinde:2002:PDH

[GS01b] Ewald Geschwinde and Hans-Jürgen Schönig. PostgreSQL: Developer’s Handbook. Howard W. Sams, Indianapolis, IN 46268, USA, 2001. ISBN 0-672-32260-9. xi + 753 pp. LCCN QA76.9.D3 G469 2002. Edited by Patricia Barnes and Scott Meyers.

Geschwinde:2002:PPA

[GS02a] Ewald Geschwinde and Hans-Jürgen Schönig. PHP and PostgreSQL Advanced Web Programming. Howard W. Sams, Indianapolis, IN 46268, USA, 2002. ISBN 0-672-32382-6. xv + 772 pp. LCCN QA76.73.P224 G47 2002.

Geschwinde:2002:PHT

[GS^{+}02b] Ewald Geschwinde, Hans-Jürgen Schönig, et al. PostgreSQL: guide du développeur. CampusPress, Paris, France, 2002. ISBN 2-7440-1387-0 (br). x + 579 pp. LCCN ????

Gilmore:2006:BPP

[GT06] W. J. Gilmore and Robert Treat. Beginning PHP and PostgreSQL 8: from novice to professional. Apress, Berkeley, CA, USA, 2006. ISBN 1-59059-547-5. xxxi + 857 pp. LCCN ????

Gunderloy:2005:MMS

[Gun05] Mike Gunderloy. Mastering Microsoft SQL Server 2005 250. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2005. ISBN 0-7821-4402-0 (paperback). xxii + 576 pp. LCCN ????

Gurry:2002:OST

[Gur02] Mark Gurry. Oracle SQL Tuning: Pocket Reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00268-8. vi + 102 pp. LCCN QA76.73.S67 G87 2002 Stacks. US$12.95.

URL http://safari.oreilly.
com/0596002688;   http://www.
oreilly.com/catalog/orsqltunpr.

Groff:1990:US

[GW90] James R. Groff and Paul N. Weinberg. Using SQL. Osborne/ McGraw-Hill, Berkeley, CA, USA, 1990. ISBN 0-07-881524-X. xxvi + 686 pp. LCCN QA76.73.S67 G37

Groff:1999:SCR

[GW99] James R. Groff and Paul N. Weinberg. SQL, the complete reference. Osborne/McGraw-Hill, Berkeley, CA, USA, 1999. ISBN 0-07-211845-8. xxviii + 994 pp. LCCN QA76.73.S67 G758 1999.

Groff:2002:SCR

[GW02] James R. Groff and Paul N. Weinberg. SQL, the complete reference. Osborne/McGraw-Hill, Berkeley, CA, USA, second edition, 2002. ISBN 0-07-222559-9. xxiv + 1050 pp. LCCN QA76.73.S67 G758 2002. URL http://www.

loc.gov/catdir/enhancements/
fy0620/2004276000-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0620/2004276000-t.html.

Hirt:2004:MSS

[H^{+}04] Allan Hirt et al., editors. Microsoft SQL Server 2000 high availability. Microsoft Press, Bellevue, WA, USA, 2004. ISBN 0-7356-1920-4. xxiv + 754 pp. LCCN QA76.9.C55 M53215 2004.

Haigh:2009:HDG

[Hai09] Thomas Haigh. How data got its base: Information storage software in the 1950s and 1960s. IEEE Annals of the History of Computing, 31(4):6–25, October/December 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Haldar:2007:IS

[Hal07] Sibsankar Haldar. Inside SQLite.

O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2007. ISBN 0-596-55006-5. LCCN QA76.73.S67 H35 2007eb; QA76.73.S67. URL http:/

                                        /www.oreilly.com/catalog/9780596550066.

Hamilton:2005:SSR

[Ham05] Bill Hamilton. SQL Server 2005 reporting essentials. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-52941-

4. LCCN QA76.9.D3; QA76.9.D3 H36 2006eb. URL http://www.

                                        oreilly.com/catalog/9780596529413.

Hamilton:2006:SSR

[Ham06] Bill Hamilton. SQL Server 2005 reporting essentials. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-52941-4. LCCN ????

Hamilton:2007:SSI

[Ham07] Bill Hamilton. SQL Server integration services. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2007. ISBN 0-596-51375-5. LCCN ????

Hamilton:2008:ANC

[Ham08] Bill Hamilton. ADO.NET 3.5 cookbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2008. ISBN 0-596-10140-6. xxii + 955 pp. LCCN QA76.9.D26 H3583 2008.

Hartley:1992:OPP

[Har92] Tim Hartley. ORACLE/SQL: a professional programmer’s guide. McGraw-Hill, New York, NY, USA, 1992. ISBN 0-07-040775-4 (cloth), 0-07-040767-3 (softcover). xx + 543 pp. LCCN ????

Harrington:1998:SCE

[Har98a] Jan L. Harrington. SQL clearly explained. AP Professional, Boston, MA, USA, 1998. ISBN 0-12-326426-X. xiii + 279 pp. LCCN QA76.73.S67 H37 1998.

Harrington:1998:SDK

[Har98b] Jan L. Harrington. SQL dla każdego. Mikom, Warszawa, Poland, 1998. ISBN 83-87102-55-

5. 234 pp. LCCN ????

Harrington:2000:SDK

[Har00] Jan L. Harrington. SQL dla każdego. Mikom, Warszawa, Poland, wyd. 2. edition, 2000. ISBN 83-7279-036-1. 234 pp. LCCN ????

Harms:2001:JSM

[Har01a] David (David Gerhard) Harms.

JSP, servlets, and MySQL. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 2001. ISBN 0-7645-4787-9. xxiv + 499 pp. LCCN QA76.73.J38 H365 2001. URL

ftp://uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;   http:
//www.loc.gov/catdir/bios/wiley045/
2001016954.html;   http://www.
loc.gov/catdir/description/wiley038/
2001016954.html;   http://www.
loc.gov/catdir/toc/wiley021/
2001016954.html.

Harrison:2001:OSH

[Har01b] Guy Harrison. Oracle SQL highperformance tuning. The Prentice Hall PTR Oracle series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 2001. ISBN 0-13-012381-1. xxv + 630 pp. LCCN QA76.9.D3 H365154 2001.

Hartwig:2001:PPP

[Har01c] Jens Hartwig. PostgreSQL: professionell und praxisnah. Addison-Wesley, München, Germany, 2001. ISBN 3-8273-1860-2. 456 pp. LCCN ????

Harrington:2003:SCE

[Har03a] Jan L. Harrington. SQL: clearly explained. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, second edition, 2003. ISBN 1-55860-876-1 (paperback). xvi + 352 pp. LCCN QA76.73.S67 H37

2003. URL http://www.loc.gov/
catdir/description/els031/2003101672.
html;   http://www.loc.gov/catdir/
toc/els031/2003101672.html.

Harris:2003:PMP

[Har03b] Andy Harris. PHP/MySQL programming for the absolute beginner. Absolute beginner. Premier Press, Inc., Indianapolis, IN, USA, 2003. ISBN 1-931841-32-2. ???? pp.

Harris:2004:PMP

[Har04] Andrew Harris. PHP 5/MySQL programming for the absolute beginner. Thomson Course Technology, Boston, MA, USA, 2004. ISBN 1-59200-494-6. xxii + 439 pp. LCCN QA76.73.P224 H37 2004.

Harrington:2005:SDK

[Har05] Jan L. Harrington. SQL dla każdego. Mikom, Warszawa, Poland, wyd. 3. edition, 2005. ISBN 83-7279-460-X. 297 pp. LCCN ???? Translated by Piotr Nowakowski.

Hasan:1996:OSQ

[Has96] Waqar Hasan. Optimization of SQL queries for parallel machines. Lecture notes in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-62065-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xviii + 133 pp. LCCN QA76.73.S67 H38 1996.

Hausman:2001:MSA

[Hau01] Kalani Kirk Hausman. MCSE SQL 2000 administration. Coriolis Group Books, Scottsdale, AZ, USA, 2001. ISBN 1-58880-033-4. xl + 392 pp. LCCN QA76.3 .H375 2001eb.

Holbrook:20xx:SJH

[HCxx] Todd Holbrook and Dan Chudnov, editors. Search jake: PostgreSQL server. Simon Fraser University Library, Burnaby, BC, Canada, 20xx. ISBN ???? LCCN

???? URL ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/
;   http://jake.lib.sfu.ca/;
http://jake.med.yale.edu/;
http://mercury.lib.sfu.ca/~tholbroo/
sfujake-mason/search.html;
http://mercury.lib.sfu.ca/~tholbroo/
sfujake/search.cgi;   http://
www.jake-db.org/.

Harts:2011:MSS

[HDA11] Doug Harts, Jim (Jim L.) Dugan, and Tricia Wilcox Almas. Microsoft SQL Server 2008 R2 analytics and data visualization. McGraw-Hill, New York, NY, USA, 2011. ISBN 0-07-160143-0 (paperback). xxix + 545 + 8 pp. LCCN QA76.9.D343 H47 2011.

Henderson:2000:GGT

[Hen00] Ken Henderson. The guru’s guide to Transact-SQL. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-61576-2. xx + 551 pp. LCCN QA76.73.S67 H47 2000.

Henderson:2002:GGS

[Hen02] Ken Henderson. The Guru’s guide to SQL Server stored procedures, XML, and HTML. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-70046-8. ???? pp. LCCN QA76.9.C55 H45 2002.

Hertzog:2004:DHT

[Her04] Raphaël Hertzog. Debian. Editions Eyrolles, Paris, France, 2004. ISBN 2-212-11398-6. xi + 246 pp. LCCN ????

Hess:2007:SPC

[Hes07] Kenneth Hess. SQL power!: the comprehensive guide. Thomson Course Technology, Boston, MA, USA, 2007. ISBN 1-59863-212-4. xxi + 463 pp. LCCN QA76.73.S67 H49 2007.

Hewitt:2010:CDG

[Hew10] Eben Hewitt. Cassandra: the definitive guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2010. ISBN 1-4493-9041-2 (paperback), 1-4493-9976-2. xxiii + 301 pp. LCCN QA76.9.D3 H4898 2011.

Harrison:2006:MSP

[HF06] Guy Harrison and Steven Feuerstein. MySQL stored procedure programming. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-10089-2. xxiii + 609 pp. LCCN QA76.73.S67 H377 2006eb; QA76.73.S67. URL http://www.

oreilly.com/catalog/9780596100896.

Hanssen:2000:RPS

[HG00] Espen Bergvoll Hanssen and Per Ivar Gjerløw. Resultatsystem for praktisk skyting (PHP/ MySQL/Apache). Høgskolen i Sør-Trøndelag, Avdeling for teknologi, Institutt for databehandling, Trondheim, Norway, 2000. ???? pp. LCCN ????

Horn:2004:MES

[HG04] John Horn and Michael Grey.

MySQL: essential skills. McGraw-Hill/Osborne, Emeryville, CA, USA, 2004. ISBN 0-07-225513-7. xviii + 364 pp. LCCN QA76.73.S67 H676 2004.

Hursch:1988:SSQ

[HH88] Carolyn J. Hursch and Jack L.

Hursch. SQL, the structured query language. Tab Books, Blue Ridge Summit, PA, USA, 1988. ISBN 0-8306-9016-6, 0-8306-3016-3 (paperback). xi + 184 pp. LCCN QA76.73.S67 H87 1988. US$18.60.

Hursch:1989:DIS

[HH89] Jack L. Hursch and Carolyn J.

Hursch. dBase IV SQL user’s guide. Ashton-Tate, Torrance, CA, USA, 1989. ISBN 1-55519-052-9 (Ashton-Tate : paperback), 0-13-198755-0 (Brady Books : paperback). LCCN QA76.9.D3H87 1989.

Hursch:1991:OSD

[HH91a] Carolyn J. Hursch and Jack L. Hursch. Oracle SQL developer’s guide. Windcrest/McGraw-Hill, Blue Ridge Summit, PA, USA, 1991. ISBN 0-8306-2566-6 (hardcover), 0-8306-2529-1 (paperback). xv + 240 pp. LCCN QA76.73.S67 H86 1991.

Hursch:1991:SSQ

[HH91b] Carolyn J. Hursch and Jack L. Hursch. SQL, structured query language. Windcrest/McGraw-Hill, Blue Ridge Summit, PA, USA, second edition, 1991. ISBN 0-8306-8803-X, 0-8306-3803-2. xi + 203 pp. LCCN QA76.73.S67 H87 1991. US$32.95 (hardcover), US$21.60 (paperback).

Hursch:1992:ISD

[HH92] Jack L. Hursch and Carolyn J. Hursch. INGRES SQL developer’s guide. Windcrest/McGraw-Hill, Blue Ridge Summit, PA, USA, 1992. ISBN 0-8306-2564-X (hardcover), 0-8306-2528-3 (paperback). xv + 234 pp. LCCN QA76.9.D26 H87 1992. US$32.95 (hardcover), US$21.60 (paperback).

Hansen:2000:KTL

[HJL00] Ole Kristian Hansen, Fredrik Johnsen, and Inger Helene Lund. Klient… tjener løsning på web, basert på Apache og MySQL ved hjelp av Java server programmering: Support system, volume 222 of Prosjektarbeid / Høgskolen i Hedmark. HHE, Rena, Norway, 2000. 94–?? pp. LCCN ????

Haugen:1995:IGS

[HL95] Signe Haugen and Catherine Lyman. Informix guide to SQL: reference with using triggers. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-100363-1. ???? pp. LCCN QA76.73.S67 H384 1995.

Hartley:1992:OSP

[HM92] Tim Hartley and Tim Martyn. Oracle/SQL: a professional programmer’s guide. J. Ranade IBM series. McGraw-Hill, New York, NY, USA, 1992. ISBN 0-07-040775-4 (hardcover), 0-07-040767-3 (softcover). xx + 543 pp. LCCN QA76.9.D3 H3652 1992. US$49.95 (hardcover), US$34.95 (softcover).

Hardman:2005:EOP

[HM05] Ron Hardman and Michael McLaughlin. Expert Oracle PL/SQL. McGraw-Hill Professional, New York, NY, USA, 2005. ISBN 0-07-226194-3 (paperback). 496 pp. LCCN MLCM 2006/41083

(Q). URL http://www.loc.gov/
catdir/enhancements/fy0653/2006280508-
t.html;   http://www.loc.gov/
catdir/enhancements/fy0668/2006280508-
b.html;   http://www.loc.gov/
catdir/enhancements/fy0668/2006280508-
d.html.

Harris:2013:SDI

[HN13] Hershel Harris and Bert Nicol.

SQL/DS: IBM’s first RDBMS. IEEE Annals of the History of Computing, 35(2):69–71, April/ June 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Holub:2004:HPL

[Hol04] Allen I. Holub. Holub on Patterns: Learning Design Patterns by Looking at Code. Apress, Berkeley, CA, USA, 2004. ISBN 1-59059-388-X. xv + 414 pp. LCCN QA76.76.P37 H65 2004. US$49.95. URL http://www.

                                        holub.com/software/holubSQL/.

Hotka:2007:OSD

[Hot07] Dan Hotka. Oracle SQL developer handbook. McGraw-Hill, New York, NY, USA, 2007. ISBN 0-07-148474-4. xxii + 230 pp. LCCN QA76.73.S67 H6763 2007. URL http://www.

                                        loc.gov/catdir/enhancements/
                                        fy0814/2007276583-b.html;   http:
                                        //www.loc.gov/catdir/enhancements/
                                        fy0814/2007276583-d.html;   http:
                                        //www.loc.gov/catdir/toc/fy0714/
                                        2007276583.html.

Hotek:2008:MSS

[Hot08] Mike Hotek. Microsoft SQL Server 2008 step by step. Microsoft Press, Bellevue, WA, USA, 2008. ISBN 0-7356-2604-9. ???? pp. LCCN ????

Hotek:2009:MSP

[Hot09] Mike Hotek. MCTS self-paced training kit (exam 70-432): Microsoft SQL Server 2008—implementation and maintenance. Microsoft Press, Bellevue, WA, USA, 2009. ISBN 0-7356-2605-7. ???? pp. LCCN ????

Houlette:2001:SBG

[Hou01a] Forrest Houlette. SQL: a beginner’s guide. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-213096-2. xxiii + 424 pp. LCCN QA76.73.S67 H677 2001.

Houlette:2001:TS

[Hou01b] Forrest Houlette. Troubleshooting SQL. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-213489-5. xxiii + 357 pp. LCCN QA76.73.S67 H678 2001. URL

ftp://uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;   http:
//www.loc.gov/catdir/description/
mh024/2001277837.html;   http:
//www.loc.gov/catdir/toc/mh023/
2001277837.html.

Hovell:1995:ISA

[Hov95] John F. Hovell. Interactive SQL on the AS/400. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-107186-6. xiv + 160 pp. LCCN QA76.73.S67 H68 1995.

Harinath:2006:PSS

[HQ06] Sivakumar Harinath and Stephen R.

Quinn. Professional SQL Server analysis services 2005 with MDX. John Wiley and Sons, Inc., New York, NY, USA, 2006. ISBN 0-7645-7918-5 (paperback). xxvi + 828 pp. LCCN QA76.9.C55 H363 2006. URL http://www.

loc.gov/catdir/enhancements/
fy0623/2005032272-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0740/2005032272-b.html;   http:
//www.loc.gov/catdir/toc/ecip063/
2005032272.html.

Haderle:2013:HGI

[HS13] Donald J. Haderle and Cynthia M. Saracco. The history and growth of IBM’s DB2. IEEE Annals of the History of Computing, 35(2): 54–66, April/June 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hernandez:2000:SQM

[HV00] Michael J. (Michael James) Hernandez and John Viescas. SQL queries for mere mortals: a handson guide to data manipulation in SQL. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-43336-2. xxxii + 492 pp. LCCN QA76.73.S67 H48 2000.

IBM:2009:HAD

[IBM09a] IBM. High Availability and Disaster Recovery Options for DB2 on Linux, Unix, and Windows. IBM redbooks. IBM Redbooks, 11400 Burnet Road, Austin, TX 78758-3493, USA, 2009. ISBN 0-7384-3138-9. xvi + 856 pp. LCCN ????

IBM:2009:MOI

[IBM09b] IBM. Migrating from Oracle to IBM Informix Dynamic Server on Linux, UNIX, and Windows. Redbooks. IBM Redbooks, 11400 Burnet Road, Austin, TX 78758-3493, USA, 2009. ISBN 0-7384-3302-0. xiv + 378 pp. LCCN ????

Informix:1995:IGSa

[Inf95a] Informix Software, Inc. Informix guide to SQL: reference, with Using triggers. Informix Press, Menlo Park, CA, USA, 1995. ISBN 0-13-100363-1. various pp. LCCN QA76.73.S67 H384 1995.

Informix:1995:IGSb

[Inf95b] Informix Software, Inc. Informix guide to SQL, tutorial. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-100371-2. various pp. LCCN QA76.73.S67 I54 1995.

Informix:2000:IGSa

[Inf00a] Informix Software, Inc. Informix guide to SQL: reference and syntax. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 2000. ISBN 0-13-016165-9. various pp. LCCN QA76.73.S67 I53 2000.

Informix:2000:IGSb

[Inf00b] Informix Software, Inc. Informix guide to SQL: reference and syntax. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 2000. ISBN 0-13-016166-7. various pp. LCCN QA76.73.S67 I53 2000.

Informix:2000:IGSc

[Inf00c] Informix Software, Inc. Informix guide to SQL: tutorial. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-016165-9. various pp. LCCN QA76.73.S67 I54 2000.

ISO:1992:IITa

[Int92] International Organization for Standardization. ISO/IEC 9075:1992: Title: Information technology — Database languages — SQL. International Organization for Standardization, Geneva, Switzerland, 1992. ISBN ???? 587 pp. LCCN ???? CHF 348. URL http://www.

iso.ch/cate/d16663.html. Avail-

able in English only.

ISO:1995:IIIa

[Int95a] International Organization for Standardization. ISO/IEC 12227:1995: Information technology — Programming languages — SQL/ Ada Module Description Language (SAMeDL). International Organization for Standardization, Geneva, Switzerland, 1995. ISBN ???? 120 pp. LCCN ???? CHF 188. URL

http://www.ansi.org/;   http:
//www.iso.ch/cate/d2345.html.

ISO:1995:IIIk

[Int95b] International Organization for Standardization. ISO/IEC 12227:1995: Information technology — Programming languages — SQL/ Ada Module Description Language (SAMeDL). International Organization for Standardization, Geneva, Switzerland, 1995. ISBN ???? 120 pp. LCCN ???? CHF 188; US$128.00. URL http://

webstore.ansi.org/ansidocstore/
product.asp?sku=ISO%2FIEC+12227%
3A1995; http://www.iso.ch/cate/

d2345.html. Available in English only.

ISO:1995:IIIh

[Int95c] International Organization for Standardization. ISO/IEC 9075-3:1995: Information technology — Database languages — SQL — Part 3: Call-Level Interface (SQL/ CLI). International Organization for Standardization, Geneva, Switzerland, 1995. ISBN ???? 401 pp. LCCN ???? CHF 306.

URL http://www.iso.ch/cate/
d30609.html.

ISO:1996:IIIa

[Int96] International Organization for Standardization. ISO/IEC 9579-3:1996: Information technology — Open Systems Interconnection — Remote Database Access — Part 3: SQL specialization Protocol Implementation Conformance Statement (PICS) proforma. International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? ???? pp. LCCN ????

US$125. URL http://www.ansi.
org/.

ISO:1998:IIIb

[Int98] International Organization for Standardization. ISO/IEC 9579-2:1998: Information technology — Open Systems Interconnection — Remote Database Access — Part 2: SQL Specialization. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? ???? pp. LCCN ???? US$105.

URL http://www.ansi.org/.

ISO:1999:IIIu

[Int99a] International Organization for Standardization. ISO/IEC 13249-3:1999: Information technology — Database languages — SQL Multimedia and Application Packages — Part 3: Spatial. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 332 pp. LCCN ???? CHF

                                        288. URL http://www.iso.ch/
                                        cate/d28054.html.

ISO:1999:IIIn

[Int99b] International Organization for Standardization. ISO/IEC 9075-1:1999: Information technology — Database languages — SQL — Part 1: Framework (SQL/ Framework). International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 67 pp. LCCN ???? CHF 144.

                                        URL http://www.iso.ch/cate/
                                        d26196.html.

ISO:1999:IIIo

[Int99c] International Organization for Standardization. ISO/IEC 9075-2:1999: Information technology — Database languages — SQL — Part 2: Foundation (SQL/ Foundation). International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 1121 pp. LCCN ???? CHF 402.

                                        URL http://www.iso.ch/cate/
                                        d26197.html.

ISO:1999:IIIv

[Int99d] International Organization for Standardization. ISO/IEC 9075-4:1999: Information technology — Database languages — SQL — Part 4: Persistent Stored Modules (SQL/PSM). International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 152 pp. LCCN ???? CHF

                                        200. URL http://www.iso.ch/
                                        cate/d29864.html.

ISO:19xx:IIDk

[Intxxa] International Organization for Standardization. ISO/IEC 9075-5:1999: Information technology —- Database languages — SQL — Part 5: Host Language Bindings (SQL/ Bindings). International Organization for Standardization, Geneva, Switzerland, 19xx. ISBN ???? 243 pp. LCCN ???? CHF 236.

URL http://www.iso.ch/cate/
d26198.html.

ISO:19xx:IIDj

[Intxxb] International Organization for Standardization. ISO/IEC DIS 9075-10: Information technology — Database languages — SQL — Part 10: Object Language Bindings (SQL/OLB). International Organization for Standardization, Geneva, Switzerland, 19xx. ISBN ???? ???? pp. LCCN ???? URL http://www.

iso.ch/cate/d30613.html.

ISO:2000:IIIc

[Int00] International Organization for Standardization. ISO/IEC 9579:2000: Information technology — Remote database access for SQL with security enhancement. International Organization for Standardization, Geneva, Switzerland, 2000. ISBN ???? 115 pp. LCCN ???? CHF

188. URL http://www.iso.ch/
cate/d30615.html.

Iseminger:2001:MSS

[Ise01] David Iseminger, editor. Microsoft SQL Server 2000 reference library. Microsoft Press, Bellevue, WA, USA, 2001. ISBN 0-7356-1280-3 (set). ???? pp. LCCN QA76.9.C55 M5323 2001.

ISO:2008:IIIc

[ISO08a] ISO. ISO/IEC 9075-10:2008 Information technology — Database languages — SQL — Part 10: Object Language Bindings (SQL/OLB). International Organization for Standardization, Geneva, Switzerland, July 17, 2008. ISBN

????  376 pp.  LCCN

???? URL http://www.iso.org/
iso/iso_catalogue/catalogue_
tc/catalogue_detail.htm?csnumber=
38644.

ISO:2008:IIId

[ISO08b] ISO. ISO/IEC 9075-13:2008 Information technology — Database languages — SQL — Part 13: QL Routines and Types Using the Java TM Programming Language (SQL/JRT). International Organization for Standardization, Geneva, Switzerland, July 17, 2008. ISBN ???? 186 pp. LCCN

???? URL http://www.iso.org/
iso/iso_catalogue/catalogue_
tc/catalogue_detail.htm?csnumber=
38646.

ISO:2008:IIIa

[ISO08c] ISO. ISO/IEC 9075-3:2008 Information technology — Database languages — SQL — Part 3: Call-Level Interface (SQL/CLI). International Organization for Standardization, Geneva, Switzerland, July 17, 2008. ISBN

????  378 pp.  LCCN

???? URL http://www.iso.org/
iso/iso_catalogue/catalogue_

tc/catalogue_detail.htm?csnumber=
38641.

ISO:2008:IIIb

[ISO08d] ISO. ISO/IEC 9075-9:2008 Information technology — Database languages — SQL — Part 9: Management of External Data (SQL/MED). International Organization for Standardization, Geneva, Switzerland, July 17, 2008. ISBN ???? 451 pp. LCCN

???? URL http://www.iso.org/
iso/iso_catalogue/catalogue_
tc/catalogue_detail.htm?csnumber=
38643.

ISO:2011:IIIc

[ISO11a] ISO. ISO/IEC 9075-1:2011 Information technology — Database languages — SQL — Part 1: Framework (SQL/Framework). International Organization for Standardization, Geneva, Switzerland, December 15, 2011. ISBN

????  ????  pp.  LCCN

???? URL http://www.iso.org/
iso/iso_catalogue/catalogue_
tc/catalogue_detail.htm?csnumber=
53681.

ISO:2011:IIIg

[ISO11b] ISO. ISO/IEC 9075-14:2011 Information technology — Database languages — SQL — Part 14: XML-Related Specifications (SQL/XML). International Organization for Standardization, Geneva, Switzerland, December 15, 2011. ISBN

????  422 pp.  LCCN

???? URL http://www.iso.org/
iso/iso_catalogue/catalogue_
tc/catalogue_detail.htm?csnumber=
53686.

ISO:2011:IIIf

[ISO11c] ISO. ISO/IEC 9075-2:2011 Information technology — Database languages — SQL — Part 11: Information and Definition Schemas (SQL/Schemata). International Organization for Standardization, Geneva, Switzerland, December 15, 2011. ISBN ???? 282 pp. LCCN

                                        ???? URL http://www.iso.org/
                                        iso/iso_catalogue/catalogue_
                                        tc/catalogue_detail.htm?csnumber=
                                        53685.

ISO:2011:IIId

[ISO11d] ISO. ISO/IEC 9075-2:2011 Information technology — Database languages — SQL — Part 2: Foundation (SQL/Foundation). International Organization for Standardization, Geneva, Switzerland, December 15, 2011. ISBN ???? 1434 pp. LCCN ????

ISO:2011:IIIe

[ISO11e] ISO. ISO/IEC 9075-4:2011 Information technology — Database languages — SQL — Part 4: Persistent Stored Modules (SQL/PSM). International Organization for Standardization, Geneva, Switzerland, December 15, 2011. ISBN

????  164 pp.  LCCN

                                        ???? URL http://www.iso.org/
                                        iso/iso_catalogue/catalogue_
                                        tc/catalogue_detail.htm?csnumber=
                                        53684.

Jones:2005:SFP

[J^{+}05] Arie Jones et al., editors. SQL functions programmer’s reference. Wrox Press, Chicago, IL, USA; Birmingham, UK, 2005. ISBN 0-7645-6901-5 (paperback). xl + 751 pp. LCCN QA76.73.S67 S674 2005; QA76.73.S67 S674 2005eb; QA329.42 .N66 2005eb S674 2005eb.

Jepson:1998:OGM

[JH98] Brian Jepson and David J. Hughes.

Official guide to Mini SQL 2.0. John Wiley, New York, NY, USA, 1998. ISBN 0-471-24535-6 (paperback/CD-ROM). xiii + 417 pp. LCCN QA76.73.S67 J47 1998.

Jorden:2006:MMS

[Jor06] Joseph L. Jorden. MCTS: Microsoft SQL Server 2005 implementation and maintenance study guide (exam 70-431). John Wiley and Sons, Inc., New York, NY, USA, 2006. ISBN 0-470-02565-4 (paperback). xxxvi + 634 pp. LCCN Z52.5.M52 .J67 2006. URL http://www.

loc.gov/catdir/enhancements/
fy0741/2007270254-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0741/2007270254-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0741/2007270254-t.html.

Joshi:1997:CGJ

[JRC97] Daniel I. Joshi, Rodney Runolfson, and Ramesh Chandak. The Comprehensive Guide to the JDBC SQL API. Ventana Press, Chapel Hill, NC, USA, 1997. ISBN 1-56604-637-8. xxiv + 422 pp. LCCN QA76.625.J67 1997.

US$50. URL http://www.vmedia.
com/bookshop/book.html?isbn=

1566046378. Includes CD-ROM.

Jorden:2009:MSS

[JWO09] Joseph L. Jorden, Dandy Weyn, and Dave Owen. Microsoft SQL Server 2005 implementation and maintenance (70-431). John Wiley and Sons, Inc., New York, NY, USA, 2009. ISBN 0-470-11596-3. 381 pp. LCCN QA76.3 .J69 2009.

Knight:2007:PSS

[K^{+}07] Brian Knight et al., editors. Professional SQL Server 2005 administration. Wrox professional guides. Wrox/Wiley Technology, Indianapolis, IN, USA, 2007. ISBN 0-470-05520-0 (paper/website). xxx + 735 pp. LCCN QA76.9.D3 P7667 2007. URL http://www.

                                        loc.gov/catdir/enhancements/
                                        fy0741/2006032940-b.html;   http:
                                        //www.loc.gov/catdir/enhancements/
                                        fy0741/2006032940-d.html;   http:
                                        //www.loc.gov/catdir/toc/ecip072/
                                        2006032940.html.

Knight:2010:KMB

[K^{+}10] Brian Knight et al., editors. Knight’s Microsoft Business Intelligence 24-hour trainer. John Wiley and Sons, Inc., New York, NY, USA, 2010. ISBN 0-470-88963-2 (paperback). xxix + 394 pp. LCCN ???? US$29.99.

Kaplan:2001:ISS

[Kap01] Michael S. Kaplan. Internationalization with SQL Server. Macmillan Computer Publishing, Indianapolis, IN, USA, 2001. ISBN 0-672-32099-

1. ???? pp. LCCN ???? Edited by Sharon Cox.

Karwin:2010:SAA

[Kar10] Bill Karwin. SQL antipatterns:

avoiding the pitfalls of database programming. Pragmatic programmers. Pragmatic Bookshelf, Raleigh, NC, USA, 2010. ISBN 1-934356-55-7 (paperback). xii + 333 pp. LCCN QA76.73.S67 K37 2010.

Kline:1999:TSP

[KGZ99a] Kevin Kline, Lee Gould, and Andrew Zanevsky. Transact-SQL programming. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-401-0. xxvi + 807 pp. LCCN ????

Kline:1999:TP

[KGZ99b] Kevin E. Kline, Lee Gould, and Andrew Zanevsky. Transact-SQL programming. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-401-0. xxvi + 807 pp. LCCN QA76.73.S67 K55 1999.

Kline:2008:SN

[KHK08] Kevin E. Kline, Brand Hunt, and Daniel Kline. SQL in a nutshell. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2008. ISBN 0-596-51884-6. xii + 578 pp. LCCN ????

Khoshafian:1992:GDC

[Kho92] Setrag Khoshafian. A Guide to developing client/server SQL applications. Morgan Kaufmann series in data management systems. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1992. ISBN 1-55860-147-3. ISSN 1046-1698. xvii + 634 pp. LCCN QA76.9.D3 G835 1992.

Kiessling:1984:SQC

[Kie84] Werner Kiessling. SQL-like and QUEL-like correlation queries with aggregates revisited. Memorandum UCB/ERL M84/75, Electronics Research Laboratory, College of Engineering, University of California, Berkeley, Berkeley, CA, USA, September 17, 1984. 14 pp.

Kiessig:2009:UFA

[Kie09] Richard Kiessig. Ultra-fast ASP.NET: building ultra-fast and ultra-scalable Web sites using ASP.NET and SQL Server. The expert’s voice in .NET. Apress, Berkeley, CA, USA, 2009. ISBN 1-4302-2383-9 (paperback). xxi + 460 pp. LCCN TK5105.8885.A26 K56 2009.

King:2002:STT

[KJ02] Konrad King and Kris A. Jamsa. SQL tips and techniques. Premier Press, Inc., Indianapolis, IN, USA, 2002. ISBN 1-931841-45-4. xxi + 1127 pp. LCCN QA76.73.S67 K53 2002.

Kline:2001:SND

[KK01] Kevin E. Kline and Daniel Kline. SQL in a nutshell: a desktop quick reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2001. ISBN 1-56592-744-3. viii + 214 pp. LCCN QA76.73.S67 K55 2001.

Kline:2002:SAO

[KK02] Kevin E. Kline and Daniel Kline.

SQL: almanach: opis poleceń jȩzyka. Helion, Gliwice, Poland, 2002. ISBN 83-7197-595-3. 226 + 1 pp. LCCN ???? Translation by Paweł Janociński.

Kofler:2005:DGM

[KK05] Michael Kofler and David Kramer.

The definitive guide to MySQL5. The expert’s voice in open source. Apress, Berkeley, CA, USA, third edition, 2005. ISBN 1-59059-535-1 (paperback). xxxi + 748 pp. LCCN QA76.73.S67 K64 2005.

Kline:2004:SN

[KKH04] Kevin E. Kline, Daniel Kline, and Brand Hunt. SQL in a nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2004. ISBN 0-596-00481-8. xv + 691 pp. LCCN QA76.73.S67 K54 2004; QA76.73.S67 K55 2004; QA76.73.S67; QA76.73.S67 K55 2004eb. URL http://www.

oreilly.com/catalog/9780596004811.

Kreines:1999:ODA

[KL99] David C. Kreines and Brian Laskey.

Oracle Database Administration: The Essential Reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-516-

5. xvii + 560 pp. LCCN QA76.9.D3 K7344 1999. US$36.95.

URL http://www.oreilly.com/
catalog/9781565925168;   http:
//www.oreilly.com/catalog/oradba/

.

Klein:2006:PSS

[Kle06] Scott Klein. Professional SQL Server 2005 XML. Wrox professional guides. Wrox/Wiley Pub., Indianapolis, IN, 2006. ISBN 0-7645-9792-2 (paperback). xx + 526 pp. LCCN QA76.9.C55 K545 2006.

Klute:1998:JPD

[Klu98] Rainer Klute. JDBC in der Praxis: Datenbankanwendungen in Intranet und Internet; [Entwicklungssoftware und Tools: Java development kit (JDK), Java servlet development kit (JSDK), PostgreSQL u.v.m.]. Addison-Wesley Longman, Reading, MA, USA, 1998. ISBN 3-8273-1301-5. 287 (est.) pp. LCCN ???? Includes CD-ROM.

Koletzke:2007:OJF

[KM07] Peter Koletzke and Duncan Mills. Oracle JDeveloper 10g for forms and PL/SQL developers. Mc-Graw-Hill, New York, NY, USA, 2007. ISBN 0-07-225960-4 (paperback), (paperback). xxx + 562 pp. LCCN QA76.73.J38 K655 2006. URL http://www.

                                        loc.gov/catdir/enhancements/
                                        fy0806/2007298524-b.html;   http:
                                        //www.loc.gov/catdir/enhancements/
                                        fy0806/2007298524-d.html;   http:
                                        //www.loc.gov/catdir/enhancements/
                                        fy0806/2007298524-t.html.

Kofler:2001:M

[Kof01] Michael Kofler. MySQL. Books for professionals by professionals. Apress, Berkeley, CA, USA, 2001. ISBN 1-893115-57-7. xi + 659 pp. LCCN ????

Kopstad:2004:WFW

[Kop04] Magnus Kopstad. Weblogg-Forum web-system: (PHP/MySQL). Høgskolen i Sør-Trøndelag, Avdeling for informatikk og e-læring, Trondheim, Norway, 2004. ???? pp. LCCN ????

Kreines:2000:OSE

[Kre00a] David C. Kreines. Oracle SQL: the essential reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2000. ISBN 1-56592-697-8. xxxiii + 382 pp. LCCN QA76.73.S67 K74 2000.

Kreines:2000:OSH

[Kre00b] David C. Kreines. Oracle SQL: the essential reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2000. ISBN ???? xxxiii + 382 pp. LCCN ????

Kruntorad:2009:HCI

[Kru09] Judith M. Kruntorad. History of the CA IDMS Database Management System. IEEE Annals of the History of Computing, 31(4):100– 106, October/December 2009. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kriegel:2003:SB

[KT03] Alex Kriegel and Boris M. Trukhnov.

SQL bible. John Wiley and Sons, Inc., New York, NY, USA, 2003. ISBN 0-7645-2584-0 (paperback). xxvii + 831 pp. LCCN QA76.73.S67

K76 2003. URL http://www.
loc.gov/catdir/bios/wiley046/
2003101832.html;   http://www.

loc.gov/catdir/description/wiley039/
2003101832.html;   http://www.
loc.gov/catdir/toc/wiley041/
2003101832.html.

Kriegel:2008:SB

[KT08] Alex Kriegel and Boris M. Trukhnov. SQL bible. John Wiley and Sons, Inc., New York, NY, USA, second edition, 2008. ISBN 0-470-22906-3. xxx + 857 pp. LCCN ????

Kuhnel:1997:JFP

[Küh97] Ralf Kühnel. Die Java 1.1 Fibel: Programmierung von Threads und Applets, mit den neuen Bibliotheken: Beans, Security, RMI, IDL, SQL. Addison-Wesley, Reading, MA, USA, second edition, 1997. ISBN 3-8273-1233-7. 352 (or 333??) pp. LCCN ????

49,90 DM. URL http://www.
addison-wesley.de/;   http://
www.addison-wesley.de/katalog/
item.ppml?textexpr=Java&id=00212;
 http://www.addison-wesley.de/
KD/NE/1024.html. Includes CD-

ROM.

Kuznetsov:2010:DDP

[Kuz10] Alex Kuznetsov. Defensive database programming with SQL server. Red Gate Books, Cambridge, UK, 2010. ISBN 1-906434-45-X (paperback). 380 pp. LCCN ???? US$19.99.

Knight:2007:ESS

[KV07] Brian Knight and Erik Veerman. Expert SQL Server 2005 integration services. Wrox programmer to programmer. John Wiley and Sons, Inc., New York, NY, USA, 2007. ISBN 0-470-13411-9 (paper/website). xxi + 404 pp. LCCN QA76.9.D3 K6134 2007. US$31.99. URL http://www.

loc.gov/catdir/enhancements/
fy0741/2007013520-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0741/2007013520-d.html;   http:
//www.loc.gov/catdir/toc/ecip0714/
2007013520.html.

Ladanyi:2000:S

[Lad00] Hans Ladányi. SQL. Ksiȩga Eksperta. Helion, Gliwice, Poland, 2000. ISBN 83-7197-120-6. 880 pp. LCCN ????

Lafferty:1997:MTG

[Laf97] David Lafferty. MCSE training guide. SQL Server 6.5 administration. New Riders Publishing, Carmel, IN, USA, 1997. ISBN 1-56205-726-X. xvi + 704 pp. LCCN QA76.3 .M334 1997b.

Luers:1997:TYP

[LAG97] Tom Luers, Timothy Atwood, and Jonathan Gennick. Teach yourself PL/SQL in 21 days. Howard W. Sams, Indianapolis, IN 46268, USA, 1997. ISBN 0-672-31123-2. xxxi + 646 pp.

Lahdenmaki:2005:RDI

[Lah05] Tapio Lahdenmaki. Relational database index design and the optimizers DB2, Oracle, SQL Server, et al. Wiley-Interscience, New York, NY, USA, 2005. ISBN 0-471-71999-4 (cloth), 0-471-72137-9 (electronic book), 0-471-71999-4. xvii + 310 pp. LCCN QA76.9.D3 L335 2005eb.

Lam:2011:HA

[Lam11] Chuck Lam. Hadoop in action. Manning, Greenwich, CT, USA, 2011. ISBN 1-935182-19-6. xxi + 312 pp. LCCN QA76.9.D5 L36 2011.

Lancaster:2004:TDR

[Lan04a] Deac Lancaster. Transact-SQL desk reference. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2004. ISBN 0-13-029339-3 (paperback). xvii + 806 pp. LCCN QA76.9.D3 L348 2004.

Lancaster:2004:TSD

[Lan04b] Deac Lancaster. Transact-SQL desk reference. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2004. ISBN 0-13-029339-3 (paperback). xvii + 806 pp. LCCN QA76.9.D3 L348 2004.

Larson:1991:SSI

[Lar91] Bruce L. Larson. SQL solutions for IBM DBMS. VNR computer library. Van Nostrand Reinhold Co., New York, NY, USA, 1991. ISBN 0-442-00119-3. xv + 553 pp. LCCN QA76.73.S67 L37 1991.

Larson:1993:SSI

[Lar93] Bruce L. Larson. SQL solutions for IBM DBMS. Van Nostrand Reinhold Co., New York, NY, USA, 1993. ISBN 0-442-00119-3. xv + 553 pp. LCCN ????

Lardiere:2006:PHT

[Lar06a] Sébastien Lardière. PostgreSQL 8.1: administration et exploitation d’une base de données. Editions ENI, St Herblain, France, 2006. ISBN 2-7460-3468-9. 259 (est.) pp. LCCN ????

Larson:2006:DBI

[Lar06b] Brian Larson. Delivering business intelligence with Microsoft SQL Server 2005. Essential skills for database professionals. McGraw-Hill/Osborne, Emeryville, CA, USA, 2006. ISBN 0-07-148688-7, 0-07-226090-4. xiii + 776 pp. LCCN HD38.7 .L37 2006eb.

Larson:2006:MSS

[Lar06c] Brian Larson. Microsoft SQL Server 2005 reporting services. McGraw-Hill/Osborne, Emeryville, CA, USA, 2006. ISBN 0-07-226239-7 (paperback). xvi + 767 pp. LCCN ????

Lardiere:2007:PHT

[Lar07] Sébastien Lardière. PostgreSQL:

administration et exploitation d’une base de données. Editions ENI, St Herblain, France, second edition, 2007. ISBN 2-7460-3961-3. ISSN 1627-8224. 270 (est.) pp.

LCCN ???? URL http://www.
eni-livres.com.

Lobel:2008:PMS

[LBF08] Leonard Lobel, Andrew J. Brust, and Stephen Forte. Programming Microsoft SQL Server 2008. Microsoft Press, Bellevue, WA, USA, 2008. ISBN 0-7356-2599-9. ???? pp. LCCN ????

Loney:1999:OSP

[LC99] Kevin Loney and Rachel Carmichael.

Oracle SQL and PL/SQL annotated archives. McGraw-Hill, New York, NY, USA, 1999. ISBN 0-07-882536-9. xviii + 585 pp. LCCN QA76.73.S67 L66 1999.

Lorie:1991:SA

[LD91] Raymond A. Lorie and Jean-Jacques Daudenarde. SQL and its applications. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1991. ISBN 0-13-837956-4. xv + 299 pp. LCCN QA76.73.S67L67 1991.

Leffler:1991:UIS

[Lef91] Jonathan Leffler. Using INFORMIX-SQL. Addison-Wesley, Reading, MA, USA, second edition, 1991. ISBN 0-201-56509-9. xiii + 320 pp. LCCN QA76.9.D3 L429 1991. US$43.00 (est.).

Langit:2009:SBI

[LGMM09] Lynn Langit, Kevin S. Goff, Davide Mauri, and Sahil Malik. Smart business intelligence solutions with Microsoft SQL server 2008. Microsoft Press, Bellevue, WA, USA, 2009. ISBN 0-7356-2580-

8. xxx + 765 pp. LCCN HD38.7 .S63 2009. URL http://www.

loc.gov/catdir/enhancements/
fy1004/2008940532-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy1010/2008940532-t.html.

Linoff:2007:DAU

[Lin07] Gordon Linoff. Data analysis using SQL and Excel. John Wiley and Sons, Inc., New York, NY, USA, 2007. ISBN 0-470-09951-8. xli + 645 pp. LCCN QA76.73.S67

L56 2007. URL http://www.
loc.gov/catdir/enhancements/
fy0740/2007026313-b.html;   http:

//www.loc.gov/catdir/enhancements/
fy0740/2007026313-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0740/2007026313-t.html.

Linoff:2008:DAU

[Lin08] Gordon S. Linoff. Data analysis using SQL and Excel. John Wiley and Sons, Inc., New York, NY, USA, 2008. ISBN 0-470-09951-8 (paper/website). xli + 645 pp. LCCN QA76.73.S67

L56 2008. URL http://www.
loc.gov/catdir/enhancements/
fy0740/2007026313-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0740/2007026313-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0740/2007026313-t.html.

Lis:2007:S

[Lis07] Marcin Lis. SQL. Ćwiczenia Praktyczne. Wydawnictwo Helion, Gliwice, Poland, 2007. ISBN 83-246-0621-1. 151 pp. LCCN ????

Lomasky:1998:OS

[LK98] Brian Lomasky and David C.

Kreines. Oracle Scripts. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, May 1998. ISBN 1-56592-438-X. xvi + 185 pp. LCCN QA76.73.S67 L66 1998; QA76.9.D3 L648 1998. US$29.95.

URL http://www.oreilly.com/
catalog/9781565924383;   http:
//www.oreilly.com/catalog/oraclescrp.

Covers Oracle8.

Lockhart:2000:PPG

[Loc00] Thomas Lockhart, editor. PostgreSQL Programmer’s Guide. iUniverse.com, Lincoln, NE, USA, 2000. ISBN 0-595-14917-0. xxvii + 488 pp. LCCN ???? US$34.95, EUR 38.95.

Linsenbardt:2001:SSA

[LS01] Mark A. Linsenbardt and Shane Stigler. SQL Server 2000 administration. Database professional’s library. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-212618-3 (book and CD-ROM), 0-07-212616-7 (book), 0-07-212617-5 (CD-ROM). xx + 586 pp. LCCN QA76.9.C55 L56 2001.

Lacy-Thompson:1991:ITR

[LT91] Tony Lacy-Thompson. INFORMIX-SQL: a tutorial and reference. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1991. ISBN 0-13-465121-9. xii + 264 pp. LCCN QA76.9.D3 L33 1991. US$35.95.

Lusardi:1988:DEG

[Lus88a] Frank Lusardi. The database experts’ guide to SQL. McGraw-Hill computing that works. McGraw-Hill, New York, NY, USA, 1988. ISBN 0-07-039006-1, 0-07-039002-9 (paperback). xiv + 224 pp. LCCN QA76.73.S67 L87 1988.

Lusardi:DEG88

[Lus88b] Frank Lusardi. The Database Experts’ Guide to SQL. McGraw-Hill, New York, NY, USA, 1988. ISBN 0-07-039006-1 (hardcover), 0-07-039002-9 (paperback). xiv + 224 pp. LCCN QA76.73.S67 L87 1988. Prepared with TEX.

Lee:2003:OSW

[LW03] James Lee and Brent Ware.

Open source Web development with LAMP: using Linux, Apache, MySQL, Perl, and PHP. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-201-77061-X (paperback). xxxiv + 460 pp. LCCN QA76.76.D47 L435 2003.

Lynch:1990:US

[Lyn90] Elizabeth Lynch. Understanding SQL. Macmillan computer science series. Macmillan Education, Basingstoke, Hampshire, England, 1990. ISBN 0-333-52433-0 (paperback). xvii + 174 pp. LCCN QA76.73.S67 L96 1990. UK£12.99.

McClure:2006:PAN

[M^{+}06] Wallace B. McClure et al., editors.

Professional ADO.NET 2: programming with SQL Server 2005, Oracle, and MySQL. John Wiley and Sons, Inc., New York, NY, USA, 2006. ISBN 0-7645-8437-5 (paperback). xxviii + 614 pp. LCCN QA76.9.D37 P763 2006. URL http://www.

loc.gov/catdir/enhancements/
fy0653/2006278082-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0659/2006278082-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0659/2006278082-t.html.

Mackin:2007:MSS

[M^{+}07] J. C. Mackin et al., editors. Microsoft SQL Server 2005: database administrator core requirements: self-paced training kit. Microsoft Press, Bellevue, WA, USA, 2007.

ISBN 0-7356-2377-5. ???? pp. LCCN QA76.3 .M5284 2007. Three volumes.

Mitchell:2003:RWS

[MA03] Allan Mitchell and Mark Allison. Real-world SQL-DMO for SQL Server. The expert’s voice. Apress, Berkeley, CA, USA, 2003. ISBN 1-59059-040-6. xxvii + 397 pp. LCCN ????

Mackin:2007:MST

[Mac07] J. C. Mackin. MCITP self-paced training kit (Exam 70-443): designing a database server infrastructure using Microsoft SQL Server 2005. Microsoft Press, Bellevue, WA, USA, 2007. ISBN 0-7356-2173-X. xxxii + 708 pp. LCCN QA76.3 .M3224 2007.

Maguire:2009:ASA

[Mag09] John Norris Maguire. Adabas: Software AG of North America. IEEE Annals of the History of Computing, 31(4):92–99, October/ December 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Malcher:2011:ODA

[Mal11] Michelle Malcher. Oracle database administration for Microsoft SQL Server DBAs. McGraw-Hill, New York, NY, USA, 2011. ISBN 0-07-174431-2 (paperback). xiv + 334 pp. LCCN QA76.9.D3 M3325 2011. US$37.99.

Martyn:1989:DPP

[Mar89] Tim Martyn. DB2/SQL: a professional programmer’s guide. Mc-Graw-Hill, New York, NY, USA, 1989. ISBN 0-07-040666-9. xxix + 472 pp. LCCN ????

Martinsson:2001:SXW

[Mar01] Tobias Martinsson. Scripting XML and WMI for Microsoft SQL Server 2000: professional developer’s guide. Professional developer’s guide series. John Wiley, New York, NY, USA, 2001. ISBN 0-471-39951-5 (paperback). xviii + 413 pp. LCCN QA76.76.H94 M279 2001. Includes CD-ROM.

Mariel:2002:PHT

[Mar02] Stéphane Mariel. PostgreSQL: services Web professionnels avec PostgreSQL et PHP/XML. Editions Eyrolles, Paris, France, 2002. ISBN 2-212-11166-5 (br). x + 122 pp. LCCN ????

Martin:2005:RSI

[Mar05] Steve W. Martin. The real story of Informix Software and Phil White: Lessons in business and leadership for the executive team. Sand Hill Publishing, Rancho Santa Margarita, CA, USA, 2005. ISBN 0-9721822-2-5. xx + 208 pp. LCCN HD9696.65.U64.

Mauer:2001:STY

[Mau01] Lowell Mauer. Sams teach yourself Transact-SQL in 21 days. Howard W. Sams, Indianapolis, IN 46268, USA, 2001. ISBN 0-672-31967-5. xvi + 621 pp. LCCN QA76.73.S67 M338 2001.

Maxfield:2000:MPS

[Max00] Wade Maxfield. MySQL and PHP from scratch. Jesse Liberty’s from scratch programming series.

Que Corporation, Indianapolis, IN, USA, 2000. ISBN 0-7897-2440-5. xi + 372 pp. LCCN ????

Maxfield:2001:MPS

[Max01] Wade Maxfield. MySQL and PHP from scratch. Jesse Liberty’s from scratch programming series. Que Corporation, Indianapolis, IN, USA, 2001. ISBN 0-7897-2440-5. xi + 372 pp. LCCN QA76.73.S67 M327 2001.

Maslakowski:2000:STY

[MB00a] Mark Maslakowski and Tony Butcher. Sams teach yourself MySQL in 21 days. Howard W. Sams, Indianapolis, IN 46268, USA, 2000. ISBN 0-672-31914-4. xix + 532 pp. LCCN ????

Mazlakowski:2000:STY

[MB00b] Mark Mazlakowski and Tony Butcher. Sams teach yourself MySQL in 21 days. Macmillan Computer Publishing, Indianapolis, IN, USA, 2000. ISBN 0-672-31914-4. xix + 532 pp. LCCN QA76.73.S67 M328 2000.

Mishra:2002:MOS

[MB02] Sanjay Mishra and Alan Beaulieu. Mastering Oracle SQL. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00129-0. xiv + 321 pp. LCCN QA76.73.S67 M57 2002 Stacks. US$34.95.

                                        URL http://safari.oreilly.
                                        com/0596001290;   http://www.
                                        oreilly.com/catalog/mastorasql.

Mishra:2004:MOS

[MB04] Sanjay Mishra and Alan Beaulieu.

Mastering Oracle SQL. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2004. ISBN 0-596-00632-2. xvii + 472 pp. LCCN QA76.9.D3 M5787 2004.

Melton:2006:QXX

[MB06] Jim Melton and Stephen Buxton. Querying XML: XQuery, XPath, and SQL/XML in context. The Morgan Kaufmann series in data management systems. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2006. ISBN 1-55860-711-0 (paperback). xxvii + 815 pp. LCCN QA76.76.H94 M448 2006. URL http://www.

loc.gov/catdir/enhancements/
fy0668/2006298085-d.html.

McBath:2002:SSB

[McB02] Frank McBath. SQL Server backup and recovery: tools and techniques. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-062298-2 (paperback). xvi + 292 pp. LCCN QA76.9.C55 M26 2002.

McCullough:1994:OPS

[McC94] Alastair E. McCullough. Oracle PL/SQL for SQL*Forms: an introduction to PL/SQL for those using SQL *Forms version 3.0. Mc-Graw-Hill, New York, NY, USA, 1994. ISBN 0-07-707925-6. xx + 169 pp. LCCN QA76.73.S67 M33 1994. UK£22.50.

McClanahan:1996:ODG

[McC96] David McClanahan. Oracle developer’s guide. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1996. ISBN ???? xviii + 794 pp. LCCN QA76.9.D3 M39375 1996.

McCarty:1998:VDS

[McC98] Bill McCarty. Visual Developer SQL Database Programming with Java: Creating Fast, Efficient Database Applications for the Web. Coriolis Group Books, Scottsdale, AZ, USA, 1998. ISBN 1-57610-176-2. xvii + 493 pp. LCCN QA76.9.D3M39367 1998. US$39.99, CDN$55.99.

                                        URL http://www.coriolis.com/
                                        cgi-win/KBSearch.exe;   http:
                                        //www.coriolis.com/site/msie/
                                        books/ind/hpjsqudp.htm.

McClure:2005:SDS

[McC05] Russell A. McClure. SQL DOM: strongly typed dynamic SQL statements. Thesis (m.s.), University of California, San Diego, San Diego, CA, USA, 2005. viii + 47 pp.

McDonald:2003:MOP

[McD03] Connor McDonald. Mastering Oracle PL/SQL: Practical Solutions. Apress, Berkeley, CA, USA, 2003. ISBN 1-59059-217-4. 616 (est.) pp. LCCN ???? US$49.99. URL http://www.

                                        apress.com/book/bookDisplay.
                                        html?bID=276.

McGehee:1998:STY

[McG98] Brad McGehee. Sam’s teach yourself MCSE SQL Server 6.5 Administration in 14 days. Howard W. Sams, Indianapolis, IN 46268, USA, 1998. ISBN 0-672-31312-X. xiii + 484 pp.

McGehee:1999:UMS

[McG99] Brad McGehee. Using Microsoft SQL Server 7.0. Que Corporation, Indianapolis, IN, USA, 1999. ISBN 0-7897-1628-3. xvii + 746 pp.

Matthews:2003:MJD

[MCG03] Mark Matthews, Jim Cole, and Joe Gradecki. MySQL and Java developer’s guide. Java open source library. John Wiley and Sons, Inc., New York, NY, USA, 2003. ISBN 0-471-26923-9. xxii + 410 pp. LCCN QA76.3.S67 M38 2003. URL

ftp://uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;   http:
//www.loc.gov/catdir/bios/wiley045/
2002155887.html;   http://www.
loc.gov/catdir/description/wiley038/
2002155887.html;   http://www.
loc.gov/catdir/toc/wiley031/
2002155887.html.

McGee:2009:IMS

[McG09] William C. McGee. The Information Management System (IMS) program product. IEEE Annals of the History of Computing, 31(4): 66–75, October/December 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

McMahon:2001:MSS

[McM01] Richard Alan McMahon. MCSE SQL Server 2000 database design. Coriolis Group Books, Scottsdale, AZ, USA, 2001. ISBN 1-58880-034-

2. xl + 421 pp. LCCN QA76.3 .G553 2001eb.

Patrick:1997:IU

[McN97] John (John Patrick) McNally. Informix unleashed. Howard W. Sams, Indianapolis, IN 46268, USA, 1997. ISBN 0-672-30650-6. xxii + 1195 pp. LCCN QA76.9.D3 I5332 1997.

MacLennan:2008:DMM

[MCT08] Jamie MacLennan, Bogdan Crivat, and Zhaohui Tang. Data mining with Microsoft SQL Server 2008. John Wiley and Sons, Inc., New York, NY, USA, 2008. ISBN 0-470-27774-2. ???? pp. LCCN QA76.9.D343 M335 2008.

McGoveran:1992:GSS

[MD92] D. McGoveran and C. J. Date. A guide to SYBASE and SQL Server: a user’s guide to the SYBASE product (a relational database management system with application development facilities) from Sybase, Inc. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-55710-X. xxi + 548 pp. LCCN QA76.9.D3 M3953 1992.

Melton:2000:USJ

[ME00] Jim Melton and Andrew Eisenberg. Understanding SQL and Java together: a guide to SQLJ, JDBC, and related technologies. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2000. ISBN 1-55860-562-2. xxix + 514 pp. LCCN QA76.73.S67 M43 2000.

Melton:2001:SUR

[Mel01] Jim Melton. SQL: 1999 – Understanding Relational Language Components. Elsevier, Amsterdam, The Netherlands, 2001. ISBN 0-08-051760-9 (electronic book). 929 pp. LCCN ????

Meloni:2002:STY

[Mel02] Julie C. Meloni. Sams teach yourself MySQL in 24 hours. Howard W. Sams, Indianapolis, IN 46268, USA, 2002. ISBN 0-672-32349-4. xiv + 349 pp. LCCN QA76.3.S67 M45 2002.

Meloni:2003:STY

[Mel03a] Julie C. Meloni. SAMS teach yourself PHP, MySQL and Apache in 24 hours. Howard W. Sams, Indianapolis, IN 46268, USA, 2003. ISBN 0-672-32489-X. xvii + 507 pp. LCCN QA76.73.P224 M48 2003.

Melton:2003:ASU

[Mel03b] Jim Melton. Advanced SQL, 1999:

understanding object-relational and other advanced features. The Morgan Kaufmann series in data management systems. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 1-55860-677-7. xxviii + 563 pp. LCCN QA76.73.S67 M422

2003. URL http://www.loc.gov/
catdir/description/els031/2002107329.
html;   http://www.loc.gov/catdir/
toc/els031/2002107329.html.

Meloni:2004:STY

[Mel04] Julie C. Meloni. Sams teach yourself PHP, MYSQL and Apache: all in one. Howard W. Sams, Indianapolis, IN 46268, USA, 2004. ISBN 0-672-32620-5. xxiii + 616 pp. LCCN TK5105.888 .M45 2004.

Mendelsohn:2013:OS

[Men13] Andrew Mendelsohn. The Oracle story: 1984–2001. IEEE Annals of the History of Computing, 35(2): 10–23, April/June 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Meteyer:2001:QSD

[Met01] Lisa C. Meteyer. Quantification in SQL database query language. Thesis (m.s.), University of California, Santa Cruz, Santa Cruz, CA, USA, 2001. viii + 164 pp.

Martyn:1989:DSP

[MH89] Tim Martyn and Tim Hartley. DB2/SQL: a professional programmer’s guide. J. Ranade IBM series. Intertext Publications, New York, NY, USA, 1989. ISBN 0-07-040666-9. xxix + 472 pp. LCCN QA76.73.S67 M39 1989.

Millsap:2003:OOP

[MH03] Cary V. Millsap and Jeff Holt. Optimizing Oracle performance. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00527-X. xxv + 388 pp. LCCN QA76.9.D3 M577 2003.

                                        URL http://www.oreilly.com/
                                        catalog/9780596005276.

Martyn:1995:SPP

[MHJ95] Tim Martyn, Tim Hartley, and Richard Johnson. SQL/400: a professional programmer’s guide. J. Ranade IBM series. McGraw-Hill, New York, NY, USA, 1995. ISBN 0-07-040799-1. xix + 499 pp. LCCN QA76.8.I25919 M37 1995.

Mackin:2007:MSP

[MHTH07] J. C. Mackin, Mike Hotek, Tobias Thernström, and Shannon Horn. MCITP self-paced training kit (Exam 70-443): designing a database server infrastructure using Microsoft SQL Server 2005. Microsoft Press, Bellevue, WA, USA, 2007. ISBN 0-7356-2173-X. xxxii + 708 pp. LCCN QA76.3 .M3224

2007. URL http://www.loc.gov/
catdir/toc/fy0711/2006932079.
html.

Microsoft:1997:MSM

[Mic97] Microsoft Corporation. Microsoft systems management server training. Microsoft Press, Bellevue, WA, USA, 1997. ISBN 1-57231-614-4. xxix + 786 pp. LCCN TA168 .M54 1997. “Hands-on, self-paced training for version 1.2”–Cover Disc 1 includes: supplemental course materials; an HTML based course; and evaluation copies of Microsoft Systems management server and Microsoft SQL server. Disc 2 contains an evaluation copy of Microsoft Windows NT server.

Microsoft:1999:MBR

[Mic99] Microsoft Corporation. Microsoft BackOffice 4.5 resource kit. Microsoft Press, Bellevue, WA, USA, 1999. ISBN 0-7356-0583-1. ???? pp. LCCN QA76.9.C55 M525 1999.

Misner:2009:MSS

[Mis09] Stacia Misner. Microsoft SQL Server 2008 reporting services step by step. Microsoft Press, Bellevue, WA, USA, 2009. ISBN 0-7356-2647-2. ???? pp. LCCN ????

Mitchell:2003:RSS

[Mit03] Allan Mitchell. Real-world SQL-DMO for SQL Server. The expert’s voice; Books for professionals by professionals. Apress, Berkeley, CA, USA, 2003. ISBN 1-59059-040-6. xxvii + 397 pp. LCCN QA76.9.D26 M58 2003.

Machanic:2007:ESS

[MKR07] Adam Machanic, Hugo Kornelis, and Lara Rubbelke. Expert SQL Server 2005 development: [advanced SQL Server techniques for database professionals]. The expert’s voice in SQL server programming. Apress, Berkeley, CA, USA, 2007. ISBN 1-59059-729-X. xx + 449 pp. LCCN ????

Morrison:2007:CSS

[MM07] Charles Michael Morrison and Joline Morrison. Creating SQL Server 2005 applications with Visual Studio. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, 2007. ISBN 0-13-146355-1. xxvi + 573 pp. LCCN QA76.9.D3 M6585 2007. URL http://www.

                                        loc.gov/catdir/toc/ecip0616/
                                        2006021771.html.

Molinaro:2006:SC

[Mol06] Anthony Molinaro. SQL cookbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-00976-3. xxix + 595 pp. LCCN QA76.73.S67 M64

                                        2006. URL http://www.gbv.de/
                                        du/services/toc/bs/501083510;
                                         http://www.oreilly.com/catalog/
                                        9780596009762.

Momjian:2001:PEK

[Mom01a] Bruce Momjian. PostgreSQL :

Einführung und Konzepte. Addison-Wesley, München, Germany, 2001. ISBN 3-8273-1859-9. 543 (est.) pp. LCCN ???? URL

http://www.addison-wesley.de/
projector/projector.asp?page=
bookdetails&isbn=3827318599.

Momjian:2001:PIC

[Mom01b] Bruce Momjian. PostgreSQL: Introduction and Concepts. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-70331-9 (paperback). xxviii + 462 pp. LCCN QA76.9.D3 M647 2001. URL

http://www.addison-wesley.de/
projector/projector.asp?page=
bookdetails&isbn=3827318599;
http://www.ca.postgresql.org/
docs/awbook.html;   http://www.
pearsoned.co.jp/washo/db/wa_
db27-j.html.

Moncur:2003:MYV

[Mon03] Michael G. Moncur. MySQL:

your visual blueprint to open source database management. Visual read less, learn more. John Wiley and Sons, Inc., New York, NY, USA, 2003. ISBN 0-7645-1692-2. xv + 290 pp. LCCN QA76.3.S67 M66 2003. URL

ftp://uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;   http:
//www.loc.gov/catdir/bios/wiley045/
2002110260.html;   http://www.
loc.gov/catdir/description/wiley039/
2002110260.html;   http://www.
loc.gov/catdir/toc/wiley031/
2002110260.html.

Morin:20xx:PPD

[Morxxa] Rich Morin. PostgreSQL: Programming and Development. Daemon News, 560 South State Street, Suite B2 Orem, UT 84057, USA, 20xx. ISBN ???? 427 (est.) pp. LCCN ???? US$35.00. URL http://www.

                                        ptf.com/dossier/Post.shtml.

Morin:20xx:PRM

[Morxxb] Rich Morin. PostgreSQL: Reference Manual. Daemon News, 560 South State Street, Suite B2 Orem, UT 84057, USA, 20xx. ISBN ???? 414 (est.) pp. LCCN ????

                                        US$35.00. URL http://www.ptf.
                                        com/dossier/Post.shtml.

Morin:20xx:PUA

[Morxxc] Rich Morin. PostgreSQL: Use and Administration. Daemon News, 560 South State Street, Suite B2 Orem, UT 84057, USA, 20xx. ISBN ???? 427 pp. LCCN ????

                                        US$35.00. URL http://www.ptf.
                                        com/dossier/Post.shtml.

Melton:1993:UNS

[MS93] Jim Melton and Alan R. Simon. Understanding the new SQL: a complete guide. Morgan Kaufmann series in data management systems. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1993. ISBN 1-55860-245-3. xxiv + 536 pp. LCCN QA76.73.S67 M44 1993.

McEwan:1997:TYT

[MS97] Bennett Wm. (Bennett William) McEwan and David S. Solomon. Teach yourself Transact-SQL in 21 days. Howard W. Sams, Indianapolis, IN 46268, USA, 1997. ISBN 0-672-31045-7. xxvi + 548 pp.

Matthew:2000:PLP

[MS^{+}00] Neil Matthew, Richard Stones, et al. Professional Linux programming. Wrox Press, Chicago, IL, USA; Birmingham, UK, 2000. ISBN 1-86100-301-3. xviii + 1155 pp. LCCN QA76.76.O63 P754 2000.

Melton:2002:SUR

[MS02] Jim Melton and Alan R. Simon. SQL:1999: understanding relational language components. Morgan Kaufmann series in data management systems. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2002. ISBN 1-55860-456-1. xxxii + 895 pp. LCCN QA76.73.S67 M5 2002.

Matthew:2005:BDP

[MS05] Neil Matthew and Richard Stones.

Beginning databases with PostgreSQL: from novice to professional. Apress, Berkeley, CA, USA, second edition, 2005. ISBN 1-59059-478-9. xxiv + 637 pp. LCCN QA76.9.D3 S7554 2005.

Mauer:2000:STY

[MSM00] Lowell Mauer, David S. Solomon, and Bennett Wm (Bennett William) McEwan. Sams teach yourself Transact-SQL in 21 days. Howard W. Sams, Indianapolis, IN 46268, USA, second edition, 2000. ISBN 0-672-31967-5. xvi + 621 pp. LCCN QA76.73.S67 M326 2000.

Mata-Toledo:2000:SOF

[MTC00] Ramon A. Mata-Toledo and Pauline K. Cushman. Schaum’s outline of fundamentals of SQL programming. Schaum’s outline series. McGraw-Hill, New York, NY, USA, 2000. ISBN 0-07-135953-2 (paperback). ix + 314 pp. LCCN QA76.73.S67 M325 2000.

Mundy:2006:MDW

[MTK06] Joy Mundy, Warren Thornthwaite, and Ralph Kimball. The Microsoft Data Warehouse Toolkit: with SQL Server 2005 and the Microsoft Business Intelligence Toolset. John Wiley and Sons, Inc., New York, NY, USA, 2006. ISBN 0-471-26715-5 (paper/website). xliii + 746 pp. LCCN QA76.9.D37

M95 2006. URL http://www.
loc.gov/catdir/enhancements/
fy0623/2005032275-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0740/2005032275-b.html;   http:
//www.loc.gov/catdir/toc/ecip063/
2005032275.html.

MySQL:2005:MAG

[MyS05] MySQL AB. MySQL: administrator’s guide. MySQL Press, Indianapolis, IN, USA, 2005. ISBN 0-672-32634-5. xvii + 672 pp. LCCN QA76.73.S67 M98 2005.

Nielsen:2010:SSM

[N^{+}10] Paul Nielsen et al., editors. SQL Server MVP deep dives. Manning, Greenwich, CT, USA, 2010. ISBN 1-935182-04-8 (paperback). xliv + 804 pp. LCCN ????

Nanda:2005:OPD

[Nan05] Arup Nanda. Oracle PL/SQL for DBAs. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN ???? xxi + 429 pp. LCCN ????

Nanda:2007:HMS

[Nan07] Ashwani Nanda. Hands-on Microsoft SQL Server 2005 integration services. McGraw-Hill, New York, NY, USA, 2007. ISBN 0-07-226319-

9. xix + 570 pp. LCCN QA76.9.D3 N328 2007. URL http://www.

loc.gov/catdir/enhancements/
fy0810/2007297494-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0810/2007297494-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0810/2007297494-t.html.

Naramore:2005:BPA

[Nar05] Elizabeth Naramore, editor. Beginning PHP5, Apache, and MySQL Web development. John Wiley and Sons, Inc., New York, NY, USA, 2005. ISBN 0-7645-7966-

5. xviii + 798 pp. LCCN TK5105.888 .B425 2005. URL

ftp://uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;   http:
//www.loc.gov/catdir/toc/ecip054/
2004027093.html.

Nath:1990:GSS

[Nat90a] Aloke Nath. The guide to SQL Server. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-52336-

1. xxi + 498 pp. LCCN QA76.9.D3 N348 1990. US$24.95, CDN$31.95.

NCSL:1990:DLS

[Nat90b] National Computer Systems Laboratory (U.S.). Database language SQL. National Technical Information Service, Washington, DC, USA, February 2, 1990. 8 pp. Supersedes FIPS PUB 127 - 1987 March 10. Shipping list no.: 90-353-P.

Newcomer:1992:SSR

[New92] Lawrence R. Newcomer. Select– SQL: the relational database language. Maxwell Macmillan International, New York, NY, USA, 1992. ISBN 0-02-386694-2. xv + 446 pp. LCCN QA76.73.S67 N48 1992.

Newman:2005:S

[New05] Chris Newman. SQLite. Developer’s library. Howard W. Sams, Indianapolis, IN 46268, USA, 2005. ISBN 0-672-32685-X (paperback), 0-7686-6366-0 (e-book). xvi + 313 pp. LCCN QA76.73.S67 N488 2005eb. URL http:/

/proquest.safaribooksonline.
com/067232685X;   http://proquestcombo.
safaribooksonline.com/067232685X;
 http://www.contentreserve.com/
TitleInfo.asp?ID={5D380338-820F-
4EC3-BAA3-668B6BF49170}\%26Format=
50; http://www.myilibrary.com?
id=108361.

Nanda:2005:OPS

[NF05] Arup Nanda and Steven Feuerstein. Oracle PL/SQL for DBAs. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-00587-3. xxi + 429 pp. LCCN QA76.73.P258; QA76.73.P258eb.

URL http://www.oreilly.com/
catalog/9780596005870.

Nielsen:2007:SSB

[Nie07] Paul Nielsen. SQL Server 2005 bible. John Wiley and Sons, Inc., New York, NY, USA, 2007. ISBN 0-7645-4256-7 (paper/website). xlv + 1293 pp. LCCN QA76.9.C55

N55 2007. URL http://www.
loc.gov/catdir/enhancements/
fy0740/2006016151-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0740/2006016151-d.html;   http:
//www.loc.gov/catdir/toc/ecip0613/
2006016151.html.

Nies:2009:CST

[Nie09] Thomas M. (Tom) Nies. Cincom Systems’ Total. IEEE Annals of the History of Computing, 31(4):55–61, October/December 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Nystrom:2007:SID

[Nys07] Marty Nystrom. SQL injection defenses. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2007. ISBN 0-596-52964-3. LCCN ????

Otey:1999:SSD

[OC99] Michael Otey and Paul Conte.

SQL Server 7 developer’s guide. Database professional’s library. Osborne/McGraw-Hill, Berkeley, CA, USA, 1999. ISBN 0-07-882548-2. xxxii + 975 pp. LCCN QA76.9.C55 O84 1999.

Otey:2001:SSD

[OC01] Michael Otey and Paul Conte.

SQL Server 2000 developer’s guide.

Database professional’s library. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-212569-1. xxix + 1248 pp. LCCN QA76.9.C55 O843 2001. Includes CD-ROM.

Odewahn:1999:OWA

[Ode99] Andrew Odewahn. Oracle Web applications: PL/SQL developer’s introduction. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-687-0. xi + 241 pp. LCCN QA76.625 .O34 1999.

OGorman:1998:PFM

[O’G98] Kevin O’Gorman. A parallel framework for maintaining SQL views at a data warehouse. Thesis (m.s.), 1998, University of California, Santa Barbara, Santa Barbara, CA, USA, 1998. vi + 35 pp.

Hsu:2012:PRH

[OH12] Regina Obe and Leonard Hsu. PostgreSQL: Up and running. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2012. ISBN 1-4493-2632-3, 1-4493-2633-1. xiv + 145 pp. LCCN QA76.9.D3 O24 2012. EUR 16.00. URL http:

//proquest.safaribooksonline.
com/9781449326326.

ONeil:1994:DPP

[O’N94] Patrick O’Neil. Database— principles, programming, performance. The Morgan Kaufmann series in data management systems. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1994. ISBN 1-55860-219-4. xiv + 873 pp. LCCN QA76.9.D3 O489 1994.

ONeil:2001:DPP

[OO01] Patrick O’Neil and Elizabeth O’Neil. Database—principles, programming, and performance. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, second edition, 2001. ISBN 1-55860-580-0 (paperback), 1-55860-438-3 (hardcover). xxiv + 870 pp. LCCN QA76.9.D3 O489

2001. URL http://www.loc.gov/
catdir/description/els033/99089041.
html;   http://www.loc.gov/catdir/
toc/els033/99089041.html.

Oppel:2005:SD

[Opp05] Andrew J. Oppel. SQL demystified.

McGraw-Hill/Osborne, Emeryville, CA, USA, 2005. ISBN 0-07-148672-0, 0-07-226224-9. xvii + 312 pp. LCCN QA76.73.S67 O77 2005eb; QA76.9.D3 O77 2005.

OReilly:2000:OPS

[O’R00] O’Reilly & Associates, Inc. The Oracle PL/SQL CD Bookshelf. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, July 2000. ISBN 1-56592-849-0. 285 pp. LCCN ???? US$89.95. URL http://www.

oreilly.com/catalog/oraclecdbs.

Includes CD-ROM.

Oracle:1990:SAS

[Ora90a] Oracle Corporation. SQL advanced skills. Redwood Shores, CA, USA, 1990.

Oracle:1990:SAP

[Ora90b] Oracle Corporation. SQL application programming. Redwood Shores, CA, USA, 1990.

Oracle:1990:SF

[Ora90c] Oracle Corporation. SQL foundation. Redwood Shores, CA, USA, version 1.0 edition, 1990. The SQL Foundation set consists of tutorial disks (5 disks for the 5.25in environment and 3 disks for the 3.5in environment), RDBMS and SQL*PLUS disks (5 disks for the 5.25in environment and 8 disks for the 3.5in environment); the SQL* Foundation workbook; 2 quick reference cards, 1 booklet titled Introduction to SQL, and 1 booklet titled SQL Foundation start here. “SQL Foundation is a training course that teaches one to use Structured Query Language (SQL) via the online tutorial that simulates actual SQL queries and reinforces one’s knowledge with exercises. SQL can also be used through a workshop facility that provides a workbook containing exercises using an actual relational database management system (RDBMS) and sample tables.” Oracle off-the-shelf training/CBT TD System requirements: IBM AT, PS/2, or an IBM-compatible 286 or 386 computer; a hard disk and 640K of random access memory; 896K of extended memory; approximately 8.5 megabytes of free disk space on the hard disk (approximately 2 megabytes if SQL*PLUS and the RDBMS is already installed); DOS 2.1 or later; color monitor recommended.

Oracle:1990:SFH

[Ora90d] Oracle Corporation. SQL foundation. Redwood Shores, CA, USA, 1990. 21 computer disks + 1 workbook (xii + 188) + 2 booklets + 2 quick reference cards pp.

Otey:2006:MSS

[Ote06] Michael Otey. Microsoft SQL Server 2005 developer’s guide. Mc-Graw-Hill/Osborne, Emeryville, CA, USA, 2006. ISBN 0-07-148348-9, 0-07-226099-8. xx + 507 pp. LCCN ????

Otey:2010:MSS

[Ote10] Michael Otey. Microsoft SQL Server 2008 high availability with clustering and database mirroring. McGraw-Hill, New York, NY, USA, 2010. ISBN 0-07-149813-3. xii + 291 pp. LCCN QA76.9.D3 O8655 2010. URL http://www.

loc.gov/catdir/enhancements/
fy1011/2009049571-b.htm;   http:
//www.loc.gov/catdir/enhancements/
fy1011/2009049571-d.htm;   http:
//www.loc.gov/catdir/enhancements/
fy1011/2009049571-t.htm.

Owens:1996:BID

[Owe96] Kevin T. Owens. Building intelligent databases with Oracle PL/ SQL, triggers, and stored procedure. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-443631-8 (paperback). xiv + 370 pp. LCCN QA76.9.D3 O92 1996.

Owens:1998:BID

[Owe98] Kevin T. Owens. Building intelligent databases with Oracle PL/ SQL, triggers, and stored procedures. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 1998. ISBN 0-13-794314-8. xvi + 544 pp. LCCN QA76.9.D3 O92 1998.

Owens:2006:DGS

[Owe06] Michael Owens. The Definitive Guide to SQLite. Apress, Berkeley, CA, USA, 2006. ISBN 1-59059-673-0. xxi + 440 pp. LCCN QA76.73.S67 O94 2006.

                                        URL http://www.apress.com/
                                        book/bookDisplay.html?bID=10130.

Pachev:2006:UMI

[Pac06] Alexander Pachev. Understanding MySQL internals. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-00957-7. ???? pp. LCCN QA76.73.S67; Internet; QA76.73.S67 .P33eb.

Pascal:1990:SRB

[Pas90] Fabian Pascal. SQL and relational basics. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1990. ISBN 1-55851-063-X. 336 + 2 pp. LCCN QA76.73.S67 P37 1990. US$28.95.

Pascal:1993:URD

[Pas93] Fabian Pascal. Understanding relational databases with examples in SQL-92. Wiley professional computing. John Wiley and Sons, Inc., New York, NY, USA, 1993. ISBN 0-471-58538-6 (paper). xxii + 278 pp. LCCN QA76.9.D3 P3486 1993. US$29.95.

Patrick:1999:SF

[Pat99] John Patrick. SQL Fundamentals. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-096016-0. 820 pp. URL http:

//www.phptr.com/ptrbooks/ptr_
013014469X.html. Includes CD-

ROM. Covers Microsoft Access and Oracle.

Patrick:2002:SF

[Pat02] John J. Patrick. SQL fundamentals. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 2002. ISBN 0-13-066947-4. xxvi + 834 pp. LCCN QA76.73.S67 P38 2002.

Patrick:2009:IC

[Pat09] Robert L. Patrick. IMS @ Conception. IEEE Annals of the History of Computing, 31(4):62–65, October/ December 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Polk:1993:SID

[PB93] William T. Polk and Lawrence E.

Bassham. Security issues in the database language SQL. Superintendent of Documents, US GPO, Washington, DC, USA, August 1993. vi + 40 pp. US$3.25. Shipping list number 94-0041-M.

Perng:1998:SCQ

[Per98a] Chang-Shing Perng. SQL/LPP+:

a cascading query language for temporal correlation verification in time series databases. Technical report, UCLA Computer Science Department, Los Angeles, CA, 1998. 24 pp.

Perng:1998:SLC

[Per98b] Chang-Shing Perng. SQL/LPP+:

a cascading query language for temporal correlation verification in time series databases. Technical report, UCLA Computer Science Department, Los Angeles, CA, USA, 1998. 24 pp.

Perng:1998:STS

[Per98c] Chang-Shing Perng. SQL/LPP: a time series extension of SQL based on limited patience patterns. Technical report, UCLA Computer Science Department, Los Angeles, CA, 1998. 26 pp.

Perkins:2001:P

[Per01] Jeff Perkins. PostgreSQL. Premier Press, Inc., Indianapolis, IN, USA, 2001. ISBN 1-931841-42-X. xviii + 366 pp. LCCN QA76.9.D3 P459 2001. US$39.99.

Petkovic:2000:SSB

[Pet00] Dusan Petkovic. SQL Server 2000: a beginner’s guide. Database professional’s library. Osborne/McGraw-Hill, Berkeley, CA, USA, 2000. ISBN 0-07-212587-X. xxi + 663 pp. LCCN QA76.9.C55 P48 2000.

Petkovic:2005:MSS

[Pet05] Dušan Petković. Microsoft SQL Server 2005: a beginner’s guide ; [essential skills for database professionals]. McGraw-Hill/Osborne, Emeryville, CA, USA, 2005. ISBN 0-07-226093-9. xxix + 704 pp. LCCN QA76.9.C55. USD 39.99.

Petkovic:2006:MSS

[Pet06] Dušan Petković. Microsoft SQL Server 2005: a beginner’s guide. Essential skills for database professionals. McGraw-Hill/Osborne, Emeryville, CA, USA, 2006. ISBN 0-07-226093-9 (paperback). xxix + 704 pp. LCCN QA76.9.C55 P478 2006. URL http://www.

loc.gov/catdir/enhancements/
fy0665/2006279408-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0665/2006279408-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0665/2006279408-t.html.

Petkovic:2008:MSS

[Pet08] Dušan Petkovic. Microsoft SQL Server 2008: a beginner’s guide. McGraw-Hill, New York, NY, USA, 2008. ISBN 0-07-154639-1 (electronic book). LCCN ????

Petrov:2019:DID

[Pet19] Alex Petrov. Database Internals: a Deep Dive Into How Distributed Data Systems Work. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2019. ISBN 1-4920-4034-7 (paperback), 1-4920-4030-4 (e-book), 1-4920-4031-2 (e-book). xix + 349 pp. LCCN QA76.9.D3 P48 2019.

URL http://www.vlebooks.com/
vleweb/product/openreader?id=
none\%26isbn=9781492040316.

Pribyl:2002:LOP

[PF02] Bill Pribyl and Steven Feuerstein.

Learning Oracle PL/SQL. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00180-0. xix + 401 pp. LCCN QA76.73.P258 P74 2002 Stacks. US$39.95.

URL http://safari.oreilly.

                                        com/0596001800;   http://www.
                                        oreilly.com/catalog/learnoracle.

Page:1998:UO

[PH98] William G. Page and Nathan Hughes. Using Oracle8. Que Corporation, Indianapolis, IN, USA, special edition edition, 1998. ISBN 0-7897-1345-4. xix + 929 pp. LCCN QA76.9.D3P334 1998.

Pratt:2006:GM

[PL06] Philip J. Pratt and Mary Z. Last. A guide to MySQL. Thomson Course Technology, Boston, MA, USA, 2006. ISBN 1-4188-3635-4 (paperback). xvi + 288 pp. LCCN QA76.73.S67 P728 2006.

Powell:2005:OSJ

[PMD05] Gavin Powell and Carol McCullough-Dieter. Oracle SQL jumpstart with examples. Elsevier/Digital Press, Amsterdam, The Netherlands; Boston, MA, USA, 2005. ISBN 1-55558-323-7. xxix + 652 pp. LCCN QA76.73.S67

                                        P62 2005. URL http://www.
                                        loc.gov/catdir/enhancements/
                                        fy0626/2005273537-d.html.

Plugge:2010:DGM

[PMH10] Eelco Plugge, Peter Membrey, and Tim Hawkins. The Definitive Guide to MongoDB: The NoSQL Database for Cloud and Desktop Computing. Apress, Berkeley, CA, USA, 2010. ISBN 1-4302-3051-7, 1-4302-3052-5 (e-book). xx + 307 pp. LCCN QA76.9.D3 P58 2010. URL http:

                                        //proquest.safaribooksonline.
                                        com/9781430230519.

Perkins:2002:P

[POS02] Jeff Perkins, Grzegorz Tł. Olȩdzki, and Zbigniew Tł. Szalbot. PostgreSQL. Mikom, Warszawa, Poland, 2002. ISBN 83-7279-232-1. 346 pp. Polish translation of [Per01].

PGDG:2007:PRM

[Pos07] PostgreSQL Global Development Group. PostgreSQL reference manual. Network Theory Ltd., Bristol, UK, 2007. ISBN 0-9546120-2-7 (vol. 1), 0-9546120-3-5 (vol. 2), 0-9546120-4-3 (vol. 3). x + 703 (vol.

1) pp. LCCN ????

Perng:1998:SLT

[PP98] Chang-Shing Perng and Douglass Stott Parker. SQL/LPP: a time series extension of SQL based on limited patience patterns. Technical report CSD 980034, UCLA Computer Science Department, Los Angeles, CA, USA, 1998. 26 pp.

Pratt:1991:GS

[Pra91] Philip J. Pratt. A guide to SQL.

boyd & fraser, One Corporate Place, Suite 205, Danvers, MA, USA 01923, USA, second edition, 1991. ISBN 0-87835-669-X. xiii + 203 pp. LCCN QA76.73.S67 P73 1991.

Pratt:2005:GS

[Pra05] Philip J. Pratt. A guide to SQL. Thomson/Course Technology, Boston, MA, USA, seventh edition, 2005. ISBN 0-619-21674-3 (paperback). xviii + 294 pp. LCCN QA76.73.S67 P73 2005.

Preston:2007:BR

[Pre07] W. Curtis Preston. Backup and recovery. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2007. ISBN 0-596-10246-1 (paperback). xxviii + 729 pp. LCCN QA76.9.B32

                                        P74 2007. URL http://www.
                                        loc.gov/catdir/enhancements/
                                        fy0715/2007273944-d.html;   http:
                                        //www.oreilly.com/catalog/9780596102463/
                                        index.html.

Preger:2012:OSP

[Pre12] Robert Preger. The Oracle story, part 1: 1977–1986. IEEE Annals of the History of Computing, 34(4): 51–57, October/December 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Price:2001:JPO

[Pri01] Jason Price. Java Programming with Oracle SQLJ. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, September 2001. ISBN 0-596-00087-1. xiii + 381 pp. LCCN QA76.73.J38 P74 2001. US$39.95. URL http://www.

                                        oreilly.com/catalog/orasqlj.

Price:2004:ODS

[Pri04] Jason Price. Oracle Database 10g SQL. McGraw-Hill/Osborne, Emeryville, CA, USA, 2004. ISBN 0-07-225829-2 (electronic book). xxvi + 576 pp. LCCN ????

Papa:1999:MSS

[PS99] Johnny Papa and Matthew Shepker. Microsoft SQL server 7.0 programming unleashed. Unleashed series. Howard W. Sams, Indianapolis, IN 46268, USA, 1999. ISBN 0-672-31293-X (paperback). xxii + 946 pp. LCCN QA76.9.D3 P37 1999.

Purba:1994:DCS

[Pur94] Sanjiv Purba. Developing client/ server systems using Sybase SQL Server system 10. John Wiley and Sons, Inc., New York, NY, USA, 1994. ISBN 0-471-06249-9 (paper). xviii + 396 pp. LCCN QA76.9.C55 P87 1994.

Rizzo:2005:PSS

[R^{+}05] Thomas Rizzo et al., editors. Pro SQL Server 2005. The expert’s voice in SQL server. Apress, Berkeley, CA, USA, 2005. ISBN 1-59059-477-0. xxvii + 674 pp. LCCN ????

Rankins:2007:MSS

[R^{+}07] Ray Rankins et al., editors. Microsoft SQL Server 2005: unleashed. Howard W. Sams, Indianapolis, IN 46268, USA, 2007. ISBN 0-672-32824-0. xl + 1985 pp. LCCN QA76.9.D3 M57365 2007. URL http://www.

loc.gov/catdir/toc/ecip0710/
2007005947.html.

Rankins:2011:MSS

[R^{+}11] Ray Rankins et al., editors. Microsoft SQL Server 2008 R2 unleashed. Howard W. Sams, Indianapolis, IN 46268, USA, 2011. ISBN 0-672-33056-3 (paperback). xlvi + 2451 pp. LCCN ???? US$38.41.

Rabeler:2001:MTK

[Rab01] Carl Rabeler. MCSE training kit: Microsoft SQL Server 2000 system administration. Microsoft Press, Bellevue, WA, USA, 2001. ISBN 0-7356-1426-1. iii + 207 pp. LCCN QA76.3 .R32 2001.

Rainardi:2008:BDW

[Rai08] Vincent Rainardi. Building a data warehouse: with examples in SQL Server. The expert’s voice in SQL server. Apress, Berkeley, CA, USA, 2008. ISBN 1-4302-0528-8, 1-59059-931-4. xvi + 523 pp. LCCN QA76.9.D37 R35 2008.

Rankins:1996:SSS

[Ran96] Ray Rankins. Sybase SQL Server 11 unleashed. Howard W. Sams, Indianapolis, IN 46268, USA, 1996. ISBN 0-672-30909-2. xxxvi + 1151 pp. LCCN QA76.9.C55 S94 1996.

Russel:2003:ODS

[RC03] Charlie Russel and Robert Cordingley. Oracle DBA SQL quick reference. Prentice Hall PTR Oracle series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2003. ISBN 0-13-140303-6 (paperback). xiv + 529 pp. LCCN QA76.9.D3 R869 2003.

Russel:2004:ODA

[RC04] Charlie Russel and Robert Cordingley. Oracle DBA automation quick reference. Prentice Hall Oracle DBA SQL quick reference series; Prentice Hall PTR Oracle series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2004. ISBN 0-13-140301-X (paperback). xvii + 291 pp. LCCN QA76.9.D3 R865 2004.

Rizzo:2006:PSS

[RDD^{+}06] Thomas Rizzo, Louis Davidson, Robin Dewson, Adam Machanic, Jan Narkiewicz, Joseph Sack, Julian Skinner, and Rob Walters. Pro SQL Server 2005. Springer eBook Collection Professional Computing and Web Design [Dig. Serial]; Springer [Dig. Serial]. Thomas Rizzo, Adam Machanic, Julian Skinner, Louis Davidson, Robin Dewson, Jan Narkiewicz, Joseph Sack, Rob Walters, Berkeley, CA, USA, 2006. ISBN 1-4302-0094-4. LCCN ????

Reese:1997:DPJ

[Ree97] George Reese. Database Programming With JDBC and Java. Java series. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, July 1, 1997. ISBN 1-56592-270-0. 224 pp. LCCN QA76.625 .R44 1997 Sci-Eng. US$29.95. URL http://www.

ora.com/catalog/javadata/;
http://www.oreilly.com/catalog/
9781565922709;   http://www.
oreilly.com/catalog/javadata.

Java 1.1.

Reese:2003:MPR

[Ree03] George Reese. MySQL pocket reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00446-X. vi + 87 pp. LCCN QA76.73.S67 R43 2003.

Reese:2007:MPR

[Ree07] George Reese. MySQL pocket reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2007. ISBN 0-596-51426-3 (paperback). vi + 126 pp. LCCN QA76.73.S67 R43 2007.

                                        URL http://www.oreilly.com/
                                        catalog/9780596514266.

Rehman:2003:IDS

[Reh03] Rafeeq Ur Rehman. Intrusion Detection with SNORT: Advanced IDS Techniques Using SNORT, Apache, MySQL, PHP, and ACID. Bruce Perens’ Open source series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2003. ISBN 0-13-140733-3. xii + 263 pp. LCCN TK5105.59 .R44 2003. US$39.99.

Ramakrishnan:2003:DMS

[RG03] Raghu Ramakrishnan and Johannes Gehrke. Database Management Systems. McGraw-Hill, New York, NY, USA, third edition, 2003. ISBN 0-07-246563-8, 0-07-115110-9 (ISE). xxxii + 1065 pp. LCCN QA76.9.D3 R237 2002.

Riordan:2005:DED

[Rio05] Rebecca Riordan. Designing effective database systems. Addison-Wesley, Reading, MA, USA, 2005. ISBN 0-321-29093-3 (paperback). xix + 353 pp. LCCN QA76.9.D26

                                        R54 2005. URL http://www.loc.
                                        gov/catdir/toc/ecip051/2004023319.
                                        html.

Rischert:2010:OSE

[Ris10] Alice Rischert. Oracle SQL by example. The Prentice Hall Oracle series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fourth edition, 2010. ISBN 0-13-714283-8 (paperback). xxxix + 913 pp. LCCN QA76.73.S67 R57 2010.

Riggs:2010:PAC

[RK10] Simon Riggs and Hannu Krosing. PostgreSQL 9 Admin Cookbook. Packt, Birmingham, UK, 2010. ISBN 1-84951-028-8. 366 (est) pp. LCCN ????

Rosenzweig:2000:OPI

[Ros00] Benjamin Rosenzweig. Oracle PL/ SQL interactive workbook. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-015743-0 (paperback). xxi + 447 pp. LCCN QA76.73.P258 R67 2000.

Rowe:2012:HIC

[Row12] Lawrence A. Rowe. History of the Ingres Corporation. IEEE Annals of the History of Computing, 34(4): 58–70, October/December 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Reilly:2001:SSD

[RP01] Michael (Michael D.) Reilly and Michelle Poolet. SQL Server 2000 design and T-SQL programming. Database professional’s library. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-212375-3. xxv + 612 pp. LCCN QA76.9.C55 R453 2001.

Rosenzweig:2000:OPS

[RR00] Benjamin Rosenzweig and Elena Silvestrova Rakhimov. Oracle PL/ SQL interactive workbook. The Prentice Hall PTR Oracle series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-015743-0 (paperback). xxi + 447 pp. LCCN QA76.73.P258 R67 2000.

Rosenzweig:2004:OPS

[RS04] Benjamin Rosenzweig and Ellen Silvestrova. Oracle PL/SQL by Example. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, third edition, 2004. ISBN 0-13-117261-1. xxiii + 735 pp. LCCN QA76.73.P258 R68 2004. US$49.99.

Robinson:1994:TSD

[RTRT94] Vincent B. Robinson, Henry Tom, Vincent B. Robinson, and Henry Tom, editors. Towards SQL database extensions for geographic information systems. Silicon Press, 25 Beverly Road, Summit, NJ 07901, USA, 1994. ISBN 0-929306-19-8. i + 148 pp. LCCN G70.2 .T69 1994.

Redmond:2012:SDS

[RW12] Eric Redmond and Jim R. Wilson. Seven databases in seven weeks: a guide to modern databases and the NoSQL movement. The pragmatic programmers. Pragmatic Bookshelf, Raleigh, NC, USA, 2012. ISBN 1-934356-92-1 (paperback). xiii + 333 pp. LCCN QA76.9.D32 .R43 2012. Edited by Jacquelyn Carter.

Reese:2002:MUM

[RYW02] George Reese, Randy Jay Yarger, and Tim King With Hugh E. Williams. Managing & Using MySQL. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 0-596-00211-4. 425 pp. LCCN QA76.73.S67 R43 2002. US$39.95.

URL http://safari.oreilly.
com/0596002114;   http://www.
oreilly.com/catalog/msql2.

Spofford:2006:MSM

[S^{+}06] George Spofford et al., editors.

MDX solutions with Microsoft SQL Server Analysis Services 2005 and Hyperion Essbase. John Wiley and Sons, Inc., New York, NY, USA, second edition, 2006. ISBN 0-471-74808-0 (paperback). ???? pp. LCCN QA76.9.C55 S756 2006; QA76.9.C55 S756 2006eb.

Sequeira:2001:SSS

[SA01] Anthony Sequeira and Brian Alderman. SQL Server on site. Coriolis Group Books, Scottsdale, AZ, USA, 2001. ISBN 1-58880-061-X. ???? pp. LCCN QA76.9.C55 S375 2001.

Sachs:1987:SUG

[Sac87] Jonathan Sachs. SQL*plus user’s guide: version 2.0. Oracle Corporation, Redwood Shores, CA, USA, revised edition, July 1987. various pp. LCCN AA0002183382. Part number 3201-V2.0.

Sanin:1997:CPA

[San97] Leo Sanin. Client/server programming with Access and SQL server: the integrated guide for programmers and developers. Ventana Press, Chapel Hill, NC, USA, 1997. ISBN 1-56604-693-9. xxii + 398 pp. LCCN QA76.9.C55 S26 1997.

SAS:1989:SGS

[SAS89] SAS Institute. SAS guide to the SQL procedure: usage and reference: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1989. ISBN 1-55544-367-2. xviii + 210 pp. LCCN QA276.4.S2336 1989.

SAS:1990:SAId

[SAS90] SAS Institute. SAS/ACCESS interface to SQL/DS: usage and reference: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1990. ISBN 1-55544-385-0. xx + 191 pp. LCCN QA276.4 .S217 1990.

SAS:1993:SAIc

[SAS93a] SAS Institute. SAS/ACCESS interface to ODBC: SQL procedure pass-through facility. SAS technical report; P-262. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-599-3. xii + 32 pp. LCCN ????

SAS:1993:SAIi

[SAS93b] SAS Institute. SAS/ACCESS interface to SYBASE and SQL Server: usage and reference, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-552-7. xxii + 258 pp. LCCN ????

SAS:1993:USQ

[SAS93c] SAS Institute. Using the SQL query window: release 6.08. Number P-254 in SAS technical report. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-563-2. xvi + 134 pp. LCCN QA276.4.U75 1993.

SAS:1994:SAIa

[SAS94a] SAS Institute. SAS/ACCESS interface to Informix technical report: SQL procedure pass-through facility: release 6.09. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-626-4. 31 pp. LCCN ????

SAS:1994:SASa

[SAS94b] SAS Institute. SAS/ACCESS software changes and enhancements: SQL procedure pass-through facility: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-660-4. viii + 44 pp. LCCN ????

SAS:1995:SGSa

[SAS95] SAS Institute. SAS guide to the SQL Query Window: usage and reference: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-266-8. iv + 167 pp. LCCN ????

Sayles:1989:SDS

[Say89a] Jonathan Sayles. SQL for DB2 and SQL/DS application developers. QED Information Sciences, Inc., P. O. Box 82-181, Wellesley, MA 02181, USA, 1989. ISBN 0-89435-303-9. xvii + 473 pp.

LCCN QA76.73.S67 S2898 1990. US$29.95.

Sayles:1989:SSH

[Say89b] Jonathan Sayles. SQL spoken here. QED Information Sciences, Inc., P. O. Box 82-181, Wellesley, MA 02181, USA, 1989. ISBN 0-89435-262-8. xiii + 271 pp. LCCN QA76.73.S67 S291 1989. US$39.95.

Sayles:1990:ESD

[Say90a] Jonathan Sayles. Embedded SQL for DB2: application design and programming. QED Information Sciences, Inc., P. O. Box 82-181, Wellesley, MA 02181, USA, 1990. ISBN 0-89435-308-X. xiii + 345 pp. LCCN QA76.73.S67 S287 1990. US$29.95.

Sayles:1990:HUO

[Say90b] Jonathan Sayles. How to use Oracle SQL*Plus. QED Information Sciences, Inc., P. O. Box 82-181, Wellesley, MA 02181, USA, 1990. ISBN 0-89435-291-1. xviii + 406 pp. LCCN QA76.73.S67 S289 1990. US$29.95.

Sayles:1990:SDS

[Say90c] Jonathan Sayles. SQL for DB2 and SQL/DS application developers. QED Information Sciences, Wellesley, Mass., 1990. ISBN 0-89435-303-9. xvii + 473 pp. LCCN ????

Sayles:1991:PPG

[Say91] Jonathan S. Sayles. PL/I programmer’s guide to embedded SQL for DB2. QED Information Sciences, Inc., P. O. Box 82-181, Wellesley, MA 02181, USA, 1991. ISBN 0-89435-346-2. xii + 372 pp. LCCN QA76.73.S67 S288 1991.

Sanin:1997:CSP

[SC97] Leo Sanin and Renzhong Chen.

Client/server programming with Access and SQL server: the integrated guide for programmers and developers. Ventana Press, Chapel Hill, NC, USA, 1997. ISBN 1-56604-693-9. xxii + 398 pp. LCCN QA76.9.C55 S26 1997.

Smith:2009:MSS

[SC09] Bryan C. Smith and C. Ryan Clay.

Microsoft SQL Server 2008 MDX step by step. Microsoft Press, Bellevue, WA, USA, 2009. ISBN 0-7356-2618-9. ???? pp. LCCN ????

Spainhour:2003:WPC

[SEG^{+}03] Stephen Spainhour, Robert Eckstein, Danny Goodman, Hugh E. Williams, David Lane, Rasmus Lerdorf, Kevin Tatroe, David Sklar, and Adam Trachtenberg, editors. The Web programming CD bookshelf. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, version 1.0. edition, 2003. ISBN 0-596-00510-5. LCCN TK5105.888.

Seidman:2001:DMM

[Sei01] Claude Seidman. Data mining with Microsoft SQL Server 2000: technical reference. Microsoft Press, Bellevue, WA, USA, 2001. ISBN 0-7356-1271-4. xiv + 367 pp. LCCN QA76.9.D343 S45 2001.

Sadalage:2013:NDB

[SF13] Pramod J. Sadalage and Martin Fowler. NoSQL distilled: a brief guide to the emerging world of polyglot persistence. Addison-Wesley, Reading, MA, USA, 2013. ISBN 0-321-82662-0 (paperback), 0-13-303613-8. xix + 164 pp. LCCN QA76.9.D32 S228 2013. URL http:

//proquest.safaribooksonline.
com/?fpi=9780133036138.

Sayles:1991:QHU

[SG91] Jonathan Sayles and Gabriel F. Gargiulo. QMF: how to use query management facility with DB2 and SQL/DS. QED IBM Mainframe series. QED Technical Publishing Group, Boston, MA, USA, 1991. ISBN 0-89435-376-4. xvii + 362 pp. LCCN QA76.9.D3G376 1991. US$39.95.

Spetic:2002:TSC

[SG02] Ales Spetic and Jonathan Gennick. Transact-SQL Cookbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 1-56592-756-7. xviii + 282 pp. LCCN QA76.73.S67 S64 2002 Stacks. US$34.95.

URL http://safari.oreilly.
com/1565927567;   http://www.
oreilly.com/catalog/transqlcook.

Steele:1990:PDI

[SH90] Philip Steele and Robert Heydt. Professional dBASE IV 1.1 and SQL programming. John Wiley and Sons, Inc., New York, NY, USA, 1990. ISBN 0-471-50985-X. xiv + 528 pp. LCCN QA76.9.D3 S739 1990. US$27.95.

Stephenson:1992:SST

[SH92] Peter Stephenson and Glenn Hartwig. SQL: self-teaching guide. John Wiley and Sons, Inc., New York, NY, USA, 1992. ISBN 0-471-54544-9. xii + 223 pp. LCCN QA76.73.S67 S74 1992. US$22.95.

Sheldon:2003:SBG

[SH03] Robert Sheldon and Forrest Houlette.

SQL: a beginner’s guide. Mc-Graw-Hill/Osborne, Emeryville, CA, USA, second edition, 2003. ISBN 0-07-222885-7. xxvii + 530 pp. LCCN QA76.73.S67 S54

2003.  URL http://www.loc.
gov/catdir/description/mh051/
2004270049.html;   http://www.
loc.gov/catdir/toc/mh041/2004270049.
html.

Shapiro:2000:SSC

[Sha00] Jeffrey R. Shapiro. SQL Server 2000: the complete reference. Database professional’s library. Osborne/McGraw-Hill, Berkeley, CA, USA, 2000. ISBN 0-07-212588-8. 931 pp. LCCN QA76.9.D3 S498 2000.

Shapiro:2001:SSC

[Sha01] Jeffrey R. Shapiro. SQL Server 2000: the complete reference. Database professional’s library. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-212588-8. xxviii + 931 pp. LCCN ????

Sharma:2002:MSS

[Sha02] Rahul Sharma. Microsoft SQL Server 2000: a guide to enhancements and new features. Addison-Wesley, Reading, MA, USA, 2002.

ISBN 0-201-75283-2. xxvii + 557 pp. LCCN QA76.9.C55 S5338 2002.

Shapiro:2007:MSS

[Sha07] Jeffrey R. Shapiro. Microsoft SQL Server 2005: the complete reference. Essential skills for database professionals. McGraw-Hill, New York, NY, USA, 2007. ISBN 0-07-226152-8 (paperback), (paperback). xxv + 645 pp. LCCN QA76.9.C55 S5332 2007. URL http://www.

loc.gov/catdir/enhancements/
fy0814/2007274366-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0814/2007274366-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0814/2007274366-t.html.

Sheldon:2000:MRR

[She00] Robert Sheldon. MCSE readiness review. Exam 70-029. Designing and implementing databases with Microsoft SQL Server 7.0. Microsoft Press, Bellevue, WA, USA, 2000. ISBN 0-585-39285-4 (electronic book), 0-7356-0673-

0. xvii + 475 pp. LCCN QA76.3 .S517 2000. URL http:

//www.netLibrary.com/urlapi.
asp?action=summary&v=1&bookid=
54073.

Shea:2003:RWS

[She03] Linchi Shea. Real world SQL Server administration with Perl. Apress, Berkeley, CA, USA, 2003. ISBN 1-59059-097-X. xxxi + 798 pp. LCCN ????

Sinclair:2000:ASS

[Sin00] Russell Sinclair. From Access to SQL Server. Books for professionals by professionals. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 1-893115-24-0 (paperback). xiii + 340 pp. LCCN QA76.73.S67 S56 2000.

Siple:1998:CGJ

[Sip98] Matthew D. Siple. The Complete Guide to Java Database Programming with FDBC: JDBC, ODBC & SQL. Java Masters Series. Mc-Graw-Hill, New York, NY, USA, January 1998. ISBN 0-07-913286-3 (softcover). LCCN QA76.625 .S557 1998. US$44.95. URL http:

//mcgraw-hill.inforonics.com/
cgi/getarec?mgh31395%comp.

Sippl:2013:IIM

[Sip13] Roger Sippl. Informix: Information management on Unix. IEEE Annals of the History of Computing, 35(2): 42–53, April/June 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Stones:2001:BDP

[SM01] Richard Stones and Neil Matthew.

Beginning Databases with PostgreSQL. Wrox Press, Chicago, IL, USA; Birmingham, UK, 2001. ISBN 1-86100-515-6. xi + 569 pp. LCCN ???? US$39.99, EUR 47.95.

Sheldon:2005:BM

[SM05] Robert Sheldon and Geoff Moes.

Beginning MySQL. Wrox beginning guides; Programmer to programmer. John Wiley and Sons, Inc., New York, NY, USA, 2005. ISBN 0-7645-7950-9. xxiv + 840 pp. LCCN ????

Smith:2010:PHP

[Smi10] Gregory Smith. PostgreSQL 9.0 High Performance. Packt, Birmingham, UK, 2010. ISBN 1-84951-030-X. 468 (est.) pp. LCCN ????

Snodgrass:2000:DTO

[Sno00] Richard T. Snodgrass. Developing time-oriented database applications in SQL. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2000. ISBN 1-55860-436-7. xxiii + 504 pp. LCCN QA76.73.S67 S56 2000. URL

                                        ftp://uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;   http:
                                        //www.cs.arizona.edu/people/
                                        rts/tdbbook.pdf;   http://www.
                                        loc.gov/catdir/description/els033/
                                        99014298.html;   http://www.
                                        loc.gov/catdir/toc/els033/99014298.
                                        html.

Solomon:1996:MSS

[Sol96] David S. Solomon. Microsoft SQL server 6 unleashed. Howard W. Sams, Indianapolis, IN 46268, USA, 1996. ISBN 0-672-30903-3. xxv + 958 pp. LCCN QA76.9.D3 M5737 1996.

Stephens:1998:STY

[SP98] Ryan K. Stephens and Ronald R. Plew. Sam’s teach yourself SQL in 24 hours. Howard W. Sams, Indianapolis, IN 46268, USA, 1998. ISBN 0-672-31162-3, 0-672-31245-X. xxvii + 398 pp. LCCN QA76.73.S67 S73 1998.

Stephens:2000:STY

[SP00] Ryan K. Stephens and Ronald R. Plew. Sams teach yourself SQL in 21 days. Howard W. Sams, Indianapolis, IN 46268, USA, third edition, 2000. ISBN 0-672-31674-9. xx + 754 pp. Rev. ed. of: Teach yourself SQL in 21 days. c1997.

Speshock:2002:MSS

[Spe02a] Carl H. Speshock. Microsoft SQL Server 2000 database administrators guidebook. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-061430-0. xxii + 446 pp. LCCN QA76.9.C55 S755 2002.

Spetic:2002:TCH

[Spe02b] Ales Spetic. Transact-SQL cookbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN ???? xviii + 282 pp. LCCN ????

Stephens:2008:STY

[SPJ08] Ryan K. Stephens, Ronald R. Plew, and Arie Jones. Sams teach yourself SQL in 24 hours. Sams teach yourself. Howard W. Sams, Indianapolis, IN 46268, USA, fourth edition, 2008. ISBN 0-7686-8277-0, 0-672-33018-0. xiv + 476 pp. LCCN ????

Solomon:1998:MSS

[SR98] David S. Solomon and Ray Rankins. Microsoft SQL server 6.5 unleashed. Howard W. Sams, Indianapolis, IN 46268, USA, third edition, 1998. ISBN 0-672-31190-9. xl + 1222 pp. LCCN QA76.9.D3 S64 1998.

Spenik:2003:MSS

[SS03] Mark Spenik and Orryn Sledge. Microsoft SQL Server 2000 DBA survival guide. Howard W. Sams, Indianapolis, IN 46268, USA, second edition, 2003. ISBN 0-672-32468-7. xxvii + 1001 pp. LCCN QA76.9.C55 S753 2003.

Simmons:2010:PSS

[SSS10] Ken Simmons, Colin Stasiuk, and Jorge Segarra. Pro SQL Server 2008 policy-based management. The expert’s voice in SQL server. Apress, Berkeley, CA, USA, 2010. ISBN 1-4302-2910-1. xvi + 252 pp. LCCN QA76.9.D3 S56384 2010. URL http:/

/proquest.safaribooksonline.
com/?uiCode=yaleu\%26xmlId=978143022910.

Stanek:2010:MSS

[Sta10] William R. Stanek. Microsoft SQL Server 2008: administrator’s pocket consultant. Microsoft Press, Bellevue, WA, USA, second edition, 2010. ISBN 0-7356-2738-X (paperback). xxiv + 710 pp. LCCN ????

Stephens:2000:SD

[Ste00] Ryan K. Stephens. SQL in 21 days. Sams teach yourself. Howard W. Sams, Indianapolis, IN 46268, USA, 2000. ISBN 0-672-31674-9 (paperback). xx + 754 pp. LCCN ????

Steers:2005:PUT

[Ste05] Kirk Steers. PC upgrading and troubleshooting: quicksteps. Mc-Graw-Hill/Osborne, Emeryville, CA, USA, 2005. ISBN 0-07-225961-2, 0-07-226414-4. xv + 192 pp. LCCN ????

Stevens:2009:HDD

[Ste09] Orrin Stevens. The history of Datacom/DB. IEEE Annals of the History of Computing, 31(4):87–91, October/December 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Stinson:2001:PER

[Sti01] Barry Stinson. PostgreSQL Essential Reference. New Riders Publishing, Carmel, IN, USA, 2001. ISBN 0-7357-1121-6. xxiii + 371 pp. LCCN QA76.9.D3 S747 2001. US$39.99.

Stonebraker:1988:RDS

[Sto88] Michael Stonebraker, editor. Readings in Database Systems. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1988. ISBN 0-934613-65-6. xii + 644 pp. LCCN QA76.9.D3 R4 1988. US$29.95.

Stonebraker:2011:SNE

[Sto11] Michael Stonebraker. Stonebraker on NoSQL and enterprises. Communications of the ACM, 54(8): 10–11, August 2011. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stucky:2001:MBU

[Stu01] Matthew Stucky. MySQL: building user interfaces. New Riders Publishing, Carmel, IN, USA, 2001. ISBN 0-7357-1049-x. xix + 632 pp. LCCN ????

Sunderic:2003:SSS

[Sun03] Dejan Sunderic. SQL Server 2000 stored procedure and XML programming. McGraw-Hill/Osborne, Emeryville, CA, USA, second edition, 2003. ISBN 0-07-222896-2. xvi + 679 pp. LCCN QA76.9.C55

                                        S83 2003. URL http://www.loc.
                                        gov/catdir/description/mh051/
                                        2004296506.html;   http://www.
                                        loc.gov/catdir/toc/mh041/2004296506.
                                        html.

Sunderic:2006:MSS

[Sun06] Dejan Sunderic. Microsoft SQL Server 2005 stored procedure programming in T-SQL and .NET. McGraw-Hill, New York, NY, USA, 2006. ISBN 0-07-226228-1. xxxvi + 664 pp. LCCN ????

Sunderic:2001:SSS

[SW01] Dejan Sunderic and Tom Woodhead. SQL Server 2000: stored procedure programming. Database professional’s library. Osborne/Mc-Graw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-212566-7. xvii + 732 pp. LCCN QA76.9.C55 S86 2001.

Sybase:1990:STP

[Syb90] Sybase, Inc. Sybase technical publications. Sybase, Inc., Emeryville, CA, USA, 1990. various pp. LCCN QA76.9.D3 S74 1990.

Sybex:2001:SSP

[Syb01] Sybex. SQL Server programming complete. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2001. ISBN 0-7821-2974-9. xxvi + 994 pp. LCCN QA76.9.D3 S688 2001. US$19.99.

Szeliga:2002:PCP

[Sze02] Marcin Szeliga. PostgreSQL 7.2: ćwiczenia praktyczne. (Polish) [PostgreSQL 7.2: ???].

Ćwiczenia Praktyczne. Helion, Gliwice, Poland, 2002. ISBN 83-7197-866-9. 150 + 2 pp.

Schwartz:2008:HPM

[SZT^{+}08] Baron Schwartz, Peter Zaitsev, Vadim Tkachenko, Jeremy D. Zawodny, Arjen Lentz, and Derek J. Balling. High performance MySQL. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2008. ISBN 0-596-10171-6. xxii + 684 pp. LCCN ????

Turley:2004:PSS

[T^{+}04] Paul Turley et al., editors. Professional SQL Server reporting services. John Wiley and Sons, Inc., New York, NY, USA, 2004. ISBN 0-7645-6878-7 (paper/website). xxxi + 585 pp. LCCN QA76.9.C55 P79 2004eb.

Turley:2006:PSS

[T^{+}06] Paul Turley et al., editors. Professional SQL Server 2005 reporting services. John Wiley and Sons, Inc., New York, NY, USA, 2006. ISBN 0-7645-8497-9 (paperback). ???? pp. LCCN QA76.9.C55 P79 2006eb.

Turley:2008:PMS

[T^{+}08] Paul Turley et al., editors. Professional Microsoft SQL Server 2008 reporting services. John Wiley and Sons, Inc., New York, NY, USA, 2008. ISBN 0-470-24201-9. ???? pp. LCCN QA76.9.D3 P7663 2008.

Tare:1989:DPU

[Tar89] Ramkrishna S. Tare. Data processing in the UNIX environment: with INFORMIX-SQL, Embedded-SQL, C-ISAM, and TURBO. Computing that works. McGraw-Hill, New York, NY, USA, 1989. ISBN 0-07-062885-8. xxi + 438 pp. LCCN QA76.9.D3 T374 1989.

Taylor:2001:SWC

[Tay01] Allen G. Taylor. SQL weekend crash course. Hungry Minds, Indianapolis, IN, USA, 2001. ISBN 0-7645-4901-4. xxiii + 332 pp. LCCN QA76.73.S67 T393 2002. URL

ftp://uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;   http:
//www.loc.gov/catdir/description/
wiley038/2001098038.html;   http:
//www.loc.gov/catdir/toc/wiley021/
2001098038.html.

Taylor:2010:SD

[Tay10] Allen G. Taylor. SQL for dummies. –For dummies. John Wiley and Sons, Inc., New York, NY, USA, seventh edition, 2010. ISBN 0-470-55741-9, 0-470-59314-8. xvi + 440 pp. LCCN QA76.73.S67

T39 2010. URL http://site.
ebrary.com/lib/princeton/Doc?
id=10361296.

Trimble:1989:VIS

[TC89] J. Harvey Trimble and David Chappell. A visual introduction to SQL. John Wiley and Sons, Inc., New York, NY, USA, 1989. ISBN 0-471-61684-2. viii + 254 pp. LCCN QA76.73.S67 T75 1989.

Tennick:2010:PDQ

[Ten10] Art Tennick. Practical DMX queries for Microsoft SQL Server Analysis Services 2008. McGraw-Hill/Osborne, Emeryville, CA, USA, 2010. ISBN 0-07-174866-0 (paperback). xxvi + 308 pp. LCCN ???? US$37.99.

Tennick:2011:PSQ

[Ten11] Art Tennick. Practical SQL queries for Microsoft SQL Server 2008 R2. McGraw-Hill, New York, NY, USA, 2011. ISBN 0-07-174687-0 (paperback). xxx + 304 pp. LCCN QA76.73.S67 T44 2011.

Thomas:2003:OXC

[Tho03] Michael D. Thomas. Oracle XSQL:

combining SQL, Oracle text, XSLT, and Java to publish dynamic Web content. John Wiley and Sons, Inc., New York, NY, USA, 2003. ISBN 0-471-27120-9 (paper/website). xix + 572 pp. LCCN QA76.73.S67

T48 2003. URL http://www.
loc.gov/catdir/bios/wiley045/
2002014913.html;   http://www.
loc.gov/catdir/description/wiley038/
2002014913.html;   http://www.
loc.gov/catdir/toc/wiley031/
2002014913.html.

Tiwari:2011:PN

[Tiw11] Shashank Tiwari. Professional NoSQL. Wrox Programmer to Programmer. John Wiley and Sons, Inc., New York, NY, USA, 2011. ISBN 0-470-94224-X (print), 1-118-16780-5, 1-118-16781-3 (ebook). xxi + 361 pp. LCCN QA76 .T59 2011. URL http:

//proquest.safaribooksonline.
com/9781118167809.

Trueblood:2001:DMS

[TL01] Robert P. Trueblood and John N.

Lovett. Data mining and statistical analysis using SQL. Books for professionals by professionals. Apress, Berkeley, CA, USA, 2001. ISBN 1-893115-54-2. xvi + 410 pp. LCCN QA76.9.D343 T78 2001.

Tang:2005:DMS

[TM05] ZhaoHui Tang and Jamie MacLennan. Data mining with SQL Server 2005. John Wiley and Sons, Inc., New York, NY, USA, 2005. ISBN 0-471-46261-6 (paperback), 0-471-75468-4 (electronic book). xix + 460 pp. LCCN QA76.9.D343 T36 2005.

Tow:2003:ST

[Tow03a] Dan Tow. SQL tuning. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00573-3. xix + 314 pp. LCCN QA76.73.S67 T69 2003.

Tow:2003:STH

[Tow03b] Dan Tow. SQL tuning. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN ???? xix + 314 pp. LCCN ????

Trachtenberg:2004:UP

[Tra04] Adam Trachtenberg. Upgrading to PHP 5. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00636-5. xviii + 327 pp. LCCN QA76.73.P224 T73

                                        2004. URL http://www.oreilly.
                                        com/catalog/9780596006365.

Trezzo:1999:OPS

[Tre99] Joseph C. Trezzo. Oracle PL/ SQL: tips and techniques. Osborne/ McGraw-Hill, Berkeley, CA, USA, 1999. ISBN 0-07-882438-9. xxix + 943 pp. LCCN ????

Thiruvathukal:2002:WPT

[TSC02] George K. (George Kuriakose)

Thiruvathukal, John P. Shafaee, and Thomas W. Christopher. Web programming: techniques for integrating Python, Linux, Apache, and MySQL. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-041065-9. xviii + 745 pp. LCCN QA76.625 .T48

2002. URL http://www.phptr.
com/ptrbooks/ptr_0130410659.
html.

Turley:2006:BTS

[TW06] Paul Turley and Dan Wood. Beginning Transact-SQL with SQL Server 2000 and 2005. Wrox beginning guides. John Wiley and Sons, Inc., New York, NY, USA, 2006. ISBN 0-7645-7955-X (paper/website). xxii + 570 pp. LCCN QA76.9.D3 T855 2006. URL http://www.

loc.gov/catdir/enhancements/
fy0622/2005020208-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0653/2005020208-b.html;   http:
//www.loc.gov/catdir/toc/ecip0516/
2005020208.html.

Tahaghoghi:2007:LM

[TW07] Seyed M. M. Tahaghoghi and Hugh E. Williams. Learning MySQL. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2007. ISBN 0-596-00864-3 (paperback). xvii + 598 pp. LCCN QA76.7.S67; QA76.7.S67 T34 2007; QA76.7.S67 T34 2007eb. URL http://www.

                                        loc.gov/catdir/toc/fy0706/2007272020.
                                        html.

USNBS:1987:DLS

[Uni87] United States National Bureau of Standards. Database language SQL. National Technical Information Service, Washington, DC, USA, March 10, 1987. 5 pp. Shipping list number 87-330-P.

Urman:1996:OPP

[Urm96a] Scott Urman. Oracle PL/SQL programming. Osborne/McGraw-Hill, Berkeley, CA, USA, 1996. ISBN 078821762 (paperback). xxvi + 564 pp. LCCN ????

Urman:1996:OPS

[Urm96b] Scott Urman. Oracle PL/SQL programming. Oracle Press, Berkeley, CA, USA, 1996. ISBN 0-07-882176-

2. xxvi + 564 pp. LCCN QA76.9.D3

U75 1996.

Urman:1997:OPS

[Urm97] Scott Urman. Oracle8 PL/SQL programming: The essential guide for every Oracle programmer: with new information on performance and tuning. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1997. ISBN 0-07-882305-6. xxxi + 1022 pp. LCCN QA76.9.D3U758 1997.

                                        URL http://www.osborne.com/
                                        oracle/index.htm. Covers releases

7 through 8.

Urman:2000:OAP

[Urm00] Scott Urman. Oracle8i advanced PL/SQL programming. Osborne/ McGraw-Hill, Berkeley, CA, USA, 2000. ISBN 0-07-212144-0, 0-07-212146-7. xxii + 772 pp. LCCN QA76.73.P258 U76 2000.

Urman:2004:ODP

[Urm04] Scott Urman. Oracle Database 10g PL/SQL Programming. Mc-Graw-Hill/Osborne, Emeryville, CA, USA, 2004. ISBN 0-07-223066-5 (paperback). xxvii + 862 pp. LCCN ????

Utley:2000:SSW

[Utl00] Craig Utley. SQL Server 2000: Web application developer’s guide. Osborne/McGraw-Hill, Berkeley, CA, USA, 2000. ISBN 0-07-212619-1. xxiii + 663 pp. LCCN QA76.9.C55 U88 2000.

Valade:2005:PME

[Val05] Janet Valade. PHP and MySQL everyday apps for dummies. John Wiley and Sons, Inc., New York, NY, USA, 2005. ISBN 6-61027-778-8 (electronic book), 0-7645-7587-2. xviii + 436 pp. LCCN QA76.73.P224 VAL.

Valade:2006:PMY

[Val06] Janet Valade. PHP and MySQL:

your visual blueprint for creating dynamic, database driven Web sites. John Wiley and Sons, Inc., New York, NY, USA, 2006. ISBN 0-470-04839-5 (paperback). xii + 307 pp. LCCN QA76.73.P224 V353 2006.

Valade:2007:PMD

[Val07] Janet Valade. PHP and MySQL for dummies. For dummies. John Wiley and Sons, Inc., New York, NY, USA, third edition, 2007. ISBN 0-470-12114-9 (electronic book). xviii + 436 pp. LCCN QA76.73.P224 V35 2007eb.

Vang:1991:SRD

[Van91] Soren Vang. SQL and relational databases. The Data based advisor series. Microtrend Books, San Marcos, CA, USA, 1991. ISBN 0-915391-42-2. xvi + 328 pp. LCCN QA76.73.S67 V3613 1991. US$24.95, CDN$32.95. Translation of: Relationsdatabaser og SQL.

Vaswani:2005:HDE

[Vas05] Vikram Vaswani. How to do everything with PHP and MySQL. McGraw-Hill/Osborne, Emeryville, CA, USA, 2005. ISBN 0-07-146654-1, 0-07-225795-4. xvi + 381 pp. LCCN ????

Vaughn:2007:HGV

[VB07] William R. Vaughn and Peter Blackburn. Hitchhiker’s guide to Visual Studio and SQL server: best practice architectures and examples. Microsoft Windows server system series. Addison-Wesley, Reading, MA, USA, seventh edition, 2007. ISBN 0-321-24362-5 (paperback), (paperback). xxiv + 1095 pp. LCCN QA76.76.A65

V39 2007. URL http://www.
loc.gov/catdir/toc/ecip0620/
2006030355.html.

Varga:2016:IMS

[VCD16] Stacia Varga, Denny Cherry, and Joseph D’Antoni. Introducing Microsoft SQL Server 2016: missioncritical applications, depper insights, hyperscale cloud. Microsoft Press, Bellevue, WA, USA, 2016. ISBN 1-5093-0195-X. xi + 200 pp. LCCN QA76.9.C55. URL http:

//proquest.safaribooksonline.
com/9781509301959;   http://
proquest.safaribooksonline.com/
?fpi=9781509301959;   https:/
/docs.microsoft.com/en-us/sql/
linux/sql-server-linux-setup-
red-hat.

vanderLans:1988:IS

[vdL88] Rick F. van der Lans. Introduction to SQL. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-17521-5 (paperback). x + 348 pp. LCCN QA76.73.S67 L3613 1988. UK£15.95.

vanderLans:1989:SSC

[vdL89] Rick F. van der Lans. The SQL standard: a complete reference. Academic Service, Postbus 81, 2870 AB Schoonhoven, The Netherlands, 1989. ISBN 0-13-840059-8. x + 140 pp. LCCN QA76.73.S67 L37 1989. US$40.00. Translated from the Dutch.

vanderLans:1991:SGD

[vdL91] Rick F. van der Lans. An SQL guide for dBASE IV. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-54410-5. xiii + 362 pp. LCCN QA76.9.D3 L36613 1991. UK£17.95.

vanderLans:1992:SGO

[vdL92] Rick F. van der Lans. The SQL guide to Oracle. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-56545-5. xviii + 540 pp. LCCN QA76.73.S67 L358 1992. Translated from Dutch.

vanderLans:1993:IS

[vdL93] Rick F. van der Lans. Introduction to SQL. Addison-Wesley, Reading, MA, USA, second edition, 1993. ISBN 0-201-62425-7. xix + 471 pp. LCCN QA76.73.S67 L3613 1993.

vanderLans:1998:SGO

[vdL98] Rick F. van der Lans. The SQL guide to Oracle. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-56545-5. xviii + 540 pp. LCCN QA76.73.S67 L35 1996. Translated from Dutch.

vanderLans:2007:ISM

[vdL07a] Rick F. van der Lans. Introduction to SQL: mastering the relational database language. Addison-Wesley, Reading, MA, USA, fourth (20th anniversary) edition, 2007. ISBN 0-321-30596-5 (paperback), (paperback). xxvi + 1026 pp. LCCN QA76.73.S67 L3613 2007.

vanderLans:2007:SMD

[vdL07b] Rick F. van der Lans. SQL for MySQL developers: a comprehensive tutorial and reference. Addison-Wesley, Reading, MA, USA, 2007. ISBN 0-13-149735-9 (paperback). xxv + 1004 pp. LCCN QA76.73.S67 L345 2007. US$49.99,

                                        US$35.99. URL http://www.loc.
                                        gov/catdir/toc/ecip078/2007000578.
                                        html.

Veerman:2008:MSP

[Vee08] Erik Veerman, editor. MCTS selfpaced training kit (exam 70-445): Microsoft SQL Server 2005 business intelligence — implementation and maintenance. Microsoft Press, Bellevue, WA, USA, 2008. ISBN 0-7356-2341-4. xxvi + 568 pp. LCCN QA76.3 .M334863 2008.

Viescas:2008:SQM

[VH08] John Viescas and Michael J.

(Michael James) Hernandez. SQL queries for mere mortals: a handson guide to data manipulation in SQL. For mere mortals series. Addison-Wesley, Reading, MA, USA, second edition, 2008. ISBN 0-321-44443-4 (paperback). xxxiii + 631 pp. LCCN QA76.73.S67 H48 2008.

Viescas:2014:SQM

[VH14] John Viescas and Michael J.

(Michael James) Hernandez. SQL queries for mere mortals: a handson guide to data manipulation in SQL. Addison-Wesley, Reading, MA, USA, third edition, 2014. ISBN 0-321-99247-4 (paperback), 0-13-382488-8. xxx + 760 pp. LCCN QA76.73.S67 H48 2014.

Vieira:2000:PSS

[Vie00] Robert Vieira. Professional SQL Server 2000 programming. Programmer to programmer. Wrox Press, Chicago, IL, USA; Birmingham, UK, 2000. ISBN 1-86100-448-6. xxiii + 1390 pp. LCCN QA76.9.D3 V525 2000.

Vieira:2006:BSS

[Vie06] Robert Vieira. Beginning SQL Server 2005 programming. Programmer to programmer; Wrox professional guides. John Wiley and Sons, Inc., New York, NY, USA, 2006. ISBN 0-7645-8433-2 (paperback). xxv + 688 pp.

LCCN QA76.9.C55 V53 2006. US$25.99. URL http://www.

loc.gov/catdir/enhancements/
fy0653/2006297436-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0658/2006297436-t.html;   http:
//www.loc.gov/catdir/enhancements/
fy0740/2006297436-b.html.

Vieira:2007:PSS

[Vie07] Robert Vieira. Professional SQL Server 2005 programming. Wrox professional guides. John Wiley and Sons, Inc., New York, NY, USA, 2007. ISBN 0-7645-8434-0 (paper/website). xxvii + 884 pp. LCCN QA76.9.C55

V55 2007. URL http://www.
loc.gov/catdir/enhancements/
fy0740/2006023047-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0740/2006023047-d.html;   http:
//www.loc.gov/catdir/toc/ecip0617/
2006023047.html.

Widenius:2002:MRM

[WA02] Michael “Monty” Widenius and David Axmark. MySQL Reference Manual: Documentation from the Source. O’Reilly Community Press, Sebastopol, CA, USA, 2002. ISBN 0-596-00265-3. x + 802 pp. LCCN QA76.73.S67 W54 2002 Stacks. US$39.95. URL http://www.

oreilly.com/catalog/mysqlref.

Wade:2012:CSS

[Wad12] Bradford W. Wade. Compiling SQL into System/370 machine language. IEEE Annals of the History of Computing, 34(4):49–50, October/ December 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Walker:2000:LDD

[Wal00] Damond Walker. Linux database design and development with PostgreSQL. Macmillan Computer Publishing, Indianapolis, IN, USA, 2000. ISBN 0-672-31866-0. ???? pp. LCCN ???? Edited by Don Roche.

Watt:2006:MSS

[Wat06] Andrew Watt. Microsoft SQL Server 2005 for dummies. –For dummies. John Wiley and Sons, Inc., New York, NY, USA, 2006. ISBN 0-7645-7755-7 (paperback). xx + 406 pp. LCCN QA76.9.C55 W375 2006. URL http://www.

loc.gov/catdir/enhancements/
fy0629/2005935163-d.html;   http:
//www.loc.gov/catdir/toc/fy0614/
2005935163.html.

Wilton:2005:BSH

[WC05] Paul Wilton and John R. Colby.

Beginning SQL. John Wiley and Sons, Inc., New York, NY, USA, 2005. ISBN 0-7645-7732-8 (paperback). xvii + 501 pp. LCCN QA76.73.S67 W57 2005b.

Wade:2012:IRD

[WC12] Bradford W. Wade and Donald D. Chamberlin. IBM relational database systems: The early years. IEEE Annals of the History of Computing, 34(4):38–48, October/ December 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Worsley:2002:PPP

[WD02a] John C. Worsley and Joshua D.

Drake. PostgreSQL par la pratique. O’Reilly, Paris, France, 2002.

ISBN 2-84177-211-X. xiv + 614 pp. LCCN ????

Worsley:2002:PPPb

[WD02b] John C. Worsley and Joshua D. Drake. PostgreSQL par la pratique. O’Reilly, Paris, France, 2002. ISBN 2-84177-211-X. xiv + 614 pp. LCCN ????

Worsley:2002:PP

[WD02c] John C. Worsley and Joshua D. Drake. Practical PostgreSQL. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 1-56592-846-6. xiv + 619 pp. LCCN QA76.9.D3 W67 2002; QA76.9.D3 W72 2002. US$44.95.

URL http://www.oreilly.com/
catalog/ppostgresql. CD-ROM

contains LXP version 0.8.0 and PostgreSQL version 7.1.3.

Worsley:2002:PPPa

[WDM02] John C. Worsley, Joshua D. Drake, and Radosław Tł. Meryk. PostgreSQL: praktyczny przewodnik. (Polish) [Practical PostgreSQL]. Helion, Gliwice, Poland, 2002. ISBN 83-7197-754-9. 617 pp.

Weissinger:2000:AND

[Wei00] A. Keyton Weissinger. ASP in a Nutshell: a Desktop Quick Reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, July 2000. ISBN 1-56592-843-1. xv + 473 pp. LCCN TK5105.8885.A26 W45 2000. US$29.95. URL http://www.

oreilly.com/catalog/9781565928435;

 http://www.oreilly.com/catalog/
aspnut2.

Weinstabl:2004:PAE

[Wei04] Paul Weinstabl. PostgreSQL: [Administration und Einsatz ; für Linux und Windows]. Computer and Literatur. Computer- und Literatur-Verlag, Böblingen, Germany, 2004. ISBN 3-936546-22-3. 589 pp. LCCN ???? EUR 49.90.

Wellesley:1991:LS

[Wel91] Wellesley Software. Learning SQL.

Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1991. ISBN 0-13-528704-9. viii, 231 pp. LCCN QA76.73.S67 L4 1991.

Wellesley:1992:LAS

[Wel92] Wellesley Software. Learning advanced SQL. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1992. ISBN 0-13-528712-X. ix + 213 pp. LCCN QA76.73.S67 L39 1992.

Wells:2001:CCS

[Wel01] Garth Wells. Code-centric: T-SQL programming with stored procedures and triggers. Books for professionals by professionals. Apress, Berkeley, CA, USA, 2001. ISBN 1-893115-83-6 (paperback). xxiii + 692 pp. LCCN QA76.9.C55 W448 2001.

Weller:2010:BSS

[WF10] Bart Weller and Fabio Claudio Ferrachiatti. Beginning SQL Server modeling: model-driven application development in SQL Server 2008. The expert’s voice in SQL server. Apress, Berkeley, CA, USA, 2010. ISBN 1-4302-2751-6. LCCN QA76.9.D26 W45 2010.

Weinberg:2010:SCR

[WGOD10] Paul N. Weinberg, James R. Groff, Andrew J. Oppel, and Aaron Davenport. SQL, the complete reference. McGraw-Hill, New York, NY, USA, third edition, 2010. ISBN 0-07-159255-5. xxviii + 882 pp. LCCN QA76.73.S67 G758 2010.

Whitehead:2006:SSB

[Whi06] Paul Whitehead. SQL Server2005 Black Book. Pagina Förlags AB, Box 706, Travgatan 92, 194 27 Upplands-Väsby, Sweden, 2006. ISBN 1-933097-18-3. 900 (est.) pp. LCCN ????

White:2012:HDG

[Whi12] Tom (Tom E.) White. Hadoop: the definitive guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2012. ISBN 1-4493-1152-0 (paperback). xxiii + 657 pp. LCCN ???? EUR 41.00; UK £35.50. URL

                                        http://www.gbv.de/dms/weimar/
                                        toc/688614469\_toc.pdf.

Wildermuth:2006:PSS

[Wil06] Shawn Wildermuth. Programming SQL Server 2005. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN ???? xi + 571 pp. LCCN QA76.73.S67 W55 2006.

Wipper:1989:GDS

[Wip89] Fritz Wipper. Guide to DB2 and SQL/DS. McGraw-Hill, New York, NY, USA, 1989. ISBN 0-07-055170-

7. xiii + 273 pp. LCCN QA76.9.D3 W585 1989.

Webb:1991:ODS

[WL91] Kenneth Webb and Lori Lafreniere. Oracle distributed systems: a C programmer’s development guide. Windcrest/McGraw-Hill, Blue Ridge Summit, PA, USA, 1991. ISBN 0-8306-7774-7. xvi + 329 pp. LCCN QA76.73 .S67 W43 1991. US$34.95.

Williams:2001:WDA

[WL01] Hugh E. Williams and David Lane. Web Database Applications with PHP & MySQL. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2001. ISBN 0-596-00041-

3. 563 pp. LCCN QA76.73.P224 W55 2001 Stacks. US$44.95.

URL http://safari.oreilly.
com/0596000413;   http://www.
oreilly.com/catalog/webdbapps.

Williams:2002:WDA

[WL02] Hugh E. Williams and David (David John) Lane. Web database applications with PHP and MySQL. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00041-3. xvi + 563 pp. LCCN QA76.73.P224 W55 2002.

US$44.95. URL http://safari.
oreilly.com/0596000413;   http:
//www.oreilly.com/catalog/webdbapps.

Williams:2004:WDA

[WL04] Hugh E. Williams and David (David John) Lane. Web database applications with PHP and MySQL. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2004. ISBN 0-596-00543-1. xviii + 796 pp. LCCN QA76.73.P224 W55 2004.

Wood:2007:BSS

[WLT07] Dan Wood, Chris Leiter, and Paul Turley. Beginning SQL Server 2005 administration. Wrox beginning guides. Wrox-Wiley Pub., Indianapolis, IN, USA, 2007. ISBN 0-470-04704-6 (paper/website). xxvi + 574 pp. LCCN QA76.9.D3 W6642 2007. URL http://www.

loc.gov/catdir/enhancements/
fy0741/2006030328-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0741/2006030328-d.html;   http:
//www.loc.gov/catdir/toc/ecip0620/
2006030328.html.

Wood:1995:UWS

[Woo95] Charles A. Wood. Using Watcom SQL. Client/server library. Que Corporation, Indianapolis, IN, USA, special edition, 1995. xxix + 813 pp. LCCN QA76.73.S67 W66 1995.

Waymire:1999:STY

[WS99] Richard Waymire and Rick Sawtell. Sams teach yourself Microsoft SQL server 7.0 in 21 days. Howard W. Sams, Indianapolis, IN 46268, USA, 1999. ISBN 0-672-31290-5. xxi + 772 pp.

Wang:2020:CSH

[WSS^{+}20] Y. Wang, Y. Shen, C. Su, J. Ma, L. Liu, and X. Dong. Crypt-SQLite: SQLite with high data security. IEEE Transactions on Computers, 69(5):666–678, 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/document/8946540.

Welling:2001:PMW

[WT01] Luke Welling and Laura Thomson. PHP and MySQL Web development. Howard W. Sams, Indianapolis, IN 46268, USA, 2001. ISBN 0-672-31784-2. xxii + 867 pp. LCCN ????

Wunschiers:2013:CBP

[Wün13] Röbbe Wünschiers. Computational biology: a practical introduction to biodata processing and analysis with Linux, MySQL, and R. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2013. ISBN 3-642-34748-7, 3-642-34749-5 (e-book). xxix + 449 pp. LCCN QH324.2 .W86 2013.

Wynkoop:1999:UMS

[Wyn99] Stephen Wynkoop. Using Microsoft SQL server 7.0. Que Corporation, Indianapolis, IN, USA, special edition, 1999. ISBN 0-7897-1523-6. xvi + 766 pp. LCCN QA76.9.D3 W963 1999.

XOpen:1994:XOC

[X/O94] X/Open, editor. X/Open CAE specification. X/Open Co., Reading, Berkshire, UK, 1994. ISBN 1-872630-58-8, 1-872630-98-7, 0-13-353558-4 (Prentice Hall). xvi + 156, x, 66 pp. LCCN QA76.73.S67 X18 1994.

Yadava:2007:BDB

[Yad07] Himanshu Yadava. The Berkeley DB book: the definite guide to creating scalable and fault-tolerant, embedded database applications using the open source Berkeley DB library. The expert’s voice in open source. Apress, Berkeley, CA, USA, 2007. ISBN 1-59059-672-2. xx + 442 pp. LCCN QA76.9.D3 Y33 2007.

Yank:2004:BYO

[Yan04] Kevin Yank. Build your own database driven website using PHP and MySQL. SitePoint Pty. Ltd., Collingwood, VIC, Australia, third edition, 2004. ISBN 0-9752402-1-8. xiv + 359 pp. LCCN TK5105.888 .Y36 2004; TK5105.888 .Y25 2004.

                                        URL http://www.oreilly.com/
                                        catalog/9780975240212.

Yazici:2006:PIP

[Yaz06] Volkan Yazıcı. PostgreSQL ile Programlama. ????, ????, Turkey,

2006.  ISBN ????  160
pp.  LCCN ????  URL

                                        http://www.students.itu.edu.
                                        tr/~yazicivo/doc/postgresql-
                                        ile-programlama-20060214.pdf.

Yarger:1999:MM

[YRK99] Randy Jay Yarger, George Reese, and Tim King. MySQL and mSQL. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-434-7. xv + 487 pp. LCCN QA76.73.S67 Y37 1999. US$34.95.

Yazici:2001:OPG

[YŞ01] Ali Yazıcı and Aydın Şekihanov.

Oracle 8i: a practical guide to SQL, PL/SQL, DEVELOPER 6. Atılım University Publications, Ankara, Turkey, 2001. ISBN 975-6707-02-X. 356 (est.) pp. LCCN

QA76.9. URL http://library.
atilim.edu.tr/opgtspd.htm.

Zawodny:2004:HPM

[ZB04] Jeremy D. Zawodny and Derek J.

Balling. High performance MySQL: optimization, backups, replication, and load balancing. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2004. ISBN 0-596-00306-4. xvi + 276 pp. LCCN QA76.73.S67 Z39 2004.